### v1

# On the geometry of spaces of filtrations on local rings

Abstract.
We study the geometry of spaces of fitrations on a Noetherian local domain. We introduce a metric $d_{1}$ on the space of saturated filtrations, inspired by the Darvas metric in complex geometry, such that it is a geodesic metric space. In the toric case, using Newton-Okounkov bodies, we identify the space of saturated monomial filtrations with a subspace of $L^{1}_{\mathrm{loc}}$. We also consider several other topologies on such spaces and study the semi-continuity of the log canonical threshold function in the spirit of Kollár-Demailly. Moreover, there is a natural lattice structure on the space of saturated filtrations, which is a generalization of the classical result that the ideals of a ring form a lattice.
1. Introduction
Throughout this paper, we work with a Noetherian local domain $(R,\mathfrak{m},\kappa)$ which is analytically irreducible, that is, the $\mathfrak{m}$-adic completion $\hat{R}$ is a domain.
As a generalization of ideals, graded sequences of ideals have been extensively studied in the past few years, and have played an important roles in many areas of algebraic geometry, see for example [ELS03, JM12, Cut13] and [BFJ08, BdFF12, BFJ14, Li18, Blu18, Liu18, Xu20, XZ21, LXZ22].
Filtrations on a local ring $(R,\mathfrak{m})$ are the continuously indexed version of graded sequences of ideals. To be more precise, an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ on $R$ is a collection $\{\mathfrak{a}_{\lambda}\}_{\lambda\in\mathbb{R}_{>0}}$ of $\mathfrak{m}$-primary ideals, which is decreasing, multiplicative and left continuous. The three conditions mean $\mathfrak{a}_{\lambda}\subset\mathfrak{a}_{\mu}$ for $\lambda>\mu$, $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ and $\mathfrak{a}_{\lambda-\epsilon}=\mathfrak{a}_{\lambda}$ for $0<\epsilon\ll 1$, respectively.
In this paper, we study the geometry of certain spaces of such filtrations and prove some structural results. The first main result is that there is a pseudometric $d_{1}$ on the set of $\mathfrak{m}$-filtrations with positive multiplicity, which is an analogue of the Darvas metric introduced in [Dar15] in complex geometry (see [BJ21] for the non-archimedean setting), such that a suitable retraction, the set of saturated filtrations, is a geodesic metric space. In the toric setting, the subspace of saturated monomial filtrations can be identified with the space of cobounded convex sets of the dual cone with the symmetric difference metric, which is naturally a subspace of $L^{1}_{\mathrm{loc}}$. We also introduce another metric, denoted by $d_{\infty}$, and several weak topologies, and study the continuity properties of the log canonical threshold function, following Demailly-Kollár [DK01]. Several spaces of filtrations have a natural structure of a lattice, generalizing a classical result that the ideals of a ring form a modular lattice.
1.1. The metric geometry of space of filtrations
1.1.1. The Darvas metric, multiplicities and saturations
Given a compact Kähler manifold $(X,\omega)$, Darvas [Dar15] introduced various Finsler metrics on the space $\mathcal{H}$ of smooth Kähler potentials, generalizing the metric studied in [Mab87]. One of the metrics, known as the Darvas metric $d_{1}$, has found numerous applications in complex geometry. In particular, the geodesic metric space $(\mathcal{H},d_{1})$ and its completion, $(\mathcal{E}^{1},d_{1})$, the space of Kähler potentials of finite energy, are closely related to Kähler-Einstein metrics. See for example [DR17, BBEGZ19, DL20]. A comprehensive survey on this metric space and its applications, particularly in relation to Tian’s Properness Conjecture, can be found in [Dar19].
More recently, Boucksom and Jonsson [BJ21] developed a similar construction in the non-archimedean setting , where they introduced a pseudo-metric $d_{1}$, also referred to as the Darvas metric, on the space $\mathcal{N}$ of norms on the section ring $R(X,-mrK_{X})$ of a projective Fano variety $X$. Geodesics on this space have been independently studied by Reboulet [Reb22] and by Blum, Liu, Xu and Zhuang [BLXZ21], and these studies have found successful applications in the algebraic theory of $K$-stability and Kähler-Ricci soliton degenerations. For some related constructions, see also [Reb20, Wu22, Fin23].
Our primary motivation is to introduce an analogue of the aforementioned constructions on the space of filtrations on a local ring $(R,\mathfrak{m})$.
Given an $\mathfrak{m}$-filtration, following Ein, Lazarsfeld, and Smith [ELS03], the multiplicity of $\mathfrak{a}_{\bullet}$ is defined to be
$\mathrm{vol}(\mathfrak{a}_{\bullet})\coloneqq\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\ell(R/\mathfrak{a}_{m})}{m^{n}/n!}=\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}},$
where the existence of the above limit and the equality were proven in increasing generality by [ELS03, Mus02, LM09, Cut13, Cut14]. This invariant is the local counterpart of the volume of a graded linear series of a line bundle. Denote the space of all $\mathfrak{m}$-filtrations on $R$ with positive multiplicity by $\mathrm{Fil}_{R,\mathfrak{m}}$, and define a function $d_{1}:\mathrm{Fil}_{R,\mathfrak{m}}\times\mathrm{Fil}_{R,\mathfrak{m}}\to\mathbb{R}_{\geq 0}$ by
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\coloneqq 2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}),$ (1.1)
where $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$ is defined termwise. (1.1) is a similar expression to the Darvas metrics in the global case [Dar15, BJ21].
Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}_{R,\mathfrak{m}}$, the geodesic $\mathfrak{a}_{\bullet,t}$ between them, introduced in [XZ21, BLQ24], is a segment of $\mathfrak{m}$-filtrations for $t\in[0,1]$ defined by
$\mathfrak{a}_{\lambda,t}\coloneqq\sum_{\mu,\nu\geq 0,\ (1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}.$ (1.2)
The definition is a local analogue of the geodesics studied in [BLXZ21].
Our first main result is that the constructions above endow the Hausdorff quotient of $(\mathrm{Fil}_{R,\mathfrak{m}},d_{1})$ with the structure of a geodesic metric space.
Theorem 1.1.
Let $(R,\mathfrak{m})$ be a Noetherian local domain that is analytically irreducible. Denote the set of all linearly bounded $\mathfrak{m}$-filtrations on $R$ by $\mathrm{Fil}_{R,\mathfrak{m}}$. Then
  1. (1)
    the function $d_{1}$ given by (1.1) is a pesudometric on $\mathrm{Fil}_{R,\mathfrak{m}}$,
  2. (2)
    the Hausdorff quotient of $(\mathrm{Fil}_{R,\mathfrak{m}},d_{1})$ is $(\mathrm{Fil}^{s}_{R,\mathfrak{m}},d_{1})$, and
  3. (3)
    $(\mathrm{Fil}^{s}_{R,\mathfrak{m}},d_{1})$ is a convex metric space. Moreover, if $R$ contains a field, then it is a geodesic metric space, and a geodesic between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ can be given by $\mathfrak{a}_{\bullet,t}$ for $t\in[0,1]$, where $\mathfrak{a}_{\bullet,t}$ is defined by (1.2).
Here $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is the subset of $\mathrm{Fil}_{R,\mathfrak{m}}$ consisting of saturated $\mathfrak{m}$-filtrations. For the definition of saturated filtrations, see section 2.2. For definitions related to metric spaces, see section 2.6.
1.1.2. The toric case
In the case where $(R,\mathfrak{m})$ is the local ring of a toric singularity, we can concretely describe the subspace $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ of saturated monomial filtrations.
Let $\mathbbm{k}$ be a field, $N$ a free abelian group of rank $n\geq 1$ and $M=N^{*}$ its dual. Let $\sigma\subset N_{\mathbb{R}}=N\otimes_{\mathbb{Z}}\mathbb{R}$ be a strongly convex rational polyhedral cone of maximal dimension. This gives us an affine normal toric variety $X_{\sigma}=\mathrm{Spec}R_{\sigma}=\mathrm{Spec}\mathbbm{k}[\sigma^{\vee}\cap M]$ with a unique torus invariant point $x$, where $\sigma^{\vee}\subset M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$. Let $R$ be the local ring of $X$ at $x$ and $\mathfrak{m}$ its maximal ideal.
Let $\mathrm{Fil}^{s\mathrm{mon}}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ be the subset of all monomial filtrations (i.e. filtrations $\mathfrak{a}_{\bullet}$ such that every $\mathfrak{a}_{\lambda}$ is monomial). Via the Okounkov body construction, we identify the metric space $(\mathrm{Fil}^{S,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ with a subspace of the space of co-bounded subsets of $\sigma^{\vee}$ with the symmetric difference metric (sometimes also called the Fréchet-Nykodym-Aronszyan metric).
Theorem 1.2.
If $(R,\mathfrak{m})$ is the local ring of a toric singularity $x\in X=\mathrm{Spec}R_{\sigma}$. Then
  1. (1)
    Taking Newton-Okounkov body gives an isometry $P:(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})\to(\mathcal{P}(\sigma^{\vee}),d)$, and
  2. (2)
    the metric space $(\mathrm{Fil}^{s,\mathrm{mon}}_{X,x},d_{1})$ is complete.
In the above theorem, $\mathcal{P}(\sigma^{\vee})$ denotes the set of all closed convex subsets of $\sigma^{\vee}$ with bounded complement, and the metric $d$ is defined to be the Euclidean volume of the symmetric difference of two subsets. For more details, we refer to Section 3.4. Note that $P\mapsto\chi_{P}$ gives an embedding $\mathcal{P}(\sigma^{\vee})\to L^{1}_{\mathrm{loc}}(\sigma^{\vee})$.
1.1.3. The supnorm metric and homogeneous norms
As in the global case [BJ21], there is also a correspondence between filtrations and norms in the local setting, see Definition 2.7 for the definition of norms and Definition-Lemma 2.8, Lemma 2.12 and Lemma 2.29 for the dictionary. Despite the equivalence, it is sometimes more convenient to use one of the languages.
As an example, we now consider another pseudometric $d_{\infty}$ on the space $\mathcal{N}_{R,\mathfrak{m}}$ of $\mathfrak{m}$-norms on $R$, which behaves like the uniform metric on the space of continuous functions on a compact space. This can also be viewed as the local analogue of the $d_{\infty}$ metric introduced in [BJ21].
Fix a homogeneous norm $\rho\in\mathcal{N}^{h}_{R,\mathfrak{m}}$, define
$d_{\infty,\rho}(\chi,\chi^{\prime})\coloneqq\limsup_{\lambda\to\infty}\sup_{\rho(f)\geq\lambda}\frac{|\chi(f)-\chi^{\prime}(f)|}{\rho(f)}.$ (1.3)
We will denote the function $d_{\infty,\chi_{0}}$ by $d_{\infty}$.
Theorem 1.3.
Let $(R,\mathfrak{m})$ be a Noetherian local domain that is analytically irreducible. Then
  1. (1)
    the function $d_{\infty}$ is a pseudometric on $\mathcal{N}_{R,\mathfrak{m}}$, and
  2. (2)
    the restriction of $d_{\infty}$ to $\mathcal{N}^{h}_{R,\mathfrak{m}}$ is a metric, and the metric space $(\mathcal{N}^{h}_{R,\mathfrak{m}},d_{\infty})$ is complete.
Here $\mathcal{N}^{h}_{R,\mathfrak{m}}$ is the subset of $\mathcal{N}_{R,\mathfrak{m}}$ of homogeneous norms. See Section 2.3 for the definition and more details.
1.2. Weak topologies and semicontinuity of log canonical thresholds
Besides the metrics introduced, we also consider the filtrations, or equivalently, norms, as functions on various spaces, and thus define some weak topologies on the spaces of filtrations. We sketch the ideas here. For more details, see Section 2.7.1.
The first perspective is to view a norm $\chi$ as a function on the ring $R$, and thus define the weak topology to be the product topology, that is, the weakest topology such that for any $f\in R$, the function $f\mapsto\chi(f)$ is continuous. Since a real valuation (centered at $\mathfrak{m}$) is a ($\mathfrak{m}$-)norm, this topological space contains the valuation space with the weak topology introduced in [JM12] as a subspace.
Another point of view is to consider a filtration $\mathfrak{a}_{\bullet}$ as a function on certain valuation spaces by $v\mapsto v(\mathfrak{a}_{\bullet})$, where $v(\mathfrak{a}_{\bullet})\coloneqq\lim_{\lambda\to\infty}\frac{v(\mathfrak{a}_{\lambda})}{\lambda}$. We call the topology defined this way using the space $\mathrm{Val}^{+}_{R,\mathfrak{m}}$ of valuations with positive volume the $+$-topology.
We prove that the log canonical threshold satisfies certain semicontinuity on the space of filtrations, and that it is locally Lipschitz continuous with respect to the $d_{\infty}$-topology, which is similar to [DK01, Theorem 3.3].
Theorem 1.4.
Let $(X,x)=(\mathrm{Spec}R,\mathfrak{m})$ be a klt singularity over a field $\mathbbm{k}$ of characteristic 00. Then
  1. (1)
    If $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}$ converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$, then
    $\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\liminf_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}),$
  2. (2)
    if $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ converges to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the $+$-topology, then
    $\mathrm{lct}(\mathfrak{a}_{\bullet})\geq\limsup_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}),$
    and
  3. (3)
    given $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$, for any $\epsilon>0$, there exists $\delta:=\delta(\mathfrak{a}_{\bullet},\epsilon)>0$ such that for any $\mathfrak{b}_{\bullet}$ with $d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})<\delta$, we have
    $|\mathrm{lct}(\mathfrak{b}_{\bullet})-\mathrm{lct}(\mathfrak{a}_{\bullet})|\leq\epsilon.$
1.3. Miscellaneous results about filtrations
A classical result in ring theory is that the semiring of all ideals of a ring $R$ with the partial order by inclusion forms a modular lattice, where meet is given by intersection and join is given by sum. We denote the lattice of all ideals (all $\mathfrak{m}$-primary ideals when $(R,\mathfrak{m})$ is a local ring, resp.) by $\mathcal{I}_{R}$ ($\mathcal{I}_{R,\mathfrak{m}}$, resp.).
The spaces of filtrations are equipped with a natural partial order by inclusion, that is, $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ if and only if $\mathfrak{a}_{\lambda}\subset\mathfrak{b}_{\lambda}$ for any $\lambda\in\mathbb{R}_{>0}$. One can similarly define the intersection of two filtration termwisely, and in Definition-Lemma 2.3 (Definition-Lemma 2.22, resp.), we define the join $\vee$ (saturated join $\vee_{s}$, resp.), which is the analogue of the sum of two ideals. The next theorem asserts that these operations define natural lattice structures on the spaces of $\mathfrak{m}$-filtrations.
Theorem 1.5.
Let $(R,\mathfrak{m})$ be a Noetherian local domain that is analytically irreducible. Then
  1. (1)
    the set of all $\mathfrak{m}$-filtrations with the partial order by inclusion is a lattice, where the meet is given by $\cap$ and the join is given by $\vee$, and there is an injective join morphism from $\mathcal{I}_{R,\mathfrak{m}}$ to its sub-lattice $(\mathrm{Fil}_{R,\mathfrak{m}},\subset,\cap,\vee)$,
  2. (2)
    $(\mathrm{Fil}^{s}_{R,\mathfrak{m}},\subset,\cap,\vee_{s})$ is a distributive lattice. Moreover, $\mathfrak{a}\mapsto\widetilde{\mathfrak{a}^{\bullet}}$ is an injective join morphism from $\mathcal{I}^{\mathrm{ic}}_{R,\mathfrak{m}}$, the set of integrally closed $\mathfrak{m}$-primary ideals, and the saturation $\mathrm{Fil}_{R,\mathfrak{m}}\to\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is a surjective join morphism.
This theorem, together with Theorem 1.1, can be viewed as a higher-dimensional analogue of the results of [FJ04].
We will frequently use the following characterization for saturated filtrations, which can be viewed as an analogue of the Berkovich Maximum Modulus Principle. See also Lemma 2.32.
Proposition 1.6 (= Proposition 2.20).
A filtration $\mathfrak{a}_{\bullet}\in\mathrm{Fil}_{R,\mathfrak{m}}$ is saturated if and only if there exists a (non-empty) subset $\Sigma\subset\mathrm{DivVal}_{R,\mathfrak{m}}$ such that
$\mathfrak{a}_{\bullet}=\cap_{v\in\Sigma}\mathfrak{a}_{\bullet}(v).$
As an example of the applications, when $\mathfrak{a}$ is an $\mathfrak{m}$-primary ideal, we prove that the saturation of the $\mathfrak{a}$-adic filtration is determined by the set of Rees valuations of $\mathfrak{a}$, as expected.
Proposition 1.7 (= Proposition 2.27).
Let $\mathfrak{a}\in\mathcal{I}_{R,\mathfrak{m}}$ be an $\mathfrak{m}$-primary ideal. Then
$\widetilde{\mathfrak{a}^{\bullet}}=\cap_{v\in\mathrm{RV}(\mathfrak{a})}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a})}),$
where $\mathrm{RV}(\mathfrak{a})$ is the set of Rees valuations of $\mathfrak{a}$.
We refer to [HS06, Chapter 10] for more details about Rees valuations.
Organization of the paper. In Section 2, we recall some definitions and basic properties that will be needed. Along the way we also prove some basic facts about filtrations. In Section 3, we introduce the Darvas metric $d_{1}$ and consider its properties. In Section 4, we define the metric(s) $d_{\infty}$, and compare the different topologies on the space of filtrations. In Section 5 we prove the main results Theorem 1.1 through Theorem 1.5. In Section 6, we discuss the relation of the results in this paper with results in the global case, give several further examples, and propose some related questions.
Notation. Throughout the paper, $(R,\mathfrak{m},\kappa)$ denotes an $n$-dimensional Noetherian local domain which is analytically irreducible, that is, the $\mathfrak{m}$-adic completion $\hat{R}$ is again a domain. $\kappa:=R/\mathfrak{m}$ denotes the residue field of $R$. We will always denote the Krull dimension of $R$ by $n$. When $R$ contains a field $\mathbbm{k}$, we sometimes write $X=\mathrm{Spec}R$ and $x\in X$ the closed point corresponding to $\mathfrak{m}$. In this case, we call $x\in X$ a singularity over the field $\mathbbm{k}$.
We write $c\coloneqq c(a,\alpha,\ldots)$ to mean that $c\in\mathbb{R}_{>0}$ is a constant depending only on $a,\alpha\ldots$.
Acknowledgement. A large portion of this paper is derived from the author’s doctoral thesis. The author would like to thank his advisor, Professor Chenyang Xu, for his constant support, encouragement and numerous inspiring conversations. Part of this work is inspired by an earlier collaboration with Harold Blum and Yuchen Liu. The author owes them special thanks for their insights on the topic.
The author would like to thank Zhiyuan Chen, Colin Fan, Yujie Luo, Tomasso de Fernex, Rémi Reboulet, Xiaowei Wang and Ziquan Zhuang for fruitful discussions. The author would also like to thank Tamás Darvas, Jingjun Han, Jihao Liu, Minghao Miao, Junyao Peng, Linsheng Wang, Lingyao Xie, Qingyuan Xue, Tong Zhang and Junyan Zhao for kind comments. Special thanks are also due to Rémi Reboulet and Mattias Jonsson for reviewing a preliminary version of this paper and offering valuable suggestions. Part of this work is done while the author visited the Northwestern University, the University of Utah, Fudan University and BICMR at Peking University. The author would like to thank their hospitality and the amazing environment they provided.
2. Preliminaries
In this section, we recall definitions and basic properties that will be needed in the paper. Several results in Section 2.2, including a characterization of saturated filtrations Propositon 2.20, the lattice structure Definition-Lemma 2.22 and the distributivity Proposition 2.24, seem to be new.
2.1. Filtrations and norms
2.1.1. Filtrations
Definition 2.1.
A filtration on $R$ is a collection $\mathfrak{a}_{\bullet}=(\mathfrak{a}_{\lambda})_{\lambda\in\mathbb{R}_{>0}}$ of ideals of $R$ such that
  1. (1)
    (decreasing) $\mathfrak{a}_{\lambda}\subset\mathfrak{a}_{\mu}$ for any $\lambda>\mu$,
  2. (2)
    (left-continuouity) $\mathfrak{a}_{\lambda}=\mathfrak{a}_{\lambda-\epsilon}$ for any $\lambda\in\mathbb{R}_{>0}$ and $0<\epsilon\ll 1$, and
  3. (3)
    (multiplicativity) $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ for any $\lambda,\mu\in\mathbb{R}_{>0}$.
A filtration $\mathfrak{a}_{\bullet}$ is called an $\mathfrak{m}$-filtration if $\mathfrak{a}_{\lambda}$ is $\mathfrak{m}$-primary for any $\lambda\in\mathbb{R}_{>0}$. By convention, we always set $\mathfrak{a}_{0}\coloneqq R$.
The definition is a local analogue of a filtration on the section ring of a polarized variety in [BHJ17]. We will exclusively work with $\mathfrak{m}$-filtrations in the sequel, though some of the results also hold for general filtrations on $R$.
For $\lambda\in\mathbb{R}_{\geq 0}$, set $\mathfrak{a}_{>\lambda}:=\cup_{\mu>\lambda}\mathfrak{a}_{\mu}$. If $\lambda\in\mathbb{R}_{>0}$ satisfies $\mathfrak{a}_{>\lambda}\subsetneq\mathfrak{a}_{\lambda}$, then we call $\lambda$ a jumping number of $\mathfrak{a}_{\bullet}$. The scaling of an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ by $c\in\mathbb{R}_{>0}$ is $\mathfrak{a}_{c\bullet}:=(\mathfrak{a}_{c\lambda})_{\lambda\in\mathbb{R}_{>0}}$, which is an $\mathfrak{m}$-filtration. For two $\mathfrak{m}$-filtrations $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$, we say that $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ if $\mathfrak{a}_{\lambda}\subset\mathfrak{b}_{\lambda}$ for any $\lambda\in\mathbb{R}_{>0}$. This defines a partial order on the set of all $\mathfrak{m}$-filtrations. There is a maximal element among all $\mathfrak{m}$-filtrations, $\{\mathfrak{m}\}_{\lambda\in\mathbb{R}_{>0}}$. We will sometimes abuse the notation and denote this $\mathfrak{m}$-filtration by $\mathfrak{m}$.
Let $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ be two $\mathfrak{m}$-filtrations. Their intersection is $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}:=(\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})_{\lambda\in\mathbb{R}_{>0}}$. It is not hard to check that this is again an $\mathfrak{m}$-filtrations.
Definition-Lemma 2.2.
Let $\{\mathfrak{a}_{\bullet,i}\}_{i\in I}$ be an arbitrary non-empty set of $\mathfrak{m}$-filtrations. Assume that there exists an $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}$ such that $\mathfrak{b}_{\bullet}\subset\mathfrak{a}_{\bullet,i}$ for any $i\in I$. Then the collection of $\mathfrak{m}$-primary ideals, $(\cap_{i\in I}\mathfrak{a}_{\lambda,i})_{\lambda\in\mathbb{R}_{>0}}$, is an $\mathfrak{m}$-filtration, called the intersection of the $\mathfrak{a}_{\bullet,i}$, and is denoted by $\cap_{i\in I}\mathfrak{a}_{\bullet,i}$. Moreover, if an $\mathfrak{m}$-filtration $\mathfrak{c}_{\bullet}$ satisfies $\mathfrak{c}_{\bullet}\subset\mathfrak{a}_{\bullet,i}$ for any $i\in I$, then we have $\mathfrak{c}_{\bullet}\subset\cap_{i\in I}\mathfrak{a}_{\bullet,i}$.
Proof.
By assumption, for any $\lambda>0$ we have $\mathfrak{b}_{\lambda}\subset\mathfrak{a}_{\lambda,i}$ and hence $\cap_{i\in I}\mathfrak{a}_{\lambda,i}$ is $\mathfrak{m}$-primary. It is straightforward to verify conditions (1) and (3) in the definition.
To prove (2), note that $\mathfrak{b}_{\lambda}\subset\cap_{i}\mathfrak{a}_{\lambda,i}\subset\cap_{i}\mathfrak{a}_{\mu,i}$ for any $\mu\leq\lambda$. In $R/\mathfrak{b}_{\lambda}$ we have
$\cap_{i\in I}\mathfrak{a}_{\lambda,i}=$ $\cap_{i\in I}\cap_{\epsilon>0}\mathfrak{a}_{\lambda-\epsilon,i}=\cap_{\epsilon>0}\cap_{i\in I}\mathfrak{a}_{\lambda-\epsilon,i}$
$=$ $\cap_{l\in\mathbb{Z}_{>1/\lambda}}\cap_{i\in I}\mathfrak{a}_{\lambda-\frac{1}{l},i}=\cap_{i\in I}\mathfrak{a}_{\lambda-\frac{1}{l},i}$
for some $l\in\mathbb{Z}_{>1/\lambda}$, where the first equality follows from the left-continuouity of $\mathfrak{a}_{\bullet,i}$, and the last equality uses the fact that $R/\mathfrak{b}_{\lambda}$ is Artinian. This proves that condition (2) holds for $\{\cap_{i\in I}\mathfrak{a}_{\lambda,i}\}$, and thus it is an $\mathfrak{m}$-filtration.
The last assertion follows by definition, and the proof is finished. ∎
Definition-Lemma 2.3.
Let $\{\mathfrak{a}_{\bullet,i}\}_{i\in I}$ be an arbitrary set of $\mathfrak{m}$-filtrations. Then there is a filtration, called the join of the $\mathfrak{a}_{\bullet,i}$, denoted by $\vee_{i\in I}\mathfrak{a}_{\bullet,i}$, such that $\mathfrak{a}_{\bullet,i}\subset\vee_{i\in I}\mathfrak{a}_{\bullet,i}$ for any $i\in I$. Moreover, if an $\mathfrak{m}$-filtration $\mathfrak{c}_{\bullet}$ satisfies $\mathfrak{a}_{\bullet,i}\subset\mathfrak{c}_{\bullet}$ for any $i\in I$, then we have $\vee_{i\in I}\mathfrak{a}_{\bullet,i}\subset\mathfrak{c}_{\bullet}$.
Proof.
If $I=\emptyset$, then $\vee_{i\in I}\mathfrak{a}_{\bullet,i}$ is the maximal $\mathfrak{m}$-filtration $\mathfrak{m}$.
Assume now $I\neq\emptyset$ and fix $0\in I$. The set
$J:=\{\mathfrak{b}_{\bullet}\mid\mathfrak{a}_{\bullet,i}\subset\mathfrak{b}_{\bullet}\text{ for any }i\in I\}$
is non-empty, since the maximal $\mathfrak{m}$-filtration $\mathfrak{m}\in J$. Moreover, by definition $\mathfrak{a}_{\bullet,0}\subset\mathfrak{b}_{\bullet}$ for any $\mathfrak{b}_{\bullet}\in J$. Hence we can apply Definition-Lemma 2.2 to get an $\mathfrak{m}$-filtration
$\vee_{i\in I}\mathfrak{a}_{\bullet,i}\coloneqq\cap_{\mathfrak{b}_{\bullet}\in J}\mathfrak{b}_{\bullet}.$
The last assertion follows by definition, and the proof is finished. ∎
Example 2.4.
Given an $\mathfrak{m}$-primary ideal $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$, we have an $\mathfrak{m}$-filtration $\mathfrak{a}^{\bullet}$, the $\mathfrak{a}$-adic filtration, defined by $\{\mathfrak{a}^{\lceil\lambda\rceil}\}_{\lambda\in\mathbb{R}_{>0}}$. Thus we get an injection from the set $\mathcal{I}_{\mathfrak{m}}$ to the set of all $\mathfrak{m}$-filtrations by $\mathfrak{a}\mapsto\mathfrak{a}^{\bullet}$ In particular, there is a canonical $\mathfrak{m}$-filtration $\mathfrak{m}^{\bullet}$.
For $\mathfrak{a},\mathfrak{b}\in\mathcal{I}_{\mathfrak{m}}$, we claim that $\mathfrak{a}^{\bullet}\vee\mathfrak{b}^{\bullet}=(\mathfrak{a}+\mathfrak{b})^{\bullet}$. Indeed, since $\mathfrak{a}^{\bullet}\subset(\mathfrak{a}+\mathfrak{b})^{\bullet}$ and $\mathfrak{b}^{\bullet}\subset(\mathfrak{a}+\mathfrak{b})^{\bullet}$, by Definition-Lemma 2.3 we have $\mathfrak{c}_{\bullet}\coloneqq\mathfrak{a}^{\bullet}\vee\mathfrak{b}^{\bullet}\subset(\mathfrak{a}+\mathfrak{b})^{\bullet}$. On the other hand, we have $\mathfrak{a}\subset\mathfrak{c}_{1}$ and $\mathfrak{b}\subset\mathfrak{c}_{1}$ by definition. Hence for any $\lambda\in\mathbb{R}_{>0}$, we have
$(\mathfrak{a}+\mathfrak{b})^{\lceil\lambda\rceil}\subset\mathfrak{c}_{1}^{\lceil\lambda\rceil}\subset\mathfrak{c}_{\lceil\lambda\rceil}\subset\mathfrak{c}_{\lambda}.$
This proves $(\mathfrak{a}+\mathfrak{b})^{\bullet}\subset\mathfrak{c}_{\bullet}$ and thus the equality holds.
However, in general we only have $(\mathfrak{a}\cap\mathfrak{b})^{\bullet}\subset\mathfrak{a}^{\bullet}\cap\mathfrak{b}^{\bullet}$ and the inclusion can be strict. Indeed, let $R=\mathbbm{k}[\![x,y]\!]$ for some field $\mathbbm{k}$, $\mathfrak{a}=(x,y^{2})$ and $\mathfrak{b}=(x^{2},y)$. Then it is not hard to see that $\mathfrak{a}\cap\mathfrak{b}=(x^{2},xy,y^{2})=\mathfrak{m}^{2}$ and $(\mathfrak{a}\cap\mathfrak{b})^{2}=\mathfrak{m}^{4}$. Hence $x^{2}y\in\mathfrak{a}^{2}\cap\mathfrak{b}^{2}\backslash(\mathfrak{a}\cap\mathfrak{b})^{2}$.
Unlike the intersection, it can be hard to determine the join of two general filtrations. In Section 2.2, we will define a similar notion, the saturated join, and give a formula for it. The following proposition, which is a combination of Definition-Lemma 2.2 and Defnition-Lemma 2.3, justifies the term join.
Proposition 2.5.
The set of all $\mathfrak{m}$-filtrations with the partial order by inclusion form a lattice, where the meet of two elements is given by their intersection, and the join is defined as in Definition-Lemma 2.3. ∎
Remark 2.6.
Note that by Example 2.4, the lattice $\mathcal{I}_{\mathfrak{m}}$ of $\mathfrak{m}$-primary ideals of $R$ does not embed in the lattice of all $\mathfrak{m}$-filtrations, since the meet is not preserved.
2.1.2. Norms and valuations
We recall the correspondence between filtrations and norms on a local ring. See also [HS06, BFJ14] for some related results.
Definition 2.7.
A seminorm on $R$ is a function $\chi:R\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ such that
  1. (1)
    $\chi(0)=+\infty$ and $\chi(-f)=\chi(f)$,
  2. (2)
    $\chi(f+g)\geq\min\{\chi(f),\chi(g)\}$ for any $f,g\in R$, and
  3. (3)
    (Sub-multiplicativity) $\chi(fg)\geq\chi(f)+\chi(g)$ for any $f,g\in R$.
A seminorm $\chi$ is a norm if $\chi(f)=+\infty$ if and only if $f=0$. A (semi)norm $\chi$ is called an $\mathfrak{m}$-(semi)norm if it further satisfies
  1. (4)
    $\chi(f)\geq 0$ for any $f\in R$ and $\chi(f)>0$ if and only if $f\in\mathfrak{m}$.
If $\chi$ is a ($\mathfrak{m}$-)norm on $R$, and we replace condition (3) above by
  1. (3)
    (Multiplicativity) $\chi(fg)=\chi(f)+\chi(g)$ for any $f,g\in R$,
then $\chi$ is a real valuation on $R$ (centered at $\mathfrak{m}$). Denote the space of real valuations on $R$ centered at $\mathfrak{m}$ by $\mathrm{Val}_{R,\mathfrak{m}}$.
If $R$ is a $\mathbbm{k}$-algebra for a field $\mathbbm{k}$, then we will always assume that $\chi$ is a $\mathbbm{k}$-norm, that is, $\chi(af)=\chi(f)$ for any $a\in\mathbbm{k}^{\times}$ and $f\in R$.
As in the global case, we adopt the additive terminology here. For a norm $\chi$ on $R$, $\|\cdot\|_{\chi}:=e^{-\chi(\cdot)}$ defines a non-Archimedean norm on $R$ in the sense of [BGR84].
There is a correspondence between seminorms and filtrations, given by the lemma below.
Definition-Lemma 2.8.
Given an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$, its order function $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is defined by
$\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f):=\sup\{\lambda\in\mathbb{R}_{>0}\mid f\in\mathfrak{a}_{\lambda}\},$ (2.1)
which is a seminorm on $R$. If $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda\ll 1$, then it is an $\mathfrak{m}$-seminorm.
Conversely, given an $\mathfrak{m}$-seminorm $\chi$, define $\mathfrak{a}_{\bullet}(\chi)$ by
$\mathfrak{a}_{\lambda}(\chi):=\{f\in R\mid\chi(f)\geq\lambda\}.$ (2.2)
Then $\mathfrak{a}_{\bullet}(\chi)$ is an $\mathfrak{m}$-filtration, called the associated filtration of $\chi$.
The above maps give a 1-1 correspondence between $\mathfrak{m}$-seminorms and $\mathfrak{m}$-filtrations satisfying $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda\ll 1$. Moreover, $\mathfrak{m}$-norms correspond to $\mathfrak{m}$-filtrations with $\cap_{\lambda>0}\mathfrak{a}_{\lambda}=\{0\}$.
The associated filtration of a real valuation $v\in\mathrm{Val}_{R,\mathfrak{m}}$ is also called the valuation ideals of $v$.
Proof.
Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration. Clearly $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ satisfies condition (1) of Definition 2.7. Condition (2) follows from the fact that each $\mathfrak{a}_{\lambda}$ is an ideal. Condition (3) follows from the multiplicativity of a filtration. If $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda\ll 1$, then condition (4) follows from the convention $\mathfrak{a}_{0}=R$.
Conversely, let $\chi$ be an $\mathfrak{m}$-seminorm. We first show that for any $\lambda>0$, $\mathfrak{a}_{\lambda}(\chi)$ defined in (2.2) is an $\mathfrak{m}$-primary ideal. It is an ideal since $\chi(f-g)\geq\min\{\chi(f),\chi(-g)\}=\min\{\chi(f),\chi(g)\}$ and $\chi(fg)\geq\chi(f)+\chi(g)\geq\min\{\chi(f),\chi(g)\}$. Choose a set of generators $f_{1},\ldots,f_{r}$ of $\mathfrak{m}$, then for any $f\in\mathfrak{m}$ we have $\chi(f)\geq\chi(\mathfrak{m})\coloneqq\min_{1\leq i\leq r}\{\chi(f_{i})\}>0$. Let $C\coloneqq\lceil\lambda/\chi(\mathfrak{m})\rceil$, then $\chi(f^{C})\geq C\chi(f)\geq\lambda$, that is, $\mathfrak{m}^{C}\subset\mathfrak{a}_{\lambda}(\chi)$. Note that the above argument also shows that $\mathfrak{a}_{\lambda}(\chi)=\mathfrak{m}$ for $0<\lambda<\chi(\mathfrak{m})$. Clearly the collection of $\mathfrak{m}$-primary ideals $\{\mathfrak{a}_{\lambda}(\chi)\}$ is decreasing, and it is multiplicative by condition (3) of Definition 2.7. To show left continuity, note that we can write $\mathfrak{a}_{\lambda}(\chi)$ as an decreasing intersection
$\mathfrak{a}_{\lambda}(\chi)=\cap_{\epsilon>0}\mathfrak{a}_{\lambda-\epsilon}(\chi),$
which terminates since the equality holds in the Artinian ring $R/\mathfrak{a}_{\lambda}(\chi)$. ∎
Note that the supremum in the definition is indeed a maximum if the value is finite.
Example 2.9.
The canonical $\mathfrak{m}$-filtration $\mathfrak{m}^{\bullet}=\{\mathfrak{m}^{\lceil\lambda\rceil}\}_{\lambda\in\mathbb{R}_{>0}}$ corresponds to the canonical $\mathfrak{m}$-norm $\mathrm{ord}_{\mathfrak{m}}:R\to\mathbb{Z}_{\geq 0}\cup\{+\infty\}$ given by
$\mathrm{ord}_{\mathfrak{m}}(f):=\sup\{k\in\mathbb{Z}_{\geq 0}\mid f\in\mathfrak{m}^{k}\},$
which has been considered in [BFJ14, Section 4].
Remark 2.10.
By Lemma 2.29, a homogeneous filtration satisfies the additional condition above. In particular, we may identify the spaces of filtrations and norms whenever we consider the metric properties.
Indeed, since we only consider the asymptotic behavior of a filtration, we can modify the terms of $\mathfrak{a}_{\bullet}$ in the following way, such that the additional condition is always satisfied. Since $\{\mathfrak{a}_{1/n}\}$ is an increasing sequence of $\mathfrak{m}$-primary ideals, there exists $N\in\mathbb{Z}_{>0}$ such that $\mathfrak{a}_{1/n}=\mathfrak{a}_{1/N}\eqqcolon\mathfrak{a}$ for any $n\geq N$. If $\mathfrak{a}=\mathfrak{m}$ then we are done. Otherwise, choose $C\in\mathbb{Z}_{>0}$ such that $\mathfrak{m}^{C}\subset\mathfrak{a}$ and define
$\mathfrak{b}_{\lambda}\coloneqq\left\{\begin{aligned} &\mathfrak{m},&0<\lambda\leq 1/NC,\\ &\mathfrak{a}_{\lambda},&\text{otherwise.}\end{aligned}\right.$
Then clearly $\mathfrak{b}_{\bullet}$ has the same asymptotic information as $\mathfrak{a}_{\bullet}$.
2.1.3. Divisorial valuations
A valuation $v\in\mathrm{Val}_{R,\mathfrak{m}}$ is divisorial if
${\rm tr.deg}_{\kappa}(\kappa_{v})=n-1.$
We write $\mathrm{DivVal}_{R,\mathfrak{m}}\subset\mathrm{Val}_{R,\mathfrak{m}}$ for the set of such valuations.
Divisorial valuations appear geometrically. If $\mu:Y\to X$ is a proper birational morphism with $Y$ normal and $E\subset Y$ a prime divisor, then there is an induced valuation $\mathrm{ord}_{E}\colon{\rm Frac}(R)^{\times}\to\mathbb{Z}$. If $\mu(E)=x$ and $c\in\mathbb{R}_{>0}$, then $c\cdot\mathrm{ord}_{E}\in\mathrm{DivVal}_{R,\mathfrak{m}}$. When $R$ is excellent, all divisorial valuations are of this form; see e.g. [CS22, Lemma 6.5].
2.1.4. Quasi-monomial valuations
In the following construction, we always assume $R$ contains a field. Let $\mu\colon Y:=\mathrm{Spec}(S)\to X=\mathrm{Spec}(R)$ be a birational morphism with $R\to S$ finite type and $\eta\in Y$ a not necessarily closed point such that $\mathcal{O}_{Y,\eta}$ is regular and $\mu(\eta)=x$. Given a regular system of parameters $y_{1},\ldots,y_{r}$ of $\mathcal{O}_{Y,\eta}$ and $\bm{\alpha}=(\alpha_{1},\cdots,\alpha_{r})\in\mathbb{R}_{\geq 0}^{r}\setminus{\bf 0}$, we define a valuation $v_{\bf\alpha}$ as follows. For $0\neq f\in\mathcal{O}_{Y,\eta}$, we can write $f$ in $\widehat{\mathcal{O}}_{Y,\eta}\simeq k(\eta)[[y_{1},\ldots,y_{r}]]$ as $\sum_{\bm{\beta}\in\mathbb{Z}_{\geq 0}^{r}}c_{\bm{\beta}}{y}^{\bm{\beta}}$ and set
$v_{\bm{\alpha}}(f)\coloneqq\min\{\langle\bm{\alpha},{\bm{\beta}}\rangle\,|\,c_{\bm{\beta}}\neq 0\}.$
A valuation of the above form is called quasi-monomial.
Let $D=D_{1}+\cdots+D_{r}$ be a reduced divisor on $Y$ such that $y_{i}=0$ locally defines $D_{i}$ and $\mu(D_{i})=x$ for each $i$. We call $\eta\in(Y,D)$ a log smooth birational model of $X$. We write $\mathrm{QM}_{\eta}(Y,E)\subset\mathrm{Val}_{X,x}$ for the set of quasi-monomial valuations that can be described at $\eta$ with respect to $y_{1},\ldots,y_{r}$ and note that $\mathrm{QM}_{\eta}(Y,D)\simeq\mathbb{R}^{r}_{\geq 0}\setminus\bm{0}$.
2.1.5. Multiplicity of a filtration
Following [ELS03], the multiplicity of a graded sequence of $\mathfrak{m}$-primary ideals $\mathfrak{a}_{\bullet}$ is
$\mathrm{e}(\mathfrak{a}_{\bullet})\coloneqq\lim_{m\to\infty}\frac{\ell(R/\mathfrak{a}_{m})}{m^{n}/n!}\in[0,\infty).$
By [ELS03, Mus02, LM09, Cut13, Cut14] in increasing generality, the above limit exists and
$\mathrm{e}(\mathfrak{a}_{\bullet})=\lim_{m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}}=\inf_{m}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}};$ (2.3)
see in particular [Cut14, Theorem 6.5]. Also defined in [ELS03], the volume of a valuation $v\in\mathrm{Val}_{R,\mathfrak{m}}$ is $\mathrm{vol}(v):=\mathrm{e}(\mathfrak{a}_{\bullet}(v))$. Denote $\mathrm{Val}_{R,\mathfrak{m}}^{+}\coloneqq\{v\in\mathrm{Val}_{R,\mathfrak{m}}\mid\mathrm{vol}(v)>0\}\subset\mathrm{Val}_{R,\mathfrak{m}}$.
2.1.6. Linearly boundedness
Definition 2.11.
An $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ is linearly bounded if there exists a constant $c\in\mathbb{R}_{>0}$ such that $\mathfrak{a}_{\bullet}\subset\mathfrak{m}^{c\bullet}$.
An $\mathfrak{m}$-norm $\chi$ is linearly bounded if there exists $c\in\mathbb{R}_{>0}$ such that $\chi(f)\leq c^{-1}\mathrm{ord}_{\mathfrak{m}}(f)$ for any $f\in R$.
By Krull intersection theorem, a linearly bounded $\mathfrak{m}$-filtration satisfies $\cap_{\lambda\in\mathbb{R}_{\geq 0}}\mathfrak{a}_{\lambda}=\{0\}$. In particular, by Definitin-Lemma 2.8, we have the following
Lemma 2.12.
An $\mathfrak{m}$-filtration is linearly bounded if and only if $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is a linearly bounded $\mathfrak{m}$-norm.
We will consider filtrations with positive multiplicity, which are exactly those which are linearly bounded by the following lemma.
Lemma 2.13.
[BLQ24, Corollary 3.18] An $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ is linearly bounded if and only if $\mathrm{e}(\mathfrak{a}_{\bullet})>0$.
Denote the set of all linearly bounded $\mathfrak{m}$-filtrations on $R$ by $\mathrm{Fil}_{R,\mathfrak{m}}$, and the set of all linearly bounded $\mathfrak{m}$-norms on $R$ by $\mathcal{N}_{R,\mathfrak{m}}$. When there is no ambiguity, we will write $\mathrm{Fil}$ and $\mathcal{N}$ respectively. By Definition-Lemma 2.8 and Lemma 2.12, $\mathcal{N}$ is a subset of $\mathrm{Fil}$.
There is a partial order $\leq$ on $\mathcal{N}$ defined pointwisely on $R$, that is,
$\chi\leq\chi^{\prime}\text{ if and only if }\chi(f)\leq\chi^{\prime}(f)\text{ for any }f\in R,$
which corresponds to the partial order defined by inclusion on filtrations. Moreover, the natural $\mathbb{R}_{>0}$-action of scaling on $\mathcal{N}$ corresponds to the scaling on $\mathfrak{m}$-filtrations is given by
$(c\cdot\chi)(f):=c^{-1}\chi(f).$
2.1.7. Asymptotic invariants
Lemma 2.14.
(cf. [JM12, Lemma 2.4]) Let $\chi\in\mathcal{N}_{R,\mathfrak{m}}$ and $w\in\mathrm{Val}_{R,\mathfrak{m}}$. Then
$w(\mathfrak{a}_{\bullet}(\chi))=\inf_{f\in\mathfrak{m}}\frac{w(f)}{\chi(f)}.$
Proof.
Let $c\coloneqq\inf_{f\in\mathfrak{m}}\frac{w(f)}{\chi(f)}$. Then for any $f\in\mathfrak{m}$ we have $w(f)\geq c\cdot\chi(f)$, hence by definition, $w(\mathfrak{a}_{\lambda}(\chi))\geq c\cdot\chi(\mathfrak{a}_{\lambda}(\chi))\geq c\lambda$ for any $\lambda\in\mathbb{R}_{>0}$. Letting $\lambda\to\infty$ and dividing by $\lambda$, we get $w(\mathfrak{a}_{\bullet}(\chi))\geq c$.
Conversely, for any $\epsilon>0$, by definition there exists $f\in\mathfrak{m}$ such that $0<w(f)<(c+\epsilon)\chi(f)$. Note that for any $k\in\mathbb{Z}_{>0}$, $f^{d}\in\mathfrak{a}_{d\cdot\chi(f)}(\chi)$ as $\chi(f^{d})\geq d\cdot\chi(f)$. Thus for $d\in\mathbb{Z}_{>0}$ we have
$\frac{w(\mathfrak{a}_{d\cdot\chi(f)}(\chi))}{d\cdot\chi(f)}\leq\frac{w(f^{d})}{d\cdot\chi(f)}=\frac{d\cdot w(f)}{d\cdot\chi(f)}<c+\epsilon.$
Letting $d\to\infty$ we get $w(\mathfrak{a}_{\bullet}(\chi))\leq c+\epsilon$. This finishes the proof. ∎
2.2. Saturated filtrations
The saturation of an $\mathfrak{m}$-filtration was introduced in [BLQ24]. We recall some facts about saturations and provide a characterization for saturated filtrations in Proposition 2.20. Several of the results in this section seem to be new.
Definition 2.15.
The saturation $\widetilde{\mathfrak{a}}_{\bullet}$ of an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ is defined by
$\widetilde{\mathfrak{a}}_{\lambda}\coloneqq\{f\in\mathfrak{m}\mid v(f)\geq\lambda\cdot v(\mathfrak{a}_{\bullet})\text{ for all }v\in\mathrm{DivVal}_{R,\mathfrak{m}}\}$
for each $\lambda\in\mathbb{R}_{>0}$. We say that $\mathfrak{a}_{\bullet}$ is saturated if $\mathfrak{a}_{\bullet}=\widetilde{\mathfrak{a}}_{\bullet}$.
Denote the subset of $\mathrm{Fil}$ of saturated filtrations by $\mathrm{Fil}^{s}\coloneqq\{\mathfrak{a}_{\bullet}\in\mathrm{Fil}\mid\mathfrak{a}_{\bullet}\text{ is saturated}\}$. There is a canonical element in $\mathrm{Fil}^{s}$, that is, the saturation $\widetilde{\mathfrak{m}}^{\bullet}$ of the $\mathfrak{m}$-adic filtration. Denote its order function by $\chi_{0}:=\mathrm{ord}_{\widetilde{\mathfrak{m}}^{\bullet}}=\widetilde{\mathrm{ord}}_{\mathfrak{m}}$.
The following lemma asserts that for saturated filtrations, the partial order by inclusion can be detected by evaluating on all divisorial valuations, which is not the case for general filtrations.
Lemma 2.16.
Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration and $\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{s}$. Then $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ if and only if $v(\mathfrak{a}_{\bullet})\geq v(\mathfrak{b}_{\bullet})$ for any $v\in\mathrm{DivVal}_{X,x}$.
Proof.
If $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ then by definition $v(\mathfrak{a}_{\bullet})\geq v(\mathfrak{b}_{\bullet})$ for any $v\in\mathrm{Val}_{X,x}$.
Conversely, assume that $\mathfrak{a}_{\bullet}\nsubseteq\mathfrak{b}_{\bullet}$. Then there exists $\lambda\in\mathbb{R}_{>0}$ and $f\in\mathfrak{a}_{\lambda}$ such that $f\notin\mathfrak{b}_{\lambda}$. This implies that there exists $\epsilon>0$ and $v\in\mathrm{DivVal}_{X,x}$ such that $v(f)<\lambda\cdot v(\mathfrak{b}_{\bullet})-\epsilon$. Hence for any $m\in\mathbb{Z}_{>0}$, we have
$v(\mathfrak{a}_{m\lambda})\leq v(\mathfrak{a}_{\lambda}^{m})=mv(\mathfrak{a}_{\lambda})\leq mv(f)<m(\lambda\cdot v(\mathfrak{b}_{\bullet})-\epsilon),$
where the first inequality follows from $\mathfrak{a}_{\lambda}^{m}\subset\mathfrak{a}_{m\lambda}$. Dividing by $m\lambda$ and letting $m\to\infty$, we get $v(\mathfrak{a}_{\bullet})\leq v(\mathfrak{b}_{\bullet})-\epsilon/\lambda<v(\mathfrak{b}_{\bullet})$. ∎
The saturation can be alternatively defined using all valuations with positive volume.
Lemma 2.17.
[BLQ24, Proposition 3.19] If $\mathfrak{a}_{\bullet}\in\mathrm{Fil}_{R,\mathfrak{m}}$ and $\lambda\in\mathbb{R}_{>0}$, then
$\widetilde{\mathfrak{a}}_{\lambda}=\{f\in\mathfrak{m}\mid v(f)\geq\lambda\cdot v(\mathfrak{a}_{\bullet})\text{ for all }v\in\mathrm{Val}^{+}_{R,\mathfrak{m}}\}.$
One of the main properties of saturations is the following generalization of Rees’ theorem on multiplicities of ideals.
Theorem 2.18.
[BLQ24, Theorem 1.4] For $\mathfrak{m}$-filtrations $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$, $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})$ if and only if $\widetilde{\mathfrak{a}}_{\bullet}=\widetilde{\mathfrak{b}}_{\bullet}$.
In other words, $\widetilde{\mathfrak{a}}_{\bullet}$ is the unique maximal $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}\supset\mathfrak{a}_{\bullet}$ such that $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})$. We now prove a formula for saturations.
Lemma 2.19.
Let $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$. Then we have
$\widetilde{\mathfrak{a}}_{\bullet}=\bigcap_{v(\mathfrak{a}_{\bullet})=1}\mathfrak{a}_{\bullet}(v)=\bigcap_{v\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}}\mathfrak{a}_{\bullet}(v),$
where the first intersection is taken over all $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$ with $v(\mathfrak{a}_{\bullet})=1$ and the second over all $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$ such that $v(f)\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)$ for any $f\in\mathfrak{m}$.
Proof.
Note that $v(\mathfrak{a}_{\bullet})>0$ for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$, by Lemma 2.13. For $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$ with $v(\mathfrak{a}_{\bullet})=1$, by [BLQ24, Proposition 3.9] we know that $v(\widetilde{\mathfrak{a}}_{\bullet})=1$, so $\widetilde{\mathfrak{a}}_{\bullet}\subset\mathfrak{a}_{\bullet}(v)$ by definition. Hence $\widetilde{\mathfrak{a}}_{\bullet}\subset\cap_{v(\mathfrak{a}_{\bullet})=1}\mathfrak{a}_{\bullet}(v)$. Conversely, if $f\in\cap_{v(\mathfrak{a}_{\bullet})=1}\mathfrak{a}_{\lambda}(v)$, then for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$, since $v^{\prime}(\mathfrak{a}_{\bullet})=1$, where $v^{\prime}:=v/v(\mathfrak{a}_{\bullet})$, we have
$v(f)=\frac{v}{v(\mathfrak{a}_{\bullet})}(f)\cdot v(\mathfrak{a}_{\bullet})\geq\lambda\cdot v(\mathfrak{a}_{\bullet}).$
Hence $\cap_{v(\mathfrak{a}_{\bullet})=1}\mathfrak{a}_{\lambda}(v)\subset\widetilde{\mathfrak{a}}_{\bullet}$ and this proves the first equality.
If $v(\mathfrak{a}_{\bullet})=1$ then for any $f\in\mathfrak{a}_{m}$, we have $v(f)\geq v(\mathfrak{a}_{m})\geq mv(\mathfrak{a}_{\bullet})$. Hence $v(f)\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)$, that is, $v\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}$, and we get $\cap_{v(\mathfrak{a}_{\bullet})=1}\mathfrak{a}_{m}(v)\supset\cap_{v\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}}\mathfrak{a}_{m}(v)$. Conversely, if $v\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}$, then $v(\mathfrak{a}_{m})\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}(\mathfrak{a}_{m})\geq m$, which implies $v(\mathfrak{a}_{\bullet})\geq 1$, that is, $v\geq v^{\prime}\coloneqq v/v(\mathfrak{a}_{\bullet})$, where $v^{\prime}(\mathfrak{a}_{\bullet})=1$. Thus we get
$\cap_{v(\mathfrak{a}_{\bullet})=1}\mathfrak{a}_{m}(v)\subset\cap_{v^{\prime}}\mathfrak{a}_{m}(v^{\prime})\subset\cap_{v\geq\mathrm{ord}_{\mathfrak{a}_{\bullet}}}\mathfrak{a}_{m}(v).$
This proves the second equality. ∎
Based on the observation above, we give a characterization for saturated filtrations.
Proposition 2.20.
A filtration $\mathfrak{a}\in\mathrm{Fil}$ is saturated if and only if it is of the form
$\mathfrak{a}_{\bullet}=\cap_{v\in\Sigma}\mathfrak{a}_{\bullet}(v),$
where $\Sigma$ is a nonempty subset of either $\mathrm{DivVal}_{R,\mathfrak{m}}$ or $\mathrm{Val}^{+}_{R,\mathfrak{m}}$.
Proof.
If $\mathfrak{a}_{\bullet}=\cap_{v\in\Sigma}\mathfrak{a}_{\bullet}(v)$ for some $\Sigma\subset\mathrm{Val}^{+}_{R,\mathfrak{m}}$, then by [BLQ24, Lemma 3.20], $\mathfrak{a}_{\bullet}$ is saturated.
Conversely, if $\mathfrak{a}_{\bullet}$ is saturated, then by Lemma 2.19, $\mathfrak{a}_{\bullet}=\widetilde{\mathfrak{a}}_{\bullet}=\cap_{v\in\Sigma_{v}}\mathfrak{a}_{\bullet}(v)$, where
$\Sigma_{v}\coloneqq\{v\in\mathrm{DivVal}_{R,\mathfrak{m}}\mid v(\mathfrak{a}_{\bullet})=1)\}\subset\mathrm{DivVal}_{R,\mathfrak{m}}.$
Since $\mathrm{DivVal}_{R,\mathfrak{m}}\subset\mathrm{Val}^{+}_{R,\mathfrak{m}}$, the proof is finished. ∎
Corollary 2.21.
Let $\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}^{s}$ for $i\in I$. If $\cap_{i\in I}\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}$, then $\cap_{i\in I}\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}^{s}$.
Proof.
By Proposition 2.20, for any $i\in I$ we may write $\mathfrak{a}_{\bullet,i}=\cap_{v\in\Sigma_{i}}\mathfrak{a}_{\bullet}(v)$ for $\Sigma_{i}\subset\mathrm{DivVal}_{X,x}$. Thus
$\cap_{i\in I}\mathfrak{a}_{\bullet,i}=\cap_{v\in\cup_{i\in I}\Sigma_{i}}\mathfrak{a}_{\bullet}(v)\in\mathrm{Fil}^{s}$
by Proposition 2.20 again. ∎
Definition-Lemma 2.22.
Let $\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}$ for $i\in I\neq\emptyset$. If there exists $\mathfrak{b}_{\bullet}\in\mathrm{Fil}$ such that $\mathfrak{a}_{\bullet,i}\subset\mathfrak{b}_{\bullet}$ for any $i\in I$, then there exists an $\mathfrak{m}$-filtration, called the saturated join of the $\mathfrak{a}_{\bullet,i}$, denoted by $\vee_{s,i\in I}\mathfrak{a}_{\bullet,i}$, such that the following holds. If $\mathfrak{c}_{\bullet}\in\mathrm{Fil}^{s}$ satisfies $\mathfrak{a}_{\bullet,i}\subset\mathfrak{c}_{\bullet}$ for any $i\in I$, then we have $\vee_{s,i\in I}\mathfrak{a}_{\bullet,i}\subset\mathfrak{c}_{\bullet}$. Moreover, $\vee_{s,i\in I}\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}^{s}$.
Proof.
Fix $0\in I$. By assumption, the set
$J:=\{\mathfrak{c}_{\bullet}\in\mathrm{Fil}^{s}\mid\mathfrak{a}_{\bullet,i}\subset\mathfrak{c}_{\bullet}\text{ for any }i\in I\}\subset\mathrm{Fil}^{s}$
is non-empty, and $\mathfrak{a}_{\bullet,0}\subset\mathfrak{c}_{\bullet}$ for any $\mathfrak{c}_{\bullet}\in J$. Thus by Definition-Lemma 2.2, the intersection
$\vee_{s,i\in I}\mathfrak{a}_{\bullet,i}\coloneqq\cap_{\mathfrak{c}_{\bullet}\in J}\mathfrak{c}_{\bullet}\in\mathrm{Fil},$
where the linearly boundedness is because $\vee_{s,i\in I}\mathfrak{a}_{\bullet,i}\subset\mathfrak{b}_{\bullet}$ by construction. Now $\vee_{s,i\in I}\mathfrak{a}_{i}\in\mathrm{Fil}^{s}$ by Corollary 2.21. ∎
We now relate the join, the saturated join and the saturation, and give a formula for the saturated join of two saturated filtrations.
Proposition 2.23.
If $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}$, then $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}=\widetilde{\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}}=\widetilde{\mathfrak{a}}_{\bullet}\vee_{s}\widetilde{\mathfrak{b}}_{\bullet}$, and
$\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}}).$ (2.4)
Proof.
By definition, $\widetilde{\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}}$ is a saturated filtration containing both $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$, so by Definition-Lemma 2.22 we have $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}\subset\widetilde{\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}}$. Conversely, by Definition-Lemma 2.22 again $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$ is saturated and $\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}\subset\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$ by definition, hence $\widetilde{\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}}\subset\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$. The inclusion $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}\subset\widetilde{\mathfrak{a}}_{\bullet}\vee_{s}\widetilde{\mathfrak{b}}_{\bullet}$ is obvious. Conversely, if $\mathfrak{a}_{\bullet}\subset\mathfrak{c}_{\bullet}$ for $\mathfrak{c}_{\bullet}\in\mathrm{Fil}^{s}$ then $\widetilde{\mathfrak{a}}_{\bullet}\subset\mathfrak{c}_{\bullet}$ and similarly for $\mathfrak{b}_{\bullet}$. Thus $\widetilde{\mathfrak{a}}_{\bullet}\vee_{s}\widetilde{\mathfrak{b}}_{\bullet}\subset\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$. This proves the first equality.
To prove the formula (2.4), we first rewrite Lemma 2.19 as
$\mathfrak{a}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a}_{\bullet})})$ (2.5)
for $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$. Let $\mathfrak{c}_{\bullet}$ be the $\mathfrak{m}$-filtration defined by the right-handed side of (2.4). Since
$\frac{v}{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}}=\max\{\frac{v}{v(\mathfrak{a}_{\bullet})},\frac{v}{v(\mathfrak{b}_{\bullet})}\}=\max\{\frac{v}{v(\widetilde{\mathfrak{a}}_{\bullet})},\frac{v}{v(\widetilde{\mathfrak{b}}_{\bullet})}\},$
applying the formula (2.5) to $\widetilde{\mathfrak{a}}_{\bullet}$ and $\widetilde{\mathfrak{b}}_{\bullet}$ respectively, we know that $\mathfrak{a}_{\bullet}\subset\widetilde{\mathfrak{a}}_{\bullet}\subset\mathfrak{c}_{\bullet}$ and similarly $\mathfrak{b}_{\bullet}\subset\mathfrak{c}_{\bullet}$. Hence $v(\mathfrak{c}_{\bullet})\leq\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$ for any $v\in\mathrm{Val}_{R,\mathfrak{m}}$. By (2.4), for any $v\in\mathrm{DivVal}_{X,x}$ we have $v(\mathfrak{c}_{\bullet})\geq v(\mathfrak{a}_{\bullet}(v/\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}))=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$. Thus
$v(\mathfrak{c}_{\bullet})=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}.$ (2.6)
We now verify that $\mathfrak{c}_{\bullet}\in\mathrm{Fil}$. Take $C\in\mathbb{R}_{>0}$ such that $\mathfrak{a}_{C\bullet}\subset\mathfrak{m}^{\bullet}$ and $\mathfrak{b}_{C\bullet}\subset\mathfrak{m}^{\bullet}$. Since $\mathfrak{m}^{\bullet}\subset\widetilde{\mathfrak{m}}^{\bullet}$, for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$, we have $Cv(\mathfrak{a}_{\bullet})\geq v(\mathfrak{m}^{\bullet})=v(\widetilde{\mathfrak{m}}^{\bullet})$ and similarly $Cv(\mathfrak{b}_{\bullet})\geq v(\widetilde{\mathfrak{m}}^{\bullet})$. By (2.6), $Cv(\mathfrak{c}_{\bullet})\geq v(\widetilde{\mathfrak{m}}^{\bullet})$. So by Lemma 2.16 $\mathfrak{c}_{\bullet}\subset\widetilde{\mathfrak{m}}^{\bullet}$ is linearly bounded.
By Corollary 2.21, $\mathfrak{c}_{\bullet}\in\mathrm{Fil}^{s}$, which implies $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}\subset\mathfrak{c}_{\bullet}$. By definition and (2.6) we have $v(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet})\leq\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}=v(\mathfrak{c}_{\bullet})$ for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$. Since $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{s}$ by Definition-Lemma 2.22, we may apply Lemma 2.16 to conclude that $\mathfrak{c}_{\bullet}\subset\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$. This proves the formula (2.4) and finishes the proof of the Proposition. ∎
We now show that the lattice operations on $\mathrm{Fil}^{s}$ satisfy the distributive law.
Proposition 2.24.
If $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}\in\mathrm{Fil}^{s}$, then we have $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})=(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet})\cap(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})$.
Proof.
We have $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})\subset(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet})\cap(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})$, since $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})\subset\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$ and $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})\subset\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet}$.
Conversely, assume that $f\in(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet})_{\lambda}\cap(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})_{\lambda}$. Then for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$ we know that
$v(f)/\lambda\geq$ $\max\{v(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}),v(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})\}$
$\geq$ $\max\{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\},\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{c}_{\bullet})\}\}$
$\geq$ $\min\{v(\mathfrak{a}_{\bullet}),\max\{v(\mathfrak{b}_{\bullet}),v(\mathfrak{c}_{\bullet})\}\},$
where the second inequality follows from Proposition 2.23, and the last inequality is easy to check. By Proposition 2.23 again, we get $f\in(\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet}))_{\lambda}$. This shows that $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{c}_{\bullet}\cap\mathfrak{b}_{\bullet})\supset(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})\cap\mathfrak{b}_{\bullet}$ and the equality must hold. ∎
Example 2.25.
We continue to consider the adic filtration $\mathfrak{a}^{\bullet}$ of $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$. By definition,
$(\widetilde{\mathfrak{a}^{\bullet}})_{\lambda}=\{f\in\mathfrak{m}\mid v(f)\geq\lambda\cdot v(\mathfrak{a}^{\bullet})\text{ for any }v\in\mathrm{DivVal}_{R,\mathfrak{m}}\}.$
Since $v(\mathfrak{a}^{\bullet})=\lim_{\lambda\to\infty}v(\mathfrak{a}^{\lceil\lambda\rceil})/\lambda=\lim_{\lambda\to\infty}\lceil\lambda\rceil v(\mathfrak{a})/\lambda=v(\mathfrak{a})$, for $m\in\mathbb{Z}_{>0}$ we have
$(\widetilde{\mathfrak{a}^{\bullet}})_{m}=\{f\in\mathfrak{m}\mid v(f)\geq mv(\mathfrak{a})=v(\mathfrak{a}^{m})\text{ for any }v\in\mathrm{DivVal}_{R,\mathfrak{m}}\}.$
Hence $(\widetilde{\mathfrak{a}^{\bullet})_{m}}=\overline{\mathfrak{a}^{m}}$ is the integral closure of $\mathfrak{a}^{m}$. See for example [HS06, Theorem 6.8.3]. Thus we have a map $\mathcal{I}_{\mathfrak{m}}\to\mathrm{Fil}^{s}$, $\mathfrak{a}\mapsto\widetilde{\mathfrak{a}^{\bullet}}$, which is an injection when restricted to the set $\mathcal{I}_{\mathfrak{m}}^{\mathrm{ic}}$ of integrally closed $\mathfrak{m}$-primary ideals. Note that in general, $\widetilde{\mathfrak{a}^{\bullet}}$ is not an adic filtration. For example, $\widetilde{(\mathfrak{m}^{2})^{\bullet}}$ has $1/2$ as a jumping number.
Even if $\mathfrak{a}$ is integrally closed, the powers of $\mathfrak{a}$ are in general not integrally closed, hence an $\mathfrak{a}$-adic filtration is in general not saturated. By Example 2.4, the join of two adic filtrations is still an adic filtration, which means that the inclusion $\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}\subset\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$ can be strict for general $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}$. It is unclear to the author whether $\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}=\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}$ if $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{s}$.
It is easy to see that $\widetilde{\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}}\subset\widetilde{\mathfrak{a}}_{\bullet}\cap\widetilde{\mathfrak{b}}_{\bullet}$, and the inclusion can be strict. Indeed, let $R=\mathbbm{k}[\![x,y]\!]$, $\mathfrak{a}=(x^{2},y^{10})$ and $\mathfrak{b}=(x^{10},y^{2})$, then it is not hard to get $\overline{\mathfrak{a}}=(x^{2},xy^{5},y^{10})$ and $\overline{\mathfrak{b}}=(x^{10},x^{5}y,y^{2})$. By looking at the Newton polytope of $\mathfrak{a}\cap\mathfrak{b}=(x^{2}y^{2},x^{10},y^{10})$, one can compute
$\overline{\mathfrak{a}\cap\mathfrak{b}}=(x^{2}y^{2},xy^{6},x^{6}y,x^{10},y^{10}).$
In particular, by the calculation for adic filtrations above, $f\coloneqq x^{5}y\in\overline{\mathfrak{a}}\cap\overline{\mathfrak{b}}=(\widetilde{\mathfrak{a}}^{\bullet}\cap\widetilde{\mathfrak{b}}^{\bullet})_{1}$, but $f\notin\overline{\mathfrak{a}\cap\mathfrak{b}}=(\widetilde{\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}})_{1}$.
Next we compute the saturation of an ideal-adic filtration. First note that to prove two saturated filtrations are equal, it suffices to check the equality along an arbitrary unbounded sequence.
Lemma 2.26.
Let $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{s}$. If there exists $\lambda_{k}\in\mathbb{R}_{>0}$ with $\lambda_{k}\to\infty$ such that $\mathfrak{a}_{\lambda_{k}}=\mathfrak{b}_{\lambda_{k}}$, then $\mathfrak{a}_{\bullet}=\mathfrak{b}_{\bullet}$.
Proof.
By assumption, for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$ we have
$v(\mathfrak{a}_{\bullet})=\lim_{k\to\infty}v(\mathfrak{a}_{\lambda_{k}})/\lambda_{k}=\lim_{k\to\infty}v(\mathfrak{b}_{\lambda_{k}})/\lambda_{k}=v(\mathfrak{b}_{\bullet}).$
Hence by Lemma 2.16, $\mathfrak{a}_{\bullet}=\mathfrak{b}_{\bullet}$. ∎
Recall that for an ideal $\mathfrak{a}\in\mathcal{I}_{R}$, the set $\mathrm{RV}(\mathfrak{a})$ of Rees valuations of $\mathfrak{a}$ is the minimal set of divisorial valuations such that for any $m\in\mathbb{Z}_{>0}$,
$\overline{\mathfrak{a}^{m}}=\{f\in R\mid v(f)\geq mv(\mathfrak{a})\text{ for any }v\in\mathrm{RV}(\mathfrak{a})\}.$
This is a finite set for any $\mathfrak{a}$. We refer to [HS06, Chapter 10] for more about Rees valuations.
Proposition 2.27.
Let $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$ be an $\mathfrak{m}$-primary ideal. Then
$\widetilde{\mathfrak{a}^{\bullet}}=\cap_{v\in RV(\mathfrak{a})}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a})}),$ (2.7)
where $RV(\mathfrak{a})$ is the set of Rees valuations of $\mathfrak{a}$.
Proof.
Denote $\mathfrak{b}_{\bullet}\coloneqq\cap_{v\in RV(\mathfrak{a})}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a})})$, which is saturated by Proposition 2.20. By the definition of Rees valuations and the calculation in Example 2.25, we know that $(\widetilde{\mathfrak{a}^{\bullet}})_{m}=\mathfrak{b}_{m}$ for $m\in\mathbb{Z}_{>0}$. Thus (2.7) follows from Lemma 2.26. ∎
2.3. Homogeneous filtrations
Definition 2.28.
A norm $\chi\in\mathcal{N}$ is said to be power-multiplicative, or homogeneous, if it satisfies $\chi(f^{d})=d\cdot\chi(f)$ for any $f\in R$ and $d\in\mathbb{Z}_{>0}$.
A filtration $\mathfrak{a}_{\bullet}$ is said to be homogeneous if $f^{d}\in\mathfrak{a}_{\lambda}$ implies $f\in\mathfrak{a}_{\lambda/d}$ for any $\lambda>0$ and $d\in\mathbb{Z}_{>0}$.
It is easy to see that a saturated filtration is homogeneous, and that a homogeneous filtration satisfies the additional condition in Definition-Lemma 2.8.
Lemma 2.29.
Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration. Then
  1. (1)
    if $\mathfrak{a}_{\bullet}$ is saturated, then it is homogeneous, and
  2. (2)
    if $\mathfrak{a}_{\bullet}$ is homogeneous, then $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda\ll 1$. In particular, $\mathfrak{a}_{\bullet}$ is homogeneous if and only if $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is homogeneous.
Proof.
(1) If $\mathfrak{a}_{\bullet}$ is saturated and $f^{d}\in\mathfrak{a}_{\lambda}$, then $v(f^{d})=dv(f)\geq\lambda\cdot v(\mathfrak{a}_{\bullet})$, which implies $v(f)\geq\lambda/d\cdot v(\mathfrak{a}_{\bullet})$. Hence $f\in\mathfrak{a}_{\lambda/d}$ by definition.
(2) Choose a set $\{f_{1},\ldots,f_{r}\}$ of generators of $\mathfrak{m}$ and $D\in\mathbb{Z}_{>0}$ such that $\mathfrak{m}^{D}\subset\mathfrak{a}_{1}$. Then $f_{i}^{D}\in\mathfrak{a}_{1}$ and hence $f_{i}\in\mathfrak{a}_{1/D}$ by assumption. So $\mathfrak{m}\subset\mathfrak{a}_{1/D}$ and hence $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda<1/D$. The last assertion is straightforward by Definition-Lemma 2.8. ∎
Definition 2.30.
[CP24, Definition 3.5](c.f. [BFJ14, Section 4]) Given a filtration $\mathfrak{a}_{\bullet}$ on $R$, its asymptotic Samuel function $\widehat{\mathrm{ord}}_{\mathfrak{a}_{\bullet}}:R\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is defined by $\widehat{\mathrm{ord}}_{\mathfrak{a}_{\bullet}}(f):=\lim_{k\to\infty}\frac{\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f^{k})}{k}$ for any $f\in R$.
Denote the corresponding filtration by $\widehat{\mathfrak{a}}_{\bullet}$, called the homogenization of $\mathfrak{a}_{\bullet}$.
The limit exists by Feteke’s lemma, see [CP24, Theorem 3.4]. Similar to Theorem 2.18 for the saturation, the homogenization has the following property.
Theorem 2.31.
[CP24, Theorem 5.5] Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration. Then $\widehat{\mathfrak{a}}_{\bullet}$ is the unique largest $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}$ such that $\widehat{\mathrm{ord}_{\mathfrak{a}_{\bullet}}}=\widehat{\mathrm{ord}_{\mathfrak{b}_{\bullet}}}$.
[…]
Lemma 2.32.
A norm $\chi\in\mathcal{N}$ is homogeneous if and only if it is of the form
$\chi=\inf_{v\in I}v,$
where $I$ is a non-empty subset of $X^{\mathrm{an}}_{x}$.
Proof.
If $\chi(f)=\inf_{v\in I}v(f)$, then $\chi$ is homogeneous, since the same is true for each $v$.
The reverse implication is essentially a restatement of the Berkovich maximum modulus principle, applied to the seminormed ring $(R,\chi)$. See, for example, [Jon20, Theorem 2.2.1]
Denote the subspace of $\mathrm{Fil}$ consisting of homogeneous filtrations by $\mathrm{Fil}^{h}$.
2.4. Log canonical thresholds and normalized volumes
We say $x\in(X,\Delta)$ is a klt singularity if $\Delta$ is an $\mathbb{R}$-divisor on $X$ such that $K_{X}+\Delta$ is $\mathbb{R}$-Cartier, and $(X,\Delta)$ is klt as defined in [KM98]. The following invariant was first introduced by C. Li [Li18] and plays an important role in the study of Fano cone singularities and K-stability of Fano varieties. [^3] The original motivation to introduce the metric $d_{1}$ is to study local volumes. Some results in this direction will appear elsewhere.
Definition 2.33.
For a klt singularity $x\in(X,\Delta)$, the normalized volume function $\widehat{\mathrm{vol}}_{(X,\Delta),x}:\mathrm{Val}_{X,x}\to(0,+\infty]$ is defined by
$\widehat{\mathrm{vol}}_{(X,\Delta),x}(v)\coloneqq\left\{\begin{aligned} &A_{X,\Delta}(v)^{n}\cdot\mathrm{vol}(v),&\text{ if }A_{X,\Delta}(v)<+\infty\\ &+\infty,&\text{ if }A_{X,\Delta}(v)=+\infty,\end{aligned}\right.$
where $A_{X,\Delta}(v)$ is the log discrepancy of $v$ as defined in [JM12, BdFFU15]. The local volume of a klt singularity $x\in(X,\Delta)$ is defined as
$\widehat{\mathrm{vol}}(x,X,\Delta)\coloneqq\inf_{v\in\mathrm{Val}_{X,x}}\widehat{\mathrm{vol}}_{(X,\Delta),x}(v).$
The above infimum is indeed a minimum by [Blu18, Xu20].
Denote $\mathrm{Val}^{<+\infty}_{X,x}:=\{v\in\mathrm{Val}_{X,x}\mid A_{X}(v)<+\infty\}$.
Recall that the log canonical threshold (lct) of an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ is defined to be
$\mathrm{lct}(X;\mathfrak{a}_{\bullet})\coloneqq\inf_{v\in\mathrm{Val}^{<+\infty}_{X,x}}\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}.$
We will write $\mathrm{lct}(\mathfrak{a}_{\bullet})$ when there is no ambiguity. For more properties of lct, we refer to [JM12]. We refer to [LLX20, Zhu23] for a comprehensive survey on normalized volumes.
Lemma 2.34.
(c.f. [Zhu24, Lemma 3.4]) There exists $c_{0}=c_{0}(n)>0$ depending only one $n$ such that for any $n$-dimensional klt singularity $x\in X$ and $v\in\mathrm{Val}_{X,x}$,
$\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}\leq c_{0}\cdot\frac{\widehat{\mathrm{vol}}_{x,X}(v)}{\widehat{\mathrm{vol}}(x,X)}\cdot\mathrm{lct}(X;\mathfrak{a}_{\bullet})$ (2.8)
for any $\mathfrak{a}_{\bullet}\in\mathrm{Fil}_{x}(X)$.
The following proof is inspired by communications with Z. Zhuang.
Proof.
Note that $v(\mathfrak{a}_{\bullet})=\lim_{m}v(\mathfrak{a}_{m})/m$ and $\mathrm{lct}(X;\mathfrak{a}_{\bullet})=\lim_{m}m\cdot\mathrm{lct}(X;\mathfrak{a}_{m})$, it is enough to prove (2.8) for $\mathfrak{m}$-primary ideals,
$\frac{A_{X}(v)}{v(\mathfrak{a})}\leq c_{0}\cdot\frac{\widehat{\mathrm{vol}}_{x,X}(v)}{\widehat{\mathrm{vol}}(x,X)}\cdot\mathrm{lct}(X;\mathfrak{a}).$
Assume $\mathfrak{a}=(f_{1},\ldots,f_{m})$. By lifting $\mathfrak{a}$ to a power, we may assume that $\mathrm{lct}(\mathfrak{a})<1$. Then by Bertini’s theorem, for general $c_{1},\ldots,c_{m}\in\mathbbm{k}$, $\mathrm{lct}(\sum_{i=1}^{m}c_{i}f_{i})=\mathrm{lct}(\mathfrak{a})$. Moreover, by definition, for $f\neq g\in\mathfrak{a}$, $v(f+cg)>\min\{v(f),v(g)\}$ for at most one $c\in\mathbbm{k}$. Hence one can choose the $c_{i}\in\mathbbm{k}$ such that $v(\sum c_{i}f_{i})=\min_{1\leq i\leq m}\{v(f_{i})\}=v(\mathfrak{a})$. So we are further reduced to the case of a principal ideal $\mathfrak{a}=(f)$. Now the lemma follows from [Zhu24, Lemma 3.4]. ∎
2.5. Partially ordered sets and lattices
We recall some basic definitions regarding lattices in order theory.
Definition 2.35.
A partially ordered set $(\mathcal{S},\leq)$ is called a lattice if for any $a,b\in\mathcal{S}$,
  1. (1)
    there exists $a\vee b\in\mathcal{S}$, called the supremum, or join, such that for any $c\in\mathcal{S}$, $c\geq a$ and $c\geq b$ imply $c\geq a\vee b$, and
  2. (2)
    there exists $a\wedge b\in\mathcal{S}$, called the infimum, or meet, such that for any $c\in\mathcal{S}$, $c\leq a$ and $c\leq b$ imply $c\leq a\wedge b$.
A lattice $\mathcal{S}$ is modular if for any $a,b,c\in\mathcal{S}$ with $a\leq b$, we have $a\vee(c\wedge b)=(a\wedge c)\vee b$. A lattice $\mathcal{S}$ is distributive if for any $a,b,c\in\mathcal{S}$, we have $a\vee(b\wedge c)=(a\wedge b)\vee(a\wedge c)$.
A lattice $\mathcal{S}$ is complete if any subset $A\subset\mathcal{S}$ has an infimum $a=\inf A$ and a supremum $b=\sup A$.
2.6. Metric spaces
Recall that a pseudometric on a set $S$ is a function $d:S\times S\to\mathbb{R}_{\geq 0}$ such that
  • (non-negative) $d(x,y)\geq 0$ for any $x,y\in S$,
  • (symmetric) $d(x,y)=d(y,x)$ for any $x,y\in S$, and
  • (triangle inequality) $d(x,z)\leq d(x,y)+d(y,z)$ for any $x,y,z\in S$.
A pseudometric is a metric if and only if it is non-degenerate, that is, $d(x,y)=0$ if and only if $x=y$.
Definition 2.36.
A point $z$ in a metric space $(S,d)$ is said to be between two points $x$ and $y$, if all three points are distinct, and $d(x,y)=d(x,z)+d(z,x)$.
A metric space $(S,d)$ is convex if for any two distinct points $x,y\in S$, there exists a point $z$ between $x$ and $y$.
Given a pseudometric space $(S,d)$, its Hausdorff quotient is the quotient space $(\bar{S}\coloneqq S/\sim,\bar{d})$, where $x\sim y$ if and only if $d(x,y)=0$, and $\bar{d}$ is the natural induced metric on $\bar{S}$.
Recall that the length of a continuous curve $\gamma:[a,b]\to S$ in a metric space $(S,d)$ is defined by
$\ell(\gamma):=\sup_{\mathcal{P}}\sum_{i=1}^{n}d(x_{i-1},x_{i}),$
where the supremum is taken over all finite partitions $\mathcal{P}=\{a=x_{0}\leq x_{1}\leq\ldots\leq x_{n}=b\}$ of the interval $[a,b]$.
Definition 2.37.
A metric space $(S,d)$ is a length space if $d$ coincides with the intrinsic metric, that is, $d(x,y)=\inf_{\gamma}\ell(\gamma)=:d_{I}(x,y)$ for any $x,y\in S$, where the infimum is taken over all continuous curves $\gamma:[0,1]\to S$ with $\gamma(0)=x$ and $\gamma(1)=y$.
$(S,d)$ is a geodesic space if the above infimum is a minimum, that is, the distance can always be calculated by a continuous curve.
2.7. Space of valuations and filtrations
From the previous sections we have the following subspaces of $\mathrm{Val}_{X,x}$:
  • the space of divisorial valuations, denoted by $\mathrm{DivVal}_{R,\mathfrak{m}}$,
  • the space of quasi-monomial valuations when $R$ contains a field, denoted by $\mathrm{QMVal}_{R,\mathfrak{m}}$,
  • the space of valuations with finite log discrepancy, when $x\in X=\mathrm{Spec}R$ is a klt singularity over a field of characteristic 00, denoted by $\mathrm{Val}^{<+\infty}_{X,x}$, and
  • the space of valuations with positive volume, denoted by $\mathrm{Val}^{+}_{X,x}$, which is equal to the space of linearly bounded valuations by Lemma 2.13.
Moreover, we have the following spaces of linearly bounded filtrations on $R$:
  • the space of linearly bounded $\mathfrak{m}$-filtrations, denoted by $\mathrm{Fil}_{R,\mathfrak{m}}$,
  • the space of power-multiplicative $\mathfrak{m}$-filtrations which are linearly bounded, denoted by $\mathrm{Fil}^{h}_{R,\mathfrak{m}}$, and
  • the space of saturated $\mathfrak{m}$-filtrations which are linearly bounded, denoted by $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$.
We have the following inclusions
$\mathrm{DivVal}_{X,x}\subset\mathrm{QMVal}_{X,x}\subset\mathrm{Val}^{<+\infty}_{X,x}\subset\mathrm{Val}^{+}_{X,x}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{h}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$ (2.9)
where the third inclusion follows from Li’s properness estimate [Li18, Theorem 1.4]. By Lemma 2.17, to compute the saturation of a filtration, one can use any of the above spaces of valuations.
2.7.1. Weak topologies on $\mathrm{Fil}$
There are more ways to define a topology on the space $\mathrm{Fil}$ (and $\mathrm{Fil}^{s}$ and $\mathrm{Fil}^{s}$, respectively), depending on the point of view.
For example, a filtration $\mathfrak{a}_{\bullet}$ acts naturally on the ring $R$ as a function $\chi_{\mathfrak{a}_{\bullet}}$ (or dually). The product topology, or the “weak-* topology” induced by this action will be called the weak topology, which is the coarest topology such that for any $f\in R\backslash\{0\}$, the function $\phi_{f}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)$ is continuous. In other words, a subbase of the weak topology is given by $\{\phi_{f}^{-1}(U)\mid U\subset\mathbb{R}_{\geq 0}\text{ is open, }f\in R\backslash\{0\}\}$. Under this topology, all valuation spaces in (2.9) equipped with the weak topology, defined in [JM12], are subspaces of $\mathrm{Fil}$. It is not hard to check that a sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}$ converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$ if and only of $\chi_{k}(f)\to\chi(f)$ for any $f\in R\backslash\{0\}$.
Similarly, a valuation has a natural action on the spaces of filtrations (or dually), that is, any $v\in\mathrm{Val}_{X,x}$ gives a function $\mathrm{ev}_{v}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto v(\mathfrak{a}_{\bullet})$. Hence one can consider the “weak topology” induced by this action, that is, the coarest topology on $\mathrm{Fil}^{s}$ such that $\mathrm{ev}_{v}$ is continuous for any $v\in X^{{}^{\mathrm{an}}}$, $\mathrm{Val}_{X,x}$, $\mathrm{Val}^{+}_{X,x}$ or $\mathrm{DivVal}_{X,x}$, respectively. Similarly one can check that a sequence $\{\mathfrak{a}_{\bullet,k}\}$ converges to $\mathfrak{a}_{\bullet}$ under these topologies if any only if $v(\mathfrak{a}_{\bullet,k})\to v(\mathfrak{a}_{\bullet})$ for any $v$ in the corresponding valuation space. In this paper, we will mainly consider the topology defined using $\mathrm{Val}^{+}_{X,x}$, and call it the $+$-topology. The product topology defined using $\mathrm{DivVal}_{X,x}$ as above was considered in [BdFFU15] for $b$-divisors. We will thus call it the coefficientwise topology.
Remark 2.38.
Since $v(\mathfrak{a}_{\bullet})=v(\widetilde{\mathfrak{a}}_{\bullet})$ by [BLQ24, Proposition 3.9] for $v\in\mathrm{DivVal}_{X,x}$, the functions $\mathrm{ev}_{v}$ do not distinguish $\mathfrak{a}_{\bullet}$ and its saturation. Hence the $+$-topology (and the coefficientwise topology) only makes sense on $\mathrm{Fil}^{s}$. Similar issues appear if we define the topology on $\mathrm{Fil}$ using other valuation spaces. But all the topologies are well-defined on $\mathrm{Fil}^{s}$.
Later we will define metrics on the space $\mathrm{Fil}^{s}$. See Section 4.3 for a comparison of these topologies.
3. Darvas metrics on the space of filtrations
In this section, we introduce the pseudometric $d_{1}$ and consider its first properties.
3.1. Darvas metrics and multiplicities
Throughout this section we will use the language of filtrations. Recall that $\mathrm{Fil}=\mathrm{Fil}_{R,\mathfrak{m}}$ denotes the set of all linearly bounded $\mathfrak{m}$-filtrations on $R$.
Definition 3.1.
We define a function $d_{1}:\mathrm{Fil}\times\mathrm{Fil}\to\mathbb{R}_{\geq 0}$ as follows.
Given $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}$. If $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$, then let
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}):=\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}).$ (3.1)
In general, define
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}):=d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{b}_{\bullet}).$ (3.2)
Proposition 3.2.
$d_{1}$ satisfies the triangle inequality, that is, for $\mathfrak{a}_{\bullet}$, $\mathfrak{b}_{\bullet}$, $\mathfrak{c}_{\bullet}\in\mathrm{Fil}$, we have
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$ (3.3)
Hence $d_{1}$ is a pseudometric on $\mathrm{Fil}$.
Proof.
By definition, we have
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet})=2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{c}_{\bullet})$
and
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet})=2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet})+2\mathrm{e}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{c}_{\bullet})$
Hence to prove (3.3), it suffices to show that
$\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})\leq\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})+\mathrm{e}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}),$
which will follow immediately from the following comparison on colengths
$\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{c}_{\lambda})\leq\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})+\ell(R/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})-\ell(R/\mathfrak{b}_{\lambda})$ (3.4)
for any $\lambda\in\mathbb{R}_{>0}$. But we compute
$\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})+$ $\ell(R/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})-\ell(R/\mathfrak{b}_{\lambda})-\ell(R/\mathfrak{a}_{\lambda})=\ell(\mathfrak{a}_{\lambda}/\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})+\ell(\mathfrak{b}_{\lambda}/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})$
$=$ $\ell((\mathfrak{a}_{\lambda}+\mathfrak{b}_{\lambda})/\mathfrak{b}_{\lambda})+\ell(\mathfrak{b}_{\lambda}/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})$
$=$ $\ell((\mathfrak{a}_{\lambda}+\mathfrak{b}_{\lambda})/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})$
$\geq$ $\ell((\mathfrak{a}_{\lambda}+\mathfrak{b}_{\lambda})/(\mathfrak{a}_{\lambda}+\mathfrak{b}_{\lambda})\cap\mathfrak{c}_{\lambda})$
$=$ $\ell((\mathfrak{a}_{\lambda}+\mathfrak{b}_{\lambda}+\mathfrak{c}_{\lambda})/\mathfrak{c}_{\lambda})$
$\geq$ $\ell((\mathfrak{a}_{\lambda}+\mathfrak{c}_{\lambda})/\mathfrak{c}_{\lambda})$
$=$ $\ell(\mathfrak{a}_{\lambda}/\mathfrak{a}_{\lambda}\cap\mathfrak{c}_{\lambda})=\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{c}_{\lambda})-\ell(R/\mathfrak{a}_{\lambda}).$
This proves (3.4), and hence the triangle inequality (3.3) holds.
Clearly $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet})=0$, and $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet})$ by definition. So we know that $(\mathrm{Fil},d_{1})$ is a pseudometric space. ∎
Using the identification between $\mathrm{Fil}$ and $\mathcal{N}$ Lemma 2.8, we get a pseudometric $d_{1}$ on $\mathcal{N}$. The following lemma asserts that $(\mathrm{Fil},d_{1})$ is a rooftop metric space in the sense of [Xia23], which is an analogue of [Dar19, Proposition 3.35] and [BJ21, Lemma 3.2].
Lemma 3.3.
For $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}\in\mathrm{Fil}$, we have
$d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$ (3.5)
Proof.
By definition, we have
$d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet})+\mathrm{e}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{c}_{\bullet})$
and
$d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})=\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})+\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet}).$
So by symmetry, it suffices to prove the inequality
$\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})\leq\mathrm{e}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}),$
which follows from the inclusion
$(\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})/(\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})\cong$ $(\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda}+\mathfrak{c}_{\lambda})/\mathfrak{c}_{\lambda}$
$\subset(\mathfrak{b}_{\lambda}+\mathfrak{c}_{\lambda})/\mathfrak{c}_{\lambda}\cong$ $\mathfrak{b}_{\lambda}/(\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda}).$
This finishes the proof. ∎
3.2. Measures and geodesics
Recall that given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$, the geodesic between them is defined to be the segment $\{\mathfrak{a}_{\bullet,t}\}_{t\in[0,1]}$ of $\mathfrak{m}$-filtrations, where
$\mathfrak{a}_{\lambda,t}=\sum_{(1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}.$
The goal of this section is to justify the term geodesic, in the sense that it computes the distance between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$. In other words, it is a distance-minimizing continuous curve with respect to the metric $d_{1}$.
We first give a formula for the geodesic between two saturated filtrations, and use it to show that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}^{s}$ when $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$..
Proposition 3.4.
Let $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$. Then for any $t\in(0,1)$, we have
$\mathfrak{a}_{\bullet,t}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{c_{t,v}}),$ (3.6)
where
$c_{t,v}\coloneqq\frac{v(\mathfrak{a}_{\bullet,0})v(\mathfrak{a}_{\bullet,1})}{tv(\mathfrak{a}_{\bullet,0})+(1-t)v(\mathfrak{a}_{\bullet,1})}.$
In particular, $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}^{s}$ for any $t\in(0,1)$.
Proof.
Denote the right-handed side of (3.6) by $\mathfrak{c}_{\bullet,t}$ and fix $\lambda\in\mathbb{R}_{>0}$. First assume $f\in\mathfrak{c}_{\lambda,t}$. Then for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$, $v(f)\geq\lambda c_{t,v}$. Since $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$, this implies that $f\in\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$, where $\mu\coloneqq\frac{\lambda v(\mathfrak{a}_{\bullet,1})}{tv(\mathfrak{a}_{\bullet,0})+(1-t)v(\mathfrak{a}_{\bullet,1})}$ and $\nu\coloneqq\frac{\lambda v(\mathfrak{a}_{\bullet,0})}{tv(\mathfrak{a}_{\bullet,0})+(1-t)v(\mathfrak{a}_{\bullet,1})}$. Since $(1-t)\mu+t\nu=\lambda$, we get $f\in\mathfrak{a}_{\lambda,t}$ and hence $\mathfrak{c}_{\bullet,t}\subset\mathfrak{a}_{\bullet,t}$.
Conversely, if $f\in\mathfrak{a}_{\lambda,t}$, then for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$, we have
$v(f)\geq$ $v\left(\sum_{(1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\right)=\min_{(1-t)\mu+t\nu=\lambda}\{v(\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1})\}$
$\geq$ $\min_{(1-t)\mu+t\nu=\lambda}\max\{v(\mathfrak{a}_{\mu,0}),v(\mathfrak{a}_{\nu,1})\}$
$\geq$ $\min_{0\leq\mu\leq\lambda/(1-t)}\max\{\mu v(\mathfrak{a}_{\bullet,0}),\frac{\lambda-(1-t)\mu}{t}v(\mathfrak{a}_{\bullet,1})\}.$
Now it is not hard to check that the last minimum of the piecewise linear function of $\mu$ is achieved at $\mu_{0}:=\frac{\lambda v(\mathfrak{a}_{\bullet,1})}{tv(\mathfrak{a}_{\bullet,0})+(1-t)v(\mathfrak{a}_{\bullet,1})}\in(0,\lambda/(1-t))$, with value $\mu_{0}v(\mathfrak{a}_{\bullet,0})=\lambda c_{t,v}$. This shows that $f\in\mathfrak{a}_{\lambda}(v/c_{t,v})$. Since $v$ is arbitrary, we get $f\in\mathfrak{c}_{\lambda,t}$. Thus $\mathfrak{a}_{\bullet,t}\subset\mathfrak{c}_{\bullet,t}$ and the equality holds.
It is easy to check that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}$ for any $t\in(0,1)$ (see also the proof of Proposition 3.6 below), hence by Proposition 2.20 we know that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}^{s}$. The proof is finished. ∎
3.2.1. Distance-minimizing property
We first make the following easy but useful observation.
Lemma 3.5.
Let $\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}$ for $i=0,1$, and let $\{\mathfrak{a}_{\bullet,t}\}_{t\in[0,1]}$ be the geodesic between them. For $0\leq t<t^{\prime}\leq 1$, we have $\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,t^{\prime}}\subset\mathfrak{a}_{\bullet,t}$.
Proof.
For any $\lambda\in\mathbb{R}_{>0}$, we compute
$\mathfrak{a}_{\lambda,t^{\prime}}\cap\mathfrak{a}_{\lambda,0}=$ $\left(\sum_{(1-t^{\prime})\mu+t^{\prime}\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\right)\cap\mathfrak{a}_{\lambda,0}$
$=$ $\sum_{\mu\leq\lambda}\mathfrak{a}_{\lambda,0}\cap\mathfrak{a}_{\nu,1}+\sum_{\mu>\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$
$=$ $\mathfrak{a}_{\lambda,0}\cap\mathfrak{a}_{\lambda,1}+\sum_{\mu>\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}.$
Since $\mathfrak{a}_{\lambda,0}\cap\mathfrak{a}_{\lambda,1}\subset\mathfrak{a}_{\lambda,s}$ for any $s\in[0,1]$, it suffices to show that for any $(\mu,\nu)\in\mathbb{R}_{\geq 0}^{2}$ satisfying $(1-t^{\prime})\mu+t^{\prime}\nu=\lambda$ and $\mu>\lambda$, we have
$\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\lambda,t}.$ (3.7)
To see this, let $\mu^{\prime}:=\frac{(1-t^{\prime})\mu+(t^{\prime}-t)\nu}{1-t}$, then $(1-t)\mu+t\nu=(1-t^{\prime})\mu+t^{\prime}\nu=\lambda$, and we have
$\mu-\mu^{\prime}=\frac{(1-t)\mu-(1-t^{\prime})\mu-(t^{\prime}-t)\nu}{1-t}=\frac{(t^{\prime}-t)(\mu-\nu)}{1-t}>0,$
where the last inequality follows from the condition $\mu>\lambda>\nu$. This implies
$\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\mu^{\prime},0}\cap\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\lambda,t},$
hence (3.7) holds and the proof is finished. ∎
Next we prove that the geodesic is always distance-minimizing.
Proposition 3.6.
Let $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$ and let $\{\mathfrak{a}_{\bullet,t}\}_{t\in[0,1]}$ be the geodesic between them. Then for any $0\leq t<t^{\prime}\leq 1$, we have
$d_{1}(a_{\bullet,0},\mathfrak{a}_{\bullet,t^{\prime}})=d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{a}_{\bullet,t^{\prime}}).$ (3.8)
Proof.
For simplicity, denote $\mathfrak{b}_{\bullet}:=\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,t}$, $\mathfrak{b}^{\prime}_{\bullet}:=\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,t^{\prime}}$ and $\mathfrak{c}_{\bullet}:=\mathfrak{a}_{\bullet,t}\cap\mathfrak{a}_{\bullet,t^{\prime}}$. By Lemma 3.5 we know that $\mathfrak{b}^{\prime}_{\bullet}\subset\mathfrak{b}_{\bullet}\subset\mathfrak{a}_{\bullet,0}$ and $\mathfrak{b}^{\prime}_{\bullet}\subset\mathfrak{c}_{\bullet}\subset\mathfrak{a}_{\bullet,t^{\prime}}$. We claim that for any $\lambda\in\mathbb{R}_{>0}$,
$\mathfrak{a}_{\lambda,t}/\mathfrak{b}_{\lambda}\cong\mathfrak{c}_{\lambda}/\mathfrak{b}^{\prime}_{\lambda}.$ (3.9)
Granting the claim for now, we know that $d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet,t})=d_{1}(\mathfrak{b}^{\prime}_{\bullet},\mathfrak{c}_{\bullet})$. So we get
$d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{c}_{\bullet})=d_{1}(\mathfrak{b}^{\prime}_{\bullet},\mathfrak{c}_{\bullet})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{c}_{\bullet})=d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{b}^{\prime}_{\bullet})=d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{b}^{\prime}_{\bullet}),$
which implies $d_{1}(\mathfrak{b}_{\bullet},\mathfrak{b}^{\prime}_{\bullet})=d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{c}_{\bullet})$. Now we compute
$d_{1}(a_{\bullet,0},\mathfrak{a}_{\bullet,t^{\prime}})=$ $d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{b}^{\prime}_{\bullet})+d_{1}(\mathfrak{b}^{\prime}_{\bullet},\mathfrak{a}_{\bullet,t^{\prime}})$
$=$ $d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{b}^{\prime}_{\bullet})+d_{1}(\mathfrak{b}^{\prime}_{\bullet},\mathfrak{c}_{\bullet})+d_{1}(\mathfrak{c}_{\bullet},\mathfrak{a}_{\bullet,t^{\prime}})$
$=$ $d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{c}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{c}_{\bullet},\mathfrak{a}_{\bullet,t^{\prime}})$
$=$ $d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{a}_{\bullet,t^{\prime}})$
It remains to prove the claim (3.9). Indeed, by definition, any $f\in\mathfrak{a}_{\lambda,t}$ can be written as
$f=\sum_{(1-t)\mu+t\nu=\lambda}f_{\mu,\nu},$
where $f_{\mu,\nu}\in\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$. If $\mu\geq\lambda$, then $f_{\mu,\nu}\in\mathfrak{a}_{\mu,0}\subset\mathfrak{a}_{\lambda,0}$. If $\mu\leq\lambda$, then $\nu\geq\lambda$, and hence $f_{\mu,\nu}\in\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\lambda,1}$. Thus we have
$f=\sum_{\mu\leq\lambda}f_{\mu,\nu}+\sum_{\mu\geq\lambda}f_{\mu,\nu}\in\mathfrak{a}_{\lambda,t}\cap\mathfrak{a}_{\lambda,1}+\mathfrak{a}_{\lambda,0},$
and hence $\mathfrak{c}_{\lambda}\subset\mathfrak{a}_{\lambda,t}\subset\mathfrak{a}_{\lambda,t}\cap\mathfrak{a}_{\lambda,1}+\mathfrak{a}_{\lambda,0}\subset\mathfrak{c}_{\lambda}+\mathfrak{a}_{\lambda,0}$, where the last inclusion follows from Lemma 3.5. So we get
$\mathfrak{a}_{\lambda,t}/\mathfrak{b}_{\lambda}\cong(\mathfrak{a}_{\lambda,t}+\mathfrak{a}_{\lambda,0})/\mathfrak{a}_{\lambda,0}=(\mathfrak{c}_{\lambda}+\mathfrak{a}_{\lambda,0})/\mathfrak{a}_{\lambda,0}\cong\mathfrak{c}_{\lambda}/(\mathfrak{c}_{\lambda}\cap\mathfrak{a}_{\lambda,0})=\mathfrak{c}_{\lambda}/\mathfrak{b}^{\prime}_{\lambda},$
where the last equality follows again from Lemma 3.5. This proves (3.9), and finishes the proof of the proposition. ∎
Remark 3.7.
When $R$ contains a field, the proposition can be proved (easily) using the measure as in the next section. We present the above more tedious but elementary argument, which does not rely on this assumption.
3.2.2. Duistermaat-Heckman measures and continuity
In this section $R$ contains a field $\mathbbm{k}$. We briefly recall the measure $\mu$ introduced in [BLQ24, Section 4] encoding multiplicities and its properties.
Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$. Define the function $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ by
$H(x,y):=\mathrm{e}(\mathfrak{a}_{x\bullet,0}\cap\mathfrak{a}_{y\bullet,1}),$
where we adopt the convention $\mathrm{e}(R)=0$. Fix $D\in\mathbb{Z}_{>1}$ such that $\mathfrak{a}_{D\bullet,0}\subset\mathfrak{a}_{\bullet,1}$ and $\mathfrak{a}_{D\bullet,1}\subset\mathfrak{a}_{\bullet,0}$.
Lemma 3.8.
[BLQ24] The distributional derivative
$\mu:=-\frac{\partial^{2}H}{\partial x\partial y}$
is a measure on $\mathbb{R}^{2}$ and for any $a,b\in\mathbb{R}_{>0}$, we have
  1. (1)
    $\mu(\{(1-t)x+ty<a\})=\mathrm{e}(\mathfrak{a}_{a\bullet,t})=a^{n}\mathrm{e}(\mathfrak{a}_{\bullet,t})$,
  2. (2)
    $\mu\left(\{x<a\}\cup\{y<b\}\right)=\mathrm{e}(\mathfrak{a}_{a\bullet,0}\cap\mathfrak{a}_{b\bullet,1})$,
  3. (3)
    $\mu$ is homogeneous of degree $n$, that is, $\mu(cA))=c^{n}\mu(A)$ for any Borel set $A\subset\mathbb{R}^{2}$ and $c\in\mathbb{R}_{>0}$, and
  4. (4)
    $\mathrm{supp}(\mu)\subset\{(x,y)\in\mathbb{R}_{\geq 0}^{2}\mid\frac{1}{D}x\leq y\leq Dx\}$.
Proof.
The first two statements are [BLQ24, Proposition 4.7] and the last two are [BLQ24, Proposition 4.9]
Lemma 3.9.
Fix $(a,b)\in\mathbb{R}^{2}$. Let $A:=[a+\alpha]\times[b+\beta]\subset\mathbb{R}^{2}$ be a rectangle, where $\alpha,\beta\in\mathbb{R}_{>0}$. Then there exists $M:=M(a,b)\in\mathbb{R}_{>0}$ such that $\mu(A)\leq M\alpha\beta$ when $\alpha,\beta\leq 1$.
Proof.
We may assume that $(a,b)\in\mathbb{R}_{\geq 0}^{2}$. Using the inequality
$(x+\delta)^{n}-x^{n}\leq n(x+\delta)^{n-1}\delta$
for $x\geq 0$ and $\delta>0$, we compute
$\mu(A)=$ $\mu([0,a+\alpha]\times[b,b+\beta])-\mu([0,a]\times[b,b+\beta])$
$=$ $((a+\alpha)^{n}-a^{n})\mu([0,1]\times[b,b+\beta])$
$=$ $((a+\alpha)^{n}-a^{n})(((b+\beta)^{n}-b^{n})\mu([0,1]\times[0,1])$
$\leq$ $((a+\alpha)^{n}-a^{n})(((b+\beta)^{n}-b^{n})\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})$
$\leq$ $n^{2}(a+1)^{n-1}(b+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})\cdot\alpha\beta,$
where the second and the third equality follow from Lemma 3.8(3), and the first inequality follows from Lemma 3.8(2). Hence proof is finished with $M:=n^{2}(a+1)^{n-1}(b+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})$. ∎
Now we can prove that the function $t\mapsto\mathfrak{a}_{\bullet,t}$ is Lipschitz continuous with respect to $d_{1}$.
Lemma 3.10.
Suppose $R$ contains a field. Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$, the map $[0,1]\to\mathrm{Fil}$, $t\mapsto\mathfrak{a}_{\bullet,t}$ is Lipschitz continuous, where $[0,1]$ is equipped with the Euclidean topology and $\mathrm{Fil}$ is equipped with the topology induced by the pseudometric $d_{1}$.
Proof.
Given $0\leq t_{1}<t_{2}\leq 1$, denote $\Delta_{j}:=\{(x,y)\in\mathbb{R}^{2}_{>0}\mid(1-t_{j})x+t_{j}y\leq 1\}$ for $j=1,2$. An easy calculation using Lemma 3.8(1) gives
$d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})=\mu(\Delta_{1}\backslash\Delta_{2})+\mu(\Delta_{2}\backslash\Delta_{1}).$ (3.10)
First assume that $10<t_{1}<t_{2}<$ 1. By Lemma 3.8(4) we have
$\mu(\Delta_{1}\backslash\Delta_{2})=\mu((\Delta_{1}\backslash\Delta_{2})\cap\{y\leq Dx\})=\mu(\Delta)$
where $\Delta:=\{(x,y)\in\mathbb{R}^{2}_{\geq 0}\mid 0\leq x\leq 1,y\leq Dx,(1-t_{2})x+t_{2}y\leq 1\leq(1-t_{1})x+t_{1}y\})$. By an elementary calculation, the area of $\Delta$ is at most
$\mathrm{Area}(\Delta)\leq S:=\frac{D(D-1)(t_{2}-t_{1})}{(1-t_{1}+t_{1}D)(1-t_{2}+t_{2}D)}\leq D(D-1)(t_{2}-t_{1}).$
Note that $(x,y)\in\Delta$ satisfies $x\leq 1$ and $y\leq D$. So $\Delta$ can be covered by finitely many rectangles $R_{j}$ of the form $[a_{j},a_{j}^{\prime}]\times[b_{j},b_{j}^{\prime}]$, satisfying
$a_{j}\leq 1,\ b_{j}\leq D,\text{ and }\sum_{j}\mathrm{Area}(R_{j})\leq 2S.$
Hence we may apply Lemma 3.9 to get
$\mu(\Delta_{1}\backslash\Delta_{2})\leq$ $n^{2}2^{n-1}(D+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})\sum_{j}\mathrm{Area}(R_{j})$
$=$ $n^{2}2^{n}(D+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})S$
$\leq$ $n^{2}2^{n}(D+1)^{n-1}D(D-1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})(t_{2}-t_{1})$
Similarly $\mu(\Delta_{1}\backslash\Delta_{1})$ satisfies the same estimate. Hence by (3.10), we have
$d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n+1}(D+1)^{n-1}D(D-1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})(t_{2}-t_{1}).$ (3.11)
Now assume that $t_{1}=0$. We may assume $t_{2}<\frac{1}{2}$. We have
$\mu(\Delta_{1}\backslash\Delta_{2})=$ $\mu((\Delta_{1}\backslash\Delta_{2})\cap\{y\leq Dx\})$
$\leq$ $\mu([\frac{1-t_{2}D}{1-t_{2}},1]\times[1,D])$
$\leq$ $n^{2}(\frac{1-t_{2}D}{1-t_{2}})^{n-1}2^{n-1}(D-1)\frac{(D-1)t_{2}}{1-t_{2}}$
$\leq$ $n^{2}2^{2n-1}(D-1)^{2}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})t_{2},$
where the first equality follows from Lemma 3.8(4) and the second inequality follows from Lemma 3.9, and
$\mu(\Delta_{2}\backslash\Delta_{1})\leq$ $\mu([1,\frac{1}{1-t_{2}}]\times[0,1])$
$\leq$ $n^{2}2^{n-1}\frac{t_{2}}{1-t_{2}}$
$\leq$ $n^{2}2^{n}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})t_{2}$
where the second inequality follows from Lemma 3.9. Hence
$d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n}(2^{n-1}(D-1)^{2}+1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})t_{2}.$ (3.12)
The case when $t_{2}=1$ can be treated similarly. Combining the cases together, we get
$d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})\leq L(t_{2}-t_{1}).$
where the Lipschitz constant can be taken to be
$L:=n^{2}2^{n+1}(D+1)^{n-1}D(D-1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1}),$
where we used the assumption that $D\in\mathbb{Z}_{>1}$. ∎
Remark 3.11.
It is desirable to have a proof for Lemma 3.10 without involving the measure, and thus one can remove the assumption that $R$ contains a field.
3.3. Continuity along monotonic sequences
In this section we prove some convergence result along monotonic results.
3.3.1. Okounkov bodies and continuity along increasing sequences
In this section, using the local variant of Okounkov bodies introduced in [LM09, KK14], we prove that the multiplicity function $\mathrm{e}(\bullet)$ is continuous along increasing sequences of filtrations.
Given a semigroup $S\subset\mathbb{Z}^{n+1}_{\geq 0}$, denote its section at $m\in\mathbb{Z}_{>0}$ by $S_{m}:=S\cap(\mathbb{Z}^{n}_{\geq 0}\times\{m\})$. Recall that the closed convex cone $C(S)\subset\mathbb{R}^{n+1}_{\geq 0}$ of $S$ is defined to be the closure of the set $\{\sum_{i}a_{i}s_{i}\mid a_{i}\in\mathbb{R}_{\geq 0},\ s_{i}\in S\}$, and its Newton-Okounkov body is
$\Delta(S):=C(S)\cap(\mathbb{R}^{n}_{\geq 0}\times\{1\}).$
Lemma 3.12.
[LM09] Let $\Gamma\subset\mathbb{Z}_{\geq 0}^{n+1}$ be a semigroup. If $\Gamma$ satisfies the following
  1. (1)
    $\Gamma_{0}=\{\bm{0}\}$,
  2. (2)
    there exists finitely many vectors $(v_{i},1)\in\mathbb{Z}_{\geq 0}^{n+1}$ generating a semigroup $B\subset\mathbb{Z}_{\geq 0}^{n+1}$, such that $\Gamma\subset B$, and
  3. (3)
    $\Gamma$ contains a set of generators of $\mathbb{Z}^{n+1}$ as a group,
then
$\lim_{m\to\infty}\frac{\#\Gamma_{m}}{m^{n}}=\mathrm{vol}(\Delta(\Gamma)).$
Proposition 3.13.
[Cut14] Let $(R,\mathfrak{m})$ be a Noetherian analytically irreducible local domain of dimension $n$, and $\mathfrak{a}_{\bullet}\in\mathrm{Fil}_{R,\mathfrak{m}}$. Then there exists $t=t(R)\in\mathbb{Z}_{>0}$ depending only on $R$, and $2t$ semigroups $\bar{\Gamma}^{(j)},\Gamma^{(j)}\subset\mathbb{Z}_{\geq 0}^{n+1}$, $j=1,\ldots,t$ satisfying the conditions of Lemma 3.12, such that
$\ell(R/\mathfrak{a}_{m})=\sum_{j=1}^{t}(\#\bar{\Gamma}^{(j)}_{m}-\#\Gamma^{(j)}_{m}).$ (3.13)
Proof.
We include a sketch of the constructions for the reader’s convenience. For more details about the proof, we refer to [Cut14, Section 4].
First note that we may replace $R$ by its $\mathfrak{m}$-adic completion $\hat{R}$ and each $\mathfrak{a}_{\lambda}$ by $\mathfrak{a}_{\lambda}\cdot R$ since $\ell(R/\mathfrak{a}_{\lambda})=\ell(\hat{R}/\mathfrak{a}_{\lambda}\cdot R)$. Thus we may assume that $R$ is complete, and hence excellent. Let $\pi:Z\to\mathrm{Spec}R$ be the normalized blowup of $\mathfrak{m}$. Then there exists a closed point $z\in\pi^{-1}(\mathfrak{m})$ such that $R^{\prime}:=\mathcal{O}_{Z,z}$ is regular since $Z$ is normal. Now $\pi$ is of finite type since $R$ is Nagata. Hence $R^{\prime}$ is essentially of finite type over $R$. Denote $\kappa^{\prime}:=R^{\prime}/\mathfrak{n}$, where $\mathfrak{n}$ is the maximal ideal of $R^{\prime}$, then we know that $t:=[\kappa^{\prime}:\kappa]\in\mathbb{Z}_{>0}$.
We proceed to define a valuation $v$ on the quotient field $K:=Q(R)=Q(R^{\prime})$. Fix a regular system of parameters $(x_{1},\ldots,x_{n})$ of $R^{\prime}$ and $\alpha_{1},\ldots,\alpha_{n}\in\mathbb{R}_{>1}$, which are linearly independent over $\mathbb{Q}$. By Cohen’s structure theorem, the $\mathfrak{n}$-completion $(\widehat{R^{\prime}},\hat{\mathfrak{n}})$ of $R^{\prime}$ has a coefficient ring $C$ as a regular local ring. Hence for any $f\in R^{\prime}$ and $d\in\mathbb{Z}_{>0}$, we can write
$f=\sum_{i_{1}+\ldots i_{n}<d}c_{i_{1},\ldots,i_{n}}x_{1}^{i_{1}}\cdots x_{n}^{i_{n}}+g_{d}\in\widehat{R^{\prime}},$
where $c_{i_{1},\ldots,i_{n}}\in C$ are units and $g_{d}\in\hat{\mathfrak{n}}^{d}$. Take $d_{0}$ such that $c_{I_{1},\ldots,I_{n}}\neq 0$ for some $I_{1}+\ldots+I_{n}<d_{0}$. It is not hard to check that if $d>d_{0}+\sum\alpha_{j}I_{j}$, then the number
$v(f):=\min\{\sum_{j=1}^{n}\alpha_{j}i_{j}\mid c_{i_{1},\ldots,i_{n}}\neq 0\}$
is independent on the choice of the units and $d$, hence this defines a function $v:K^{\times}\to\mathbb{R}$ by $v(f/g)=v(f)-v(g)$ for any $f,g\in R^{\prime}$. One can check that $v$ is a real valuation. Denote its value group by $\Gamma_{v}:=v(K^{\times})$ and its valuation ring by $(\mathcal{O}_{v},\mathfrak{m}_{v})$ and for any $\lambda\in\mathbb{R}_{\geq 0}$, denote
$\mathfrak{b}_{\lambda}:=\{f\in\mathcal{O}_{v}\mid v(f)\geq\lambda\}\text{ and }\mathfrak{b}_{>\lambda}:=\{f\in\mathcal{O}_{v}\mid v(f)>\lambda\}.$
Since $\alpha_{j}$ are linearly independent over $\mathbb{Q}$, we know that
$\mathfrak{b}_{\lambda}/\mathfrak{b}_{>\lambda}=\left\{\begin{aligned} &R^{\prime}/\mathfrak{n}=\kappa^{\prime},&\text{if }\lambda\in\Gamma_{v},\\ &0,&\text{otherwise.}\end{aligned}\right.$
Hence for any $m\in\mathbb{Z}_{>0}$ and $\lambda\in\mathbb{R}_{>0}$, $\dim_{\kappa}\mathfrak{a}_{m}\cap\mathfrak{b}_{\lambda}/\mathfrak{a}_{m}\cap\mathfrak{b}_{>\lambda}\leq\dim_{\kappa}R\cap\mathfrak{b}_{\lambda}/R\cap\mathfrak{b}_{>\lambda}\leq[\kappa^{\prime}:\kappa]=t$.
Fix $c_{1}\in\mathbb{Z}_{>0}$ such that $\mathfrak{m}^{c_{1}}\subset\mathfrak{a}_{1}$. By [Cut13, Lemma 4.3], there exists $c_{2}\in\mathbb{Z}_{>0}$ such that $\mathfrak{b}_{c_{2}m}\cap R\subset\mathfrak{m}^{m}$ for any $m\in\mathbb{Z}_{>0}$. Set $c:=c_{1}c_{2}$. Then for any $m\in\mathbb{Z}_{>0}$ we have
$\mathfrak{b}_{cm}\cap R\subset\mathfrak{m}^{c_{1}m}\subset\mathfrak{a}_{1}^{m}\subset\mathfrak{a}_{m}.$ (3.14)
To simplify the notation, for $\bm{\beta}:=(\beta_{1},\ldots,\beta_{n})\in\mathbb{R}_{\geq 0}^{n}$, we denote $\xi(\bm{\beta}):=\sum_{i=1}^{n}\alpha_{i}\beta_{i}$ and $|\bm{\beta}|:=\sum_{i=1}^{n}\beta_{i}$. Now for $1\leq j\leq t$, define
$\Gamma^{(j)}:=\{(\bm{\beta},m)\in\mathbb{Z}_{\geq 0}^{n+1}\mid\dim_{\kappa}\mathfrak{a}_{m}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m}\cap\mathfrak{b}_{>\xi(\bm{\beta})}\geq j\text{ and }|\bm{\beta}|\leq cm\}$
and
$\bar{\Gamma}^{(j)}:=\{(\bm{\beta},m)\in\mathbb{Z}_{\geq 0}^{n+1}\mid\dim_{\kappa}R\cap\mathfrak{b}_{\xi(\bm{\beta})}/R\cap\mathfrak{b}_{>\xi(\bm{\beta})}\geq j\text{ and }|\bm{\beta}|\leq cm\}.$
One can check that $\Gamma^{(j)}$ and $\bar{\Gamma}^{(j)}$ are semigroups satisfying the conditions of Lemma 3.12.
$\ell(R/\mathfrak{a}_{m})=$ $\ell(R/R\cap\mathfrak{b}_{cm})-\ell(\mathfrak{a}_{m}/\mathfrak{a}_{m}\cap\mathfrak{b}_{cm})$
$=$ $\sum_{0\leq\lambda<cm}\dim_{\kappa}(R\cap\mathfrak{b}_{\lambda}/R\cap\mathfrak{b}_{>\lambda})-\sum_{0\leq\lambda<cm}\dim_{\kappa}(\mathfrak{a}_{m}\cap\mathfrak{b}_{\lambda}/\mathfrak{a}_{m}\cap\mathfrak{b}_{>\lambda})$
$=$ $\sum_{\bm{\beta}\in\mathbb{Z}^{n}_{\geq 0},\xi(\bm{\beta})<cm}\dim_{\kappa}(R\cap\mathfrak{b}_{\xi(\bm{\beta})}/R\cap\mathfrak{b}_{\xi(\bm{\beta})})$
$-\sum_{\bm{\beta}\in\mathbb{Z}^{n}_{\geq 0},\xi(\bm{\beta})<cm}\dim_{\kappa}(\mathfrak{a}_{m}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m}\cap\mathfrak{b}_{\xi(\bm{\beta})})$
$=$ $\sum_{j=1}^{t}\#\{\bm{\beta}\in\mathbb{Z}^{n}_{\geq 0}\mid\dim_{\kappa}(R\cap\mathfrak{b}_{\xi(\bm{\beta})}/R\cap\mathfrak{b}_{>\xi(\bm{\beta})})\geq j,\ \xi(\bm{\beta})<cm\}$
$-\sum_{j=1}^{t}\#\{\bm{\beta}\in\mathbb{Z}^{n}_{\geq 0}\mid\dim_{\kappa}(\mathfrak{a}_{m}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m}\cap\mathfrak{b}_{>\xi(\bm{\beta})})\geq j,\ \xi(\bm{\beta})<cm\}$
$=$ $\sum_{j=1}^{t}\#\bar{\Gamma}^{(j)}_{m}-\sum_{j=1}^{t}\#\Gamma^{(j)}_{m}$
where the first equality follows from (3.14), an the last is because when $|\bm{\beta}|\leq cm$ but $\xi(\bm{\beta})\geq cm$, by (3.14) again we have
$\mathfrak{b}_{\xi(\bm{\beta})}\cap R\subset\mathfrak{b}_{cm}\cap R\subset\mathfrak{a}_{m}.$
This proves (3.13) and finishes the proof. ∎
Lemma 3.14.
Let $\{\Gamma_{k}\subset\mathbb{R}^{n+1}_{\geq 0}\}$ be an increasing sequence of semigroups with respect to inclusion. Then $C(\cup_{k}\Gamma_{k})=\overline{\cup_{k}C(\Gamma_{k})}$.
In particular, $\mathrm{vol}(\Delta(\cup_{k}\Gamma_{k}))=\mathrm{vol}(\cup_{k}\Delta(\Gamma_{k}))=\lim_{k\to\infty}\mathrm{vol}(\Delta(\Gamma_{k}))$.
Proof.
Since $C(\Gamma_{k})\subset C(\cup_{k}\Gamma_{k})$ for any $k\in\mathbb{Z}_{>0}$, we know that $\cup_{k}C(\Gamma_{k})\subset C(\cup_{k}\Gamma_{k})$. So $\overline{\cup_{k}C(\Gamma_{k})}\subset C(\cup_{k}\Gamma_{k})$.
To prove the reverse inclusion, it suffices to note that $\cup_{k}C(\Gamma_{k})$ is a convex cone, which implies $\overline{\cup_{k}C(\Gamma_{k})}=C(\cup_{k}C(\Gamma_{k}))\supset C(\cup_{k}\Gamma_{k})$.
The last equality follows immediately. ∎
Proposition 3.15.
Let $\{\mathfrak{a}_{\bullet,k}\}_{k\in\mathbb{Z}_{>0}}$ be an increasing sequence in $\mathrm{Fil}_{R,\mathfrak{m}}$ such that $\cup\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}_{R,\mathfrak{m}}$. Then
$\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet,k})=0,$
where $\mathfrak{a}_{\bullet}:=\{\cup_{k}\mathfrak{a}_{\lambda,k}\}_{\lambda\in\mathbb{R}_{>0}}$.
Proof.
We construct $2t$ sequence of semigroups $\Gamma_{,k}^{(j)}$ and $\bar{\Gamma}_{,k}^{(j)}$ associated to $\mathfrak{a}_{\bullet,k}$ as in Proposition 3.13. Let $\Gamma^{(j)}$ and $\bar{\Gamma}^{(j)}$ be the semigroups of $\mathfrak{a}_{\bullet}$. By the proof of Proposition 3.13 it is easy to see that $\bar{\Gamma}_{,k}^{(j)}$ is indeed independent on $k$, that is, $\bar{\Gamma}_{,k}^{(j)}=\bar{\Gamma}^{(j)}$ for any $1\leq j\leq t$ and $k\in\mathbb{Z}_{>0}$. We claim that for any $1\leq j\leq t$,
$\Gamma^{(j)}=\cup_{k}\Gamma_{,k}^{(j)}.$ (3.15)
Indeed, by definition $\Gamma_{,k}^{(j)}\subset\Gamma^{(j)}$ for any $k$ and $j$, so $\cup_{k}\Gamma_{,k}^{(j)}\subset\Gamma^{(j)}$. Conversely, for any $(\bm{\beta},m)\in\Gamma^{(j)}$ where $1\leq j\leq t$, we know that
$\dim_{\kappa}\mathfrak{a}_{m}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m}\cap\mathfrak{b}_{>\xi(\bm{\beta})}\geq j.$
Choose $\kappa$-linearly independent elements $\bar{f_{1}},\ldots,\bar{f_{j}}\in\mathfrak{a}_{m}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m}\cap\mathfrak{b}_{>\xi(\bm{\beta})}$ and let $f_{1},\ldots,f_{j}$ be a lifting to $\mathfrak{a}_{m}\cap\mathfrak{b}_{\xi(\bm{\beta})}$. Since $\mathfrak{a}_{m}=\cup_{k}\mathfrak{a}_{m,k}$, there exists $K\in\mathbb{Z}_{>0}$ such that $f_{1},\ldots,f_{j}\in\mathfrak{a}_{m,K}$. Now their image in $\mathfrak{a}_{m,K}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m,K}\cap\mathfrak{b}_{>\xi(\bm{\beta})}$ must be $\kappa$-linearly independent, which implies
$\dim_{\kappa}\mathfrak{a}_{m,K}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m,K}\cap\mathfrak{b}_{>\xi(\bm{\beta})}\geq j.$
This show that $(\bm{\beta},m)\in\Gamma_{,K}^{(j)}$. Thus we get $\Gamma^{(j)}\subset\cup_{k}\Gamma_{,k}^{(j)}$ and finishes the proof of the claim.
Now we can apply Lemma 3.12, Proposition 3.13 and Lemma 3.14 to get
$\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet,k})=$ $\lim_{k\to\infty}(\mathrm{e}(\mathfrak{a}_{\bullet,k})-\mathrm{e}(\mathfrak{a}_{\bullet}))$
$=$ $\lim_{k\to\infty}\lim_{m\to\infty}\frac{\ell(R/\mathfrak{a}_{m,k})-\ell(R/\mathfrak{a}_{m})}{m^{n}/n!}$
$=$ $n!\lim_{k\to\infty}\lim_{m\to\infty}\sum_{j=1}^{t}\frac{(\#\bar{\Gamma}^{(j)})-\#\Gamma_{,k}^{(j)})-(\#\bar{\Gamma}^{(j)}-\#\Gamma^{(j)})}{m^{n}}$
$=$ $n!\sum_{j=1}^{t}\lim_{k\to\infty}(\mathrm{vol}(\Delta(\Gamma^{(j)}))-\mathrm{vol}(\Delta(\Gamma_{,k}^{(j)})$
$=$ $0$
This finishes the proof. ∎
Remark 3.16.
When $x\in X$ is an isolated normal singularity, Proposition 3.15 can be obtained from [BdFF12, Theorem 4.14], since the corresponding $b$-divisors form a decreasing convergent sequence in the pointwise topology.
The following example shows that $\mathrm{e}(\bullet)$ may fail to be continuous along a decreasing sequence of filtrations. However, we will show that continuity along decreasing sequences holds under some geometric conditions in the next section.
Example 3.17.
Let $R=k[\![x]\!]$ with $\mathfrak{m}=(x)$. For $k\in\mathbb{Z}_{>0}$, define $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}$ by
$\mathfrak{a}_{\lambda,k}:=\left\{\begin{aligned} &\mathfrak{m}^{\lceil 2\lambda\rceil},&\lambda\leq k,\\ &\mathfrak{m}^{2k},&k<\lambda\leq 2k,\\ &\mathfrak{m}^{\lceil\lambda\rceil},&\lambda>2k.\end{aligned}\right.$
Then $\mathrm{e}(\mathfrak{a}_{\bullet,k})=\mathrm{e}(\mathfrak{m}^{\bullet})=1$. But since $\cap_{k}\mathfrak{a}_{\bullet,k}=\mathfrak{m}^{2\bullet}$, we have $\mathrm{e}(\cap_{k}\mathfrak{a}_{\bullet,k})=\mathrm{e}(\mathfrak{m}^{2\bullet})=2\mathrm{e}(\mathfrak{m}^{\bullet})=2$.
As an application of Proposition 3.15, we prove a convergence result which might be of its own interest.
Proposition 3.18.
Let $x\in X$ be a klt singularity over a field of characteristic 00. Let $v_{k}\in\mathrm{Val}_{X,x}$, $k\in\mathbb{Z}_{>0}$ be a sequence of valuations such that $A_{X}(v_{k})=1$ and $\widehat{\mathrm{vol}}_{x,X}(v_{k})<V$ for some $V>0$. Possibly passing to a subsequence, there exists an increasing sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}^{s}$ with $\mathfrak{a}_{\lambda,k}\subset\mathfrak{a}_{\lambda}(v_{k})$, which converges to some $v\in\mathrm{Val}_{X,x}$ both weakly and with respect to the $d_{1}$-metric.
Proof.
By [Li18, Theorem 1.1], there exists $\delta>0$ such that $v_{k}(\mathfrak{m})>\delta$ for any $k$. By [LX20, Proposition 3.9],the set
$\{v\in\mathrm{Val}_{X,x}\mid v(\mathfrak{m})\geq\delta,\ A_{X}(v)\leq 1\}$
is sequentially compact under the weak topology. Hence there is a subsequence of $\{v_{k}\}$ which converges to some $v\in\mathrm{Val}_{X,x}$. From now on, we assume that $v_{k}\to v$ weakly.
For $k,l\in\mathbb{Z}_{>0}$, define
$\mathfrak{a}_{\bullet,k}^{(l)}:=\bigcap_{j=k}^{k+l}\mathfrak{a}_{\bullet}(v_{j}).$
Then for any $k\in\mathbb{Z}_{>0}$, $\{\mathfrak{a}_{\bullet,k}^{(l)}\}_{l\in\mathbb{Z}_{>0}}$ is a decreasing sequence of filtrations. Let
$\mathfrak{a}_{\bullet,k}:=\cap_{l}\mathfrak{a}_{\bullet,k}^{(l)}\subset\mathfrak{a}_{\bullet}(v_{k})$
Clearly the sequence of filtrations $\{\mathfrak{a}_{\bullet,k}\}_{k\in\mathbb{Z}_{>0}}$ is increasing. Let $\chi_{k}=\mathrm{ord}_{\mathfrak{a}_{\bullet,k}}$ be the corresponding norm. By definition, $\chi_{k}=\inf_{j\geq k}v_{j}\leq v$. Since $v_{j}\to v$ weakly, given any $f\in\mathfrak{m}$ and any $\epsilon>0$, there exists $J\in\mathbb{Z}_{>0}$ such that $v_{j}(f)>v(f)-\epsilon$ for any $j\geq J$. Thus
$v(f)\geq\chi_{k}(f)=\inf_{j\geq k}v_{j}(f)\geq v(f)-\epsilon$
as long as $k\geq J$. This proves $\chi_{k}\to v$ weakly as an increasing sequence. So we may apply Proposition 3.15 to get $\lim_{k\to\infty}d_{1}(\chi_{k},v)=0$, and the proof is finished. ∎
3.3.2. Normalized volumes and continuity along decreasing sequences
In this section, we assume that $R$ contains an algebraically closed field $\mathbbm{k}$ with $\mathrm{char}\mathbbm{k}=0$. As usual, we denote the singularity by $(x\in X):=(\mathfrak{m}\in\mathrm{Spec}R)$. The result is not used in the proof of the main results, though it might be of its own interest.
Proposition 3.19.
Let $x\in X$ be a klt singularity. Let $\{\mathfrak{a}_{\bullet,k}\}_{k\in\mathbb{Z}_{>0}}$ be a decreasing sequence in $\mathrm{Fil}$. Assume each $\mathfrak{a}_{\bullet,k}$ is of the form $\mathfrak{a}_{\bullet,k}=\cap_{i\in I_{k}}\mathfrak{a}_{\bullet}(v_{i})$, where for each $k$, $\{v_{i}\}_{i\in I_{k}}\subset\mathrm{Val}^{+}_{X,x}$ is a set of valuations satisfying the following conditions.
For some constants $V>0$ and $C\in\mathbb{Z}_{>0}$,
  1. (1)
    for any $k$ and $i\in I_{k}$, $\widehat{\mathrm{vol}}_{x,X}(v_{i})<V$,
  2. (2)
    for any $k$ and $i\in I_{k}$, $v_{i}(\mathfrak{m})>1/C$, and
  3. (3)
    for any $k$, there exists $i_{0}\in I_{k}$ such that $\mathfrak{a}_{C\bullet}(v_{i_{0}})\subset\mathfrak{m}^{\bullet}$.
Then
$\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet,k})=0,$
where $\mathfrak{a}_{\bullet}:=\cap_{k}\mathfrak{a}_{\bullet,k}$.
Remark 3.20.
If $I_{k}$ has a common element $v_{0}\in\mathrm{Val}^{+}_{X,x}$, then condition (3) in the proposition is automatic, since by Lemma 2.13, $\mathfrak{a}_{\bullet}(v_{0})$ is linearly bounded.
Proof.
By construction, $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$ and $\mathrm{e}(\mathfrak{a}_{\bullet})\geq\mathrm{e}(\mathfrak{a}_{\bullet,k})$ for any $k$. So $\lim_{k}\mathrm{e}(\mathfrak{a}_{\bullet,k})\in\mathbb{R}_{>0}$.
For any fixed $m\in\mathbb{Z}{{}_{>}0}$, we have $\cap_{k}\mathfrak{a}_{m,k}=\mathfrak{a}_{m}$. So actually there exists $K_{m}\in\mathbb{Z}_{>0}$ such that for any $k\geq K_{m}$, $\mathfrak{a}_{m,k}=\mathfrak{a}_{m}$, since $R/\mathfrak{a}_{m}$ is Artinian. In particular, $\mathrm{e}(\mathfrak{a}_{m,k})=\mathrm{e}(\mathfrak{a}_{m})$. Now for any $\epsilon>0$, by Lemma 3.21, there exists $M=M(C,V,\epsilon)>0$ such that for any $k$ and any $m\geq M$
$\mathrm{e}(\mathfrak{a}_{\bullet,k})\leq\frac{\mathrm{e}(\mathfrak{a}_{m,k})}{m^{n}}\leq\mathrm{e}(\mathfrak{a}_{\bullet,k})+\epsilon.$
Thus for $m\geq M$ and $k\geq K_{m}$, we have
$\mathrm{e}(\mathfrak{a}_{\bullet,k})\leq\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}}\leq\mathrm{e}(\mathfrak{a}_{\bullet,k})+\epsilon.$
Letting $k\to\infty$, for $m\geq M$ we get
$\lim_{k}\mathrm{e}(\mathfrak{a}_{\bullet,k})\leq\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}}\leq\lim_{k}\mathrm{e}(\mathfrak{a}_{\bullet,k})+\epsilon.$
Taking limit with respect to $m$ and using (2.3), we get $\lim_{k}\mathrm{e}(\mathfrak{a}_{\bullet,k})\leq\mathrm{e}(\mathfrak{a}_{\bullet})\leq\lim_{k}\mathrm{e}(\mathfrak{a}_{\bullet,k})+\epsilon$. Since $\epsilon>0$ is arbitrary, we get $\mathrm{e}(\mathfrak{a}_{\bullet})=\lim_{k}\mathrm{e}(\mathfrak{a}_{\bullet,k})$, which finishes the proof. ∎
In the proof of the proposition, we need the following lemma, which is a slight generalization of the uniform approximation proved in [Blu18]. For the definition of multiplier ideals in the proof below, we also refer to ibid.
Lemma 3.21.
(c.f. [Blu18, Proposition 3.7]) Let $\epsilon,V>0$ and $C\in\mathbb{Z}_{>1}$ be constants. Then for there exists $\Lambda=\Lambda(C,V,\epsilon)\in\mathbb{R}_{>0}$ such that the following holds.
Let $\mathfrak{a}_{\bullet}=\cap_{i\in I}\mathfrak{a}_{\bullet}(v_{i})$, where $\{v_{i}\}_{i\in I}\subset\mathrm{Val}^{+}_{X,x}$ is a set of valuations such that
  1. (1)
    $\widehat{\mathrm{vol}}_{x,X}(v_{i})<V$ for any $i\in I$,
  2. (2)
    $v_{i}(\mathfrak{m})>1/C$ for any $i\in I$, and
  3. (3)
    $\mathfrak{a}_{C\bullet}(v_{i_{0}})\subset\mathfrak{m}^{\bullet}$ for some $i_{0}\in I$.
Then for any $\lambda>\Lambda$, we have
$\mathrm{e}(\mathfrak{a}_{\bullet})\leq\frac{\mathrm{e}(\mathfrak{a}_{\lambda})}{\lambda^{n}}<\mathrm{e}(\mathfrak{a}_{\bullet})+\epsilon.$
Proof.
Under the conditions, we have
$\mathrm{lct}(X;\mathfrak{a}_{\bullet})\leq C\cdot\mathrm{lct}(X;\mathfrak{m}).$
Hence by Lemma 2.34 for any $i\in I$,
$\frac{A_{X}(v_{i})}{v_{i}(\mathfrak{a}_{\bullet})}\leq c_{0}\cdot\frac{\widehat{\mathrm{vol}}_{x,X}(v_{i})}{\widehat{\mathrm{vol}}(x,X)}\cdot\mathrm{lct}(X;\mathfrak{a}_{\bullet})\leq\frac{c_{0}C\cdot V\cdot\mathrm{lct}(X;\mathfrak{m})}{\widehat{\mathrm{vol}}(x,X)}.$
By the valuative characterization of multiplier ideals [BdFFU15, Theorem 1.2], we have
$\mathcal{J}(X;\lambda\cdot\mathfrak{a}_{\bullet})\subset\mathfrak{a}_{\lambda-E},$ (3.16)
where $E:=\frac{c_{0}CV\mathrm{lct}(X;\mathfrak{m})}{\widehat{\mathrm{vol}}(x,X)}>0$ depends only on $C$ and $V$. Replacing the inclusion
$\mathcal{J}(X;m\cdot\mathfrak{a}_{\bullet}(v))\subset\mathfrak{a}_{m-E}$
in the proof of [Blu18, Proposition 3.7], which follows from the assumption $A(v)\leq E$, by (3.16), the proposition follows from the same argument therein. ∎
3.4. The toric case and Newton-Okounkov bodies
In this section, we work over a field $\mathbbm{k}$ and follow the notation in [Ful93] for toric varieties. Our goal is to identify the subspace $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ of saturated monomial filtrations on a toric singularity with a subspace of the Fréchet-Nikodym-Aronszyan metric on cobounded sets of the dual cone. See [KK14, Section 6] and [Blu18, Section 8] for some relevant constructions.
Let $N$ be a free abelian group of rank $n\geq 1$ and $M=N^{*}$ its dual. Let $\sigma\subset N_{\mathbb{R}}=N\otimes\mathbb{R}$ be a strongly convex rational polyhedral cone of maximal dimension. We get an affine toric variety $X_{\sigma}=\mathrm{Spec}R_{\sigma}=\mathrm{Spec}\mathbbm{k}[\sigma^{\vee}\cap M]$ with a unique torus invariant point $x$, where $\sigma^{\vee}\subset M_{\mathbb{R}}=M\otimes\mathbb{R}$ is the dual cone of $\sigma$. Let $R$ be the local ring of $X_{\sigma}$ at $x$ and $\mathfrak{m}$ its maximal ideal.
Recall that there is a $1$-to-$1$ correspondence between toric valuations $v_{u}\in\mathrm{Val}^{\mathrm{toric}}_{R}$ and $u\in\sigma$, given by
$v_{u}(\sum_{m\in\sigma^{\vee}\cap M}c_{m}\chi^{m})=\min\{\langle u,m\rangle\mid c_{m}\neq 0\},$
and $v_{u}$ is centered at $\mathfrak{m}$ if and only if $u\in\mathrm{Int}(\sigma)$. For $u\in\mathrm{Int}(\sigma)$ and $\lambda\in\mathbb{R}_{>0}$, denote $H_{u}(\lambda)\coloneqq\{\beta\in M_{\mathbb{R}}\mid\langle\beta,u\rangle\geq\lambda\}$. Then we have
$\mathfrak{a}_{\lambda}(v_{u})=\mathrm{span}\{\chi^{m}\mid m\in H_{u}(\lambda)\cap\sigma^{\vee}\cap M\},$
and
$\mathrm{vol}(v_{u})=n!\cdot\mathrm{vol}(\sigma^{\vee}\backslash H_{u}(1)).$
We now reproduce the generalization of the above construction to the case of monomial filtrations following [KK14], see also [Mus02]. An $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ on $R$ is called monomial if for any $\lambda\in\mathbb{R}_{>0}$, $\mathfrak{a}_{\lambda}$ is a monomial ideal. Denote the set of (resp. saturated) linearly bounded monomial $\mathfrak{m}$-filtrations by $\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$ (resp. $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$).
Given an $\mathfrak{m}$-primary monomial ideal $\mathfrak{a}$, the Newton polyhedron $P(\mathfrak{a})$ of $\mathfrak{a}$ is the convex hull of $\{m\in\sigma^{\vee}\cap M\mid\chi^{m}\in\mathfrak{a}\}$. Moreover, we have $\mathrm{e}(\mathfrak{a})=\mathrm{vol}(\sigma^{\vee}\backslash P(\mathfrak{a}))$. Recall that the Newton-Okounkov body $P(\mathfrak{a}_{\bullet})$ of a monomial $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$ is defined to be
$P(\mathfrak{a}_{\bullet})\coloneqq\overline{\cup_{\lambda\in\mathbb{R}_{>0}}\{m/\lambda\mid m\in P(\mathfrak{a}_{\lambda})\}}\subset\sigma^{\vee},$
which is a convex set such that $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})$ is bounded. It is known that the covolume of the Newton-Okounkov body computes the multiplicity of the filtration.
Lemma 3.22.
[KK14, Theorem 6.5] Let $\mathfrak{a}_{\bullet}$ be a monomial $\mathfrak{m}$-filtration on $R$. Then we have
$\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{vol}(\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})).$
The next lemma gives an alternative characterization for $P(\mathfrak{a}_{\bullet})$.
Lemma 3.23.
For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$, $P(\mathfrak{a}_{\bullet})=\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
Proof.
For $m\in\sigma^{\vee}\cap M$ with $\chi^{m}\in\mathfrak{a}_{\lambda}$, we have $\langle u,m\rangle=v_{u}(m)\geq v_{u}(\mathfrak{a}_{\lambda})\geq\lambda v(\mathfrak{a}_{\bullet})$ by definition, that is, $m/\lambda\in H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$. Hence $P(\mathfrak{a}_{\bullet})\subset\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
Conversely, assume $p\in\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$. Assume $p\notin P(\mathfrak{a}_{\bullet})$. Since $P(\mathfrak{a}_{\bullet})$ is convex, there exists $u\in N_{\mathbb{R}}$ and $b\in\mathbb{R}_{>0}$ such that $\langle u,p\rangle<b$ but $P(\mathfrak{a}_{\bullet})\subset H_{u}(b)$. Since $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})$ is bounded, we know that $u\in\mathrm{Int}(\sigma)$. Now $b>\langle u,p\rangle\geq v_{u}(\mathfrak{a}_{\bullet})$ by the choice of $p$. Hence there exists $\lambda\in\mathbb{R}_{>0}$ and $\chi^{m}\in\mathfrak{a}_{\lambda}$ such that $v_{u}(\chi^{m})=v_{u}(\mathfrak{a}_{\lambda})<\lambda\cdot b$. In particular, $m/\lambda\in P(\mathfrak{a}_{\bullet})\backslash H_{u}(b)$, a contradiction. So $\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))\subset P(\mathfrak{a}_{\bullet})$ and the proof is finished. ∎
Next we show that the saturation of a monomial filtration can be computed using only toric valuations.
Lemma 3.24.
A monomial filtration $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$ is saturated if and only if it is of the form
$\mathfrak{a}_{\bullet}=\cap_{u\in\mathrm{Int}(\sigma)}\mathfrak{a}_{\bullet}(v_{u}/a_{u})$ (3.17)
where $a_{u}:\mathrm{Int}(\sigma)\to\mathbb{R}_{>0}$ is a homogeneous function.
Proof.
If $\mathfrak{a}_{\bullet}$ is of the form (3.17), then $\mathfrak{a}_{\bullet}$ is saturated by Proposition 2.20 since $v_{u}\in\mathrm{Val}^{+}_{R,\mathfrak{m}}$.
Conversely, assume that $\mathfrak{a}_{\bullet}$ is saturated. Then we know that
$\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\coloneqq\cap_{u\in\mathrm{Int}(\sigma)}\mathfrak{a}_{\bullet}(\frac{v_{u}}{v_{u}(\mathfrak{a}_{\bullet})})$
by Lemma 2.19. Moreover, $v_{u}(\mathfrak{a}_{\bullet}^{\prime})=v_{u}(\mathfrak{a}_{\bullet})$ since $\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\subset\mathfrak{a}_{\bullet}(v_{u}/v_{u}(\mathfrak{a}_{\bullet}))$. So $P(\mathfrak{a}_{\bullet})=P(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.23, and hence $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.22. Now $\mathfrak{a}^{\prime}_{\bullet}$ is saturated by Proposition 2.20 again, so by Theorem 2.18, we get $\mathfrak{a}_{\bullet}=\mathfrak{a}^{\prime}_{\bullet}$. ∎
Now we can characterize the intersection of two monomial filtrations.
Lemma 3.25.
Let $I$ be a countable index set and $\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}^{s,\mathrm{mon}}_{X,x}$ for $i\in I$ and assume that $\cap_{i\in I}\mathfrak{a}_{\bullet,i}\in\mathrm{Fil}^{s,\mathrm{mon}}_{X,x}$. Then
$v_{u}(\cap_{i\in I}\mathfrak{a}_{\bullet,i})=\sup_{i\in I}v_{u}(\mathfrak{a}_{\bullet,i})$ (3.18)
for any $u\in\mathrm{Int}(\sigma)$. As a result, we have $P(\cap_{i\in I}\mathfrak{a}_{\bullet,i})=\cap_{i\in I}P(\mathfrak{a}_{\bullet,i})$.
Proof.
Since $\cap_{i}\mathfrak{a}_{\bullet,i}\subset\mathfrak{a}_{\bullet,i}$ for any $i$, we have $v(\cap_{i}\mathfrak{a}_{\bullet,i})\geq\sup_{i\in I}v(\mathfrak{a}_{\bullet,i})$ for any $v\in\mathrm{Val}_{X,x}$.
To prove the reverse inequality for toric valuations, we first treat the case where $I$ is finite. It suffices to prove $v_{u}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})=\max\{v_{u}(\mathfrak{a}_{\bullet}),v_{u}(\mathfrak{b}_{\bullet})\}$ for any $u\in\mathrm{Int}(\sigma)$ and apply induction. Denote $\mathfrak{c}_{\bullet}\coloneqq\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$, and $a_{u}\coloneqq v_{u}(\mathfrak{a}_{\bullet})$, $b_{u}\coloneqq v_{u}(\mathfrak{b}_{\bullet})$, $c_{u}\coloneqq v_{u}(\mathfrak{c}_{\bullet})$ and $c^{\prime}_{u}\coloneqq\max\{a_{u},b_{u}\}$. Assume to the contrary that there exists $w\in\mathrm{Int}(\sigma)$ such that $c_{w}>c^{\prime}_{w}$. Fix $\delta\in\mathbb{R}_{>0}$ such that $c_{w}(1-\delta)>c^{\prime}_{w}$. Choose $p\in P(\mathfrak{a}_{\bullet})\cap P(\mathfrak{b}_{\bullet})$ such that $\langle p,w\rangle=\min_{q\in P(\mathfrak{a}_{\bullet})\cap P(\mathfrak{b}_{\bullet})}\langle q,w\rangle$. Let
$C\coloneqq\{p\}+(\sigma^{\vee}\backslash H_{w}(\delta c^{\prime}_{w}))\subset\sigma^{\vee},$
where $+$ denotes the Minkowski sum. Take $p_{i}\in C\cap M_{\mathbb{Q}}$ such that $p_{i}\to p$ and $k_{i}\in\mathbb{Z}_{>0}$ such that $k_{i}p_{i}\in M$. By Lemma 3.23, for any $u\in\mathrm{Int}(\sigma)$, $\langle u,p_{i}\rangle\geq\langle u,p\rangle\geq c^{\prime}_{u}$. Hence $\chi^{k_{i}p_{i}}\in\mathfrak{a}_{k_{i}}\cap\mathfrak{b}_{k_{i}}$by Lemma 3.24, which implies
$v_{w}(\mathfrak{a}_{k_{i}}\cap\mathfrak{b}_{k_{i}})/k_{i}\leq\langle w,p_{i}\rangle<\langle w,p\rangle+\delta c^{\prime}_{w}=(1+\delta)c^{\prime}_{w},$
where the last equality follows from Lemma 3.23. Letting $k_{i}\to\infty$ we get $c_{w}\leq(1+\delta)c^{\prime}_{w}$. Thus we get $c^{\prime}_{w}<(1-\delta)c_{w}<(1-\delta^{2})c^{\prime}_{w}$, a contradiction.
The proof for the general case is quite similar. For simplicity let $I=\mathbb{Z}_{>0}$. By the case where $I$ is finite, we may replace $\mathfrak{a}_{\bullet,i}$ by $\cap_{j\leq i}\mathfrak{a}_{\bullet,j}$ and assume that $\mathfrak{a}_{\bullet,i}$ is decreasing. Assume to the contrary that there exists $w\in\mathrm{Int}(\sigma)$ such that $a_{w}>\sup_{i\in I}a_{w,i}$. Fix $\delta\in\mathbb{R}_{>0}$ such that $a_{w}(1-4\delta)>\sup_{i}a_{w,i}$. For each $i\in I$, there exists $\lambda_{i}\in\mathbb{R}_{>0}$ and $\chi^{m_{i}}\in\mathfrak{a}_{\lambda_{i},i}$ such that
$\lambda_{i}\cdot a_{w,i}\leq v_{w}(\chi^{m_{i}})=\langle w,m_{i}\rangle<\lambda_{i}\cdot a_{w}(1-3\delta),$
where the first inequality follows from the assumption that $\chi^{m_{i}}\in\mathfrak{a}_{\lambda_{i},i}$ and the second inequality follows from the definition $v_{w}(\mathfrak{a}_{\bullet,i})=\lim_{\lambda}v(\mathfrak{a}_{\lambda,i})/\lambda$. In particular, $m_{i}/\lambda_{i}$ lies in a bounded region and hence after passing to a subsequence, we may assume that $m_{i}/\lambda_{i}\to p\in\sigma^{\vee}\backslash H_{u}(a_{u}(1-2\delta))$. Now consider the region
$C\coloneqq\{p\}+(\sigma^{\vee}\backslash H_{w}(\delta a_{w}))\subset\sigma^{\vee}$
where $+$ denotes the Minkowski sum. Then there exists $p_{j}\in C\cap M_{\mathbb{Q}}$ such that $p_{j}\to p$. Take $k_{j}\in\mathbb{Z}_{>0}$ such that $k_{j}p_{j}\in M$. Then for any $u\in\mathrm{Int}(\sigma)$, we have $\langle u,p\rangle\leq\langle u,p_{j}\rangle$ for any $u\in\mathrm{Int}(\sigma)$ and
$\langle w,p_{j}\rangle<a_{w}(1-2\delta)+\delta a_{w}=a_{w}(1-\delta).$ (3.19)
By the first inequality, for any $u\in\mathrm{Int}(\sigma)$ we have
$\langle u,k_{j}p_{j}\rangle/k_{j}\geq\langle u,p\rangle=\lim_{i}\langle u,m_{i}/\lambda_{i}\rangle\geq\sup_{i}a_{u,i},$
where in the last inequality we used the fact that $\{\mathfrak{a}_{\bullet,i}\}$ is decreasing. Hence $\chi^{k_{j}p_{j}}\in\cap_{i}\mathfrak{a}_{k_{j},i}$ by Lemma 3.24, which implies $\langle u,p_{j}\rangle=\langle u,k_{j}p_{j}\rangle/k_{j}\geq a_{u}$ for any $u\in\mathrm{Int}(\sigma)$. In particular, we get
$a_{w}\leq\langle w,p_{j}\rangle<a_{w}-\delta,$
a contradiction. This contradiction proves (3.18) and the last assertion follows from (3.18) and Lemma 3.23. ∎
Let $\mathcal{P}(\sigma^{\vee})\coloneqq\{P\subsetneq\sigma^{\vee}\mid P\text{ is closed and convex, }\sigma^{\vee}\backslash P\text{ is bounded}\}$. We prove that taking Newton-Okounkov body gives a bijection between $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ and $\mathcal{P}(\sigma^{\vee})$.
Lemma 3.26.
$P:\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}\to\mathcal{P}(\sigma^{\vee})$ is a bijection, whose inverse is given by $P\mapsto\mathfrak{a}_{\bullet}(P)$, where
$\mathfrak{a}_{\lambda}(P)\coloneqq\mathrm{span}\{\chi^{m},m\in\sigma^{\vee}\cap M\mid\langle m,u\rangle\geq\lambda\cdot\inf_{p\in P}\langle p,u\rangle\text{ for any }u\in\mathrm{Int}(\sigma)\},$ (3.20)
Proof.
Let $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s,\mathrm{mon}}_{X,x}$. It is known that $P(\mathfrak{a}_{\bullet})$ is convex (e.g. [KK14]). Fix $u\in\mathrm{Int}(\sigma)$. Since $\mathfrak{a}_{\bullet}$ is linearly bounded, by definition there exists $c^{\prime}\in\mathbb{R}_{>0}$ such that $\mathfrak{a}_{\bullet}\subset\mathfrak{m}^{\bullet/c^{\prime}}$, which implies $v_{u}(\mathfrak{a}_{\bullet})\geq v_{u}(\mathfrak{m}^{\bullet/c^{\prime}})=c^{\prime}v_{u}(\mathfrak{m})\eqqcolon c>0$. By Lemma 3.24, we get $P(\mathfrak{a}_{\bullet})\subset H_{u}(c)\subsetneq\sigma^{\vee}$. Fix $d\in\mathbb{Z}_{>0}$ such that $\mathfrak{m}^{d}\subset\mathfrak{a}_{1}$. Then for any $m\in\mathbb{Z}_{>0}$ we have $\mathfrak{m}^{dm}\subset\mathfrak{a}_{m}$. By Proposition 2.27, there exist finitely many $u_{i}\in\mathrm{Int}(\sigma)$, $i=1,\ldots,k$ such that $\widetilde{\mathfrak{m}^{d\bullet}}=\cap_{i=1}^{k}\mathfrak{a}_{\bullet}(v_{u_{i}})$. By Lemma 3.23 and Lemma 3.25, we have
$P(\mathfrak{a}_{\bullet})\supset P(\mathfrak{m}^{d\bullet)}=P(\widetilde{\mathfrak{m}^{d\bullet}})=\cap_{i=1}^{k}P(\mathfrak{a}_{\bullet}(v_{u_{i}})=\cap_{i=1}^{k}H_{u_{i}}(1).$
Thus $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})\subset\cup_{i=1}^{k}\sigma^{\vee}\backslash H_{u_{i}}(1)$, which is bounded. This proves $P:\mathrm{Fil}^{s,\mathrm{mon}}_{X,x}\to\mathcal{P}(\sigma^{\vee})$.
Now let $P\in\mathcal{P}(\sigma^{\vee})$. We need to prove that $\mathfrak{a}_{\bullet}(P)$ is a linearly bounded monomial $\mathfrak{m}$-filtration. It is easy to see from (3.20) that $\mathfrak{a}_{\lambda}(P)$ is an $\mathfrak{m}$-primary monomial ideal. Condition (1) of Definition 2.1 follows immediately from (3.20) and (3) follows fron the linearity of the function $\langle\bullet,u\rangle$. To prove (2), it suffices to note that $\sigma^{\vee}\cap M$ is discrete, and hence for any $\lambda\in\mathbb{R}_{>0}$ and $u\in\mathrm{Int}(\sigma)$, there exists $\epsilon=\epsilon(\lambda,u)$ such that $H_{u}(\lambda-\epsilon)\backslash H_{u}(\lambda)\cap\sigma^{\vee}\cap M=\emptyset$.
It is then routine to check the four inclusions. We include the proof for the reader’s convenience.
$\mathfrak{a}_{\bullet}\subset\mathfrak{a}_{\bullet}(P(\mathfrak{a}_{\bullet}))$. If $\chi^{m}\in\mathfrak{a}_{\lambda}$, then $m/\lambda\in P(\mathfrak{a}_{\bullet})$, and hence $\chi^{m}\in\mathfrak{a}_{\lambda}(P(\mathfrak{a}_{\bullet}))$ by (3.20).
$\mathfrak{a}_{\bullet}(P(\mathfrak{a}_{\bullet}))\subset\mathfrak{a}_{\bullet}$. By Lemma 3.23, $\langle p,u\rangle\geq v_{u}(\mathfrak{a}_{\bullet})$ for any $p\in P(\mathfrak{a}_{\bullet})$ and $u\in\mathrm{Int}(\sigma)$. Hence if $\chi^{m}\in\mathfrak{a}_{\lambda}(P(\mathfrak{a}_{\bullet}))$, then
$\langle m,u\rangle\geq\lambda\cdot\inf_{p\in P(\mathfrak{a}_{\bullet})}\langle p,u\rangle\geq\lambda v_{u}(\mathfrak{a}_{\bullet})$
for any $u\in\mathrm{Int}(\sigma)$. By Lemma 3.24, we know $\chi^{m}\in\mathfrak{a}_{\lambda}$.
$P\subset P(\mathfrak{a}_{\bullet}(P))$. For $p\in P$, take $p_{i}\in P\cap M_{\mathbb{Q}}$ such that $p_{i}\to p$ and $k_{i}\in\mathbb{Z}_{>0}$ such that $m_{i}\coloneqq k_{i}p_{i}\in P\cap M$. Then by (3.20), $\chi^{m_{i}}\in\mathfrak{a}_{\bullet}(P)$, and hence $p=\lim_{i}m_{i}/k_{i}\in P(\mathfrak{a}_{\bullet}(P))$.
$P(\mathfrak{a}_{\bullet}(P))\subset P$. If $p\in P(\mathfrak{a}_{\bullet}(P))$, then by definition, there exists $m_{i}\in\sigma^{\vee}\cap M$ and $\lambda\in\mathbb{R}_{>0}$ such that $\chi^{m_{i}}\in\mathfrak{a}_{\lambda_{i}}$ and $m_{i}/\lambda_{i}\to p$. By (3.20), this implies $\langle p,u\rangle\geq\inf_{q\in P}\langle q,u\rangle$ for any $u\in\mathrm{Int}(\sigma)$. Applying the same argument as in the proof of Lemma 3.23, we know that $p\in P$. ∎
Recall that the symmetric difference metric (also called the Fréchet-Nikodym-Aronszayn distance) $d$ on $\mathcal{P}(\sigma^{\vee})$ is defined by
$d(P_{1},P_{2})\coloneqq\mathrm{vol}(P_{1}\Delta P_{2}),$
where $P_{1}\Delta P_{2}=P_{1}\backslash P_{2}\cup P_{2}\backslash P_{1}$ is the symmetric difference. We are now ready to prove
Proposition 3.27.
The map $P:(\mathrm{Fil}^{s}_{R,\mathfrak{m}},d_{1})\to(P(\sigma^{\vee},d)$, $\mathfrak{a}_{\bullet}\mapsto P(\mathfrak{a}_{\bullet})$ is an isometry.
Proof.
By Lemma 3.26, $P$ is a bijection.
For simplicity, denote $\mathrm{covol}(P)\coloneqq\mathrm{vol}(\sigma^{\vee}\backslash P)$ below. For $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, we have
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=$ $2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet})$
$=$ $2\mathrm{covol}(P(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}))-\mathrm{covol}(P(\mathfrak{a}_{\bullet}))-\mathrm{covol}(P(\mathfrak{b}_{\bullet}))$
$=$ $2\mathrm{covol}(P(\mathfrak{a}_{\bullet})\cap P(\mathfrak{b}_{\bullet}))-\mathrm{covol}(P(\mathfrak{a}_{\bullet}))-\mathrm{covol}(P(\mathfrak{b}_{\bullet}))$
$=$ $\mathrm{vol}(P(\mathfrak{a}_{\bullet})\backslash P(\mathfrak{b}_{\bullet}))+\mathrm{vol}(P(\mathfrak{b}_{\bullet})\backslash P(\mathfrak{a}_{\bullet}))$
$=$ $\mathrm{vol}(P(\mathfrak{a}_{\bullet})\Delta P(\mathfrak{b}_{\bullet}))=d(P(\mathfrak{a}_{\bullet}),P(\mathfrak{b}_{\bullet})),$
where the second equality follows from Lemma 3.23 and the third from 3.25. The proof is finished. ∎
4. Supnorm metrics on the space of filtrations
4.1. Supnorm metrics and homogenization
Throughout this section we will use the language of norms.
Definition 4.1.
Fix a norm $\rho\in\mathcal{N}^{h}$. We define a function $d_{\infty,\rho}:\mathcal{N}\times\mathcal{N}\to\mathbb{R}_{\geq 0}$ by
$d_{\infty,\rho}(\chi,\chi^{\prime}):=\limsup_{\lambda\to\infty}\sup_{\rho(f)\geq\lambda}\left\{\frac{|\chi(f)-\chi^{\prime}(f)|}{\rho(f)}\right\}.$
We will denote $d_{\infty,\chi_{0}}=:d_{\infty}$.
Lemma 4.2.
For any $\rho\in\mathcal{N}$, $d_{\infty,\rho}$ defines a pseudometric on $\mathcal{N}$. Moreover, for any $\rho,\rho^{\prime}\in\mathcal{N}$, there exists $M\in\mathbb{R}_{>0}$ such that $M^{-1}d_{\infty,\rho}\leq d_{\infty,\rho^{\prime}}\leq Md_{\infty,\rho}$.
Proof.
By definition, $d_{\infty,\rho}$ is symmetric and non-negative for any $\rho$. Given $\chi,\chi^{\prime},\chi^{\prime\prime}\in\mathcal{N}$, since $|\chi(f)-\chi^{\prime\prime}(f)|\leq|\chi(f)-\chi^{\prime}(f)|+|\chi^{\prime}(f)+\chi^{\prime\prime}(f)|$, the triangle inequality holds. This proves that $d_{\infty,\rho}$ is a pseudometric.
The second claim follows from the fact that given $\rho,\rho^{\prime}\in\mathcal{N}$, there exists $M\in\mathbb{R}_{>0}$ such that $M^{-1}\rho\leq\rho^{\prime}\leq M\rho$. ∎
Lemma 4.3.
If $\chi,\chi^{\prime}\in\mathcal{N}^{h}$, then
$d_{\infty}(\chi,\chi^{\prime})=\sup_{f\in\mathfrak{m}}\frac{|\chi(f)-\chi(f^{\prime})|}{\chi_{0}(f)}.$
Proof.
For any $\epsilon>0$, take $g\in\mathfrak{m}$ such that $\frac{|\chi(g)-\chi^{\prime}(g)|}{\chi_{0}(g)}>\sup_{f\in\mathfrak{m}}\frac{|\chi(f)-\chi(f^{\prime})|}{\chi_{0}(f)}-\epsilon\eqqcolon d-\epsilon$. Denote $a:=\chi_{0}(g)>0$. Since $\chi$, $\chi^{\prime}$ and $\chi_{0}$ are homogeneous, we have
$\sup_{\chi_{0}(f)=ka}\frac{|\chi(f)-\chi(f^{\prime})|}{\chi_{0}(f)}\geq$ $\frac{|\chi(g^{k})-\chi^{\prime}(g^{k})|}{\chi_{0}(g^{k})}=\frac{|\chi(g)-\chi^{\prime}(g)|}{\chi_{0}(g)}>d-\epsilon.$
So
$d_{\infty}(\chi,\chi^{\prime})=\limsup_{\lambda\to\infty}\sup_{\chi_{0}(f)\geq\lambda}\frac{|\chi(f)-\chi(f^{\prime})|}{\chi_{0}(f)}\geq d-\epsilon.$
Thus $d_{\infty}(\chi.\chi^{\prime})\geq d$. Since the reverse inequality is obvious, this finishes the proof. ∎
Lemma 4.4.
Let $\chi,\chi^{\prime}\in\mathcal{N}$, then $d_{\infty}(\widehat{\chi},\widehat{\chi}^{\prime})\leq d_{\infty}(\chi,\chi^{\prime})$.
Proof.
Denote $d:=d_{\infty}(\widehat{\chi},\widehat{\chi}^{\prime})$. By Lemma 4.3, we may assume that $d=\sup\frac{\widehat{\chi}(f)-\widehat{\chi}^{\prime}(f)}{\chi_{0}(f)}$. For any $\epsilon>0$, there exits $f\in\mathfrak{m}$ such that
$(d-\epsilon)\chi_{0}(f)<\widehat{\chi}(f)-\widehat{\chi}^{\prime}(f)=\lim_{k\to\infty}\frac{\chi(f^{k})}{k}-\lim_{k\to\infty}\frac{\chi^{\prime}(f^{k})}{k}.$
Thus for $k\gg 0$ we have
$\chi(f^{k})-\chi^{\prime}(f^{k})\geq(d-2\epsilon)k\chi_{0}(f)=(d-2\epsilon)\chi_{0}(f^{k})$
since $\chi_{0}$ is homogeneous. Similar to the proof of Lemma 4.3 we get $d_{\infty}(\chi,\chi^{\prime})\geq d$. ∎
Lemma 4.5.
Let $\{\chi_{k}\}_{k\in\mathbb{Z}_{>0}}$ be a Cauchy sequence in $(\mathcal{N}^{h},d_{\infty})$. Then there exists $\chi\in\mathcal{N}^{h}$ such that
$\lim_{k\to\infty}d_{\infty}(\chi_{k},\chi)=0.$
Proof.
Given any $\epsilon>0$, by assumption there exists $k\in\mathbb{Z}_{>0}$ such that for any $k,k^{\prime}>K$, $d_{\infty}(\chi_{k},\chi_{k^{\prime}})<\epsilon$. For any $f\in\mathfrak{m}$ and $k,k^{\prime}>K$, by Lemma 4.3 we know that
$|\chi_{k}(f)-\chi_{k^{\prime}}(f)|\leq\chi_{0}(f)d_{\infty}(\chi,\chi^{\prime})<\chi_{0}(f)\cdot\epsilon.$
Hence $\{\chi_{k}(f)\}$ is a Cauchy sequence and we may define $\chi(f):=\lim_{k\to\infty}\chi_{k}(f)$. It is elementary to verify that $\chi\in\mathcal{N}^{h}$. Then we can apply Lemma 4.3 to get $\lim_{k\to\infty}d_{\infty}(\chi_{k},\chi)=0$. ∎
4.2. Locally Lipschitz continuity of log canonical thresholds
In this section we prove that the log canonical thresholds is locally Lipschitz with respect to the $d_{\infty}$-topology.
Proposition 4.6.
Let $\chi\in\mathcal{N}$. For any $\epsilon>0$, there exists $\delta=\delta(\chi,\epsilon)>0\in\mathbb{R}_{>0}$ such that for any $\chi^{\prime}\in\mathcal{N}$ with $d_{\infty}(\chi,\chi^{\prime})<\delta$, we have $|\mathrm{lct}(\chi^{\prime})-\mathrm{lct}(\chi)|\leq\epsilon$.
Proof.
Since $\mathrm{lct}(\chi)=\mathrm{lct}(\widehat{\chi})$ for any $\chi\in\mathcal{N}$, by Lemma 4.4, we may assume that $\chi,\chi^{\prime}\in\mathcal{N}^{h}$. Fix $c=c(\chi):=\inf_{f\in\mathfrak{m}}\frac{\chi(f)}{\chi_{0}(f)}>0$ and denote $a:=\mathrm{lct}(\chi)>0$.
Assume that $d_{\infty}(\chi,\chi^{\prime})<\delta$ for some $\delta\in(0,c)$. By Lemma 4.3, we know that for any $f\in R$,
$\frac{\chi^{\prime}(f)}{\chi_{0}(f)}\geq\frac{\chi(f)}{\chi_{0}(f)}-\delta\geq c-\delta.$
Now we compute
$\frac{v(f)}{\chi^{\prime}(f)}=$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\chi(f)-\chi^{\prime}(f)}{\chi^{\prime}(f)}\right)$ (4.1)
$=$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\chi_{0}(f)}{\chi^{\prime}(f)}\frac{\chi(f)-\chi^{\prime}(f)}{\chi_{0}(f)}\right)$
$\leq$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\delta}{c-\delta}\right)$
For $\delta\in(0,c/2)$, by Lemma 2.14 we get that $v(\mathfrak{a}_{\bullet}(\chi^{\prime}))\leq(1+2\delta/c)v(\mathfrak{a}_{\bullet}(\chi))$ for any $v$. Let
$\delta:=\min\{\frac{c}{2},\frac{c\epsilon}{a}\}>0,$
which depends only on $\epsilon$ and $\chi$. Then for any $\chi^{\prime}\in\mathrm{Fil}^{s}$ with $d_{\infty}(\chi,\chi^{\prime})<\delta$ and any $v\in\mathrm{Val}^{<+\infty}_{X,x}$, we have
$\frac{A(v)}{v(\mathfrak{a}_{\bullet}(\chi^{\prime}))}>(1-\frac{\delta}{2c})\frac{A(v)}{v(\mathfrak{a}_{\bullet}(\chi))}\geq(1-\frac{\delta}{2c})\mathrm{lct}(\chi)=a-\frac{a}{2c}\delta>a-\epsilon.$
Taking infimum among all $v$ gives $\mathrm{lct}(\chi^{\prime})\geq a-\epsilon$.
Interchanging $\chi$ and $\chi^{\prime}$ in (4.1), by Lemma 2.14 we get $v(\mathfrak{a}_{\bullet}(\chi))\leq(1+\delta/c)v(\mathfrak{a}_{\bullet}(\chi^{\prime}))$. By [Blu18, Theorem B.1], there exists $v_{0}\in\mathrm{Val}^{<+\infty}_{X,x}$ computing the lct of $\mathfrak{a}_{\bullet}(\chi)$, thus
$\mathrm{lct}(\chi^{\prime})\leq\frac{A(v_{0})}{v_{0}(\mathfrak{a}_{\bullet}(\chi^{\prime}))}\leq(1+\frac{\delta}{c})\frac{A(v_{0})}{v_{0}(\mathfrak{a}_{\bullet}(\chi))}=(1+\frac{\delta}{c})a\leq a+\epsilon.$
The proof is finished. ∎
4.3. Comparison of different topologies on $\mathrm{Fil}^{s}$
Now we would like to compare the topologies on the space $\mathrm{Fil}^{s}$.  [^4] Since the product topology is in general not first countable and hence not obviously sequential, we avoid using the term finer for them, but only compare convergence for sequences. First we show that the a $d_{1}$-converging seqeunce converges in the $+$-topology.
Lemma 4.7.
Let $\mathfrak{a}_{\bullet}$,$\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ for $k\in\mathbb{Z}_{>0}$. If $\lim_{k}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{a}_{\bullet})=0$, then $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}$ in the $+$-topology.
Proof.
We claim that if $\mathfrak{a}_{\bullet,k}\subset\mathfrak{b}_{\bullet,k}$ satisfy $\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{b}_{\bullet,k})=0$, then for any $v\in\mathrm{Val}^{+}_{X,x}$, we have $\lim_{k\to\infty}(v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{b}_{\bullet,k}))=0$.
For any $v\in\mathrm{Val}^{+}_{X,x}$, applying the claim to $\mathfrak{a}_{\bullet,k},\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet}$ and $\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet}$ separately yields
$\lim_{k\to\infty}|v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{a}_{\bullet})|\leq\lim_{k\to\infty}|v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet})|+\lim_{k\to\infty}|v(\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet})-v(\mathfrak{a}_{\bullet})|=0.$
Now it remains to prove the claim, which follows essentially from the same argument as [BLQ24, Proosition 3.12]. We include a sketch for the reader’s convenience. Assume to the contrary that there exists $\epsilon>0$ and $v\in\mathrm{Val}^{+}_{X,x}$ such that $v(\mathfrak{a}_{\bullet,k})>v(\mathfrak{b}_{\bullet,k})+\epsilon$ for any $k$. Then for $k\in\mathbb{Z}_{>0}$, we may choose $\ell_{k}\in\mathbb{Z}_{>0}$ and $f_{k}\in\mathfrak{b}_{\ell_{k},k}$ such that
$\frac{v(f_{k})}{\ell_{k}}=\frac{v(\mathfrak{b}_{\ell_{k},k})}{\ell_{k}}<v(\mathfrak{a}_{\bullet,k})-\epsilon.$
For $m\in\mathbb{Z}_{>0}$, consider the morphism $\phi_{k,m}:R\to R/\mathfrak{a}_{m\ell_{k},k}$, $g\mapsto f_{k}^{m}g+\mathfrak{a}_{m\ell_{k},k}$. Since $f_{k}^{m}\in\mathfrak{b}_{\ell_{k},k}^{m}\subset\mathfrak{b}_{m\ell_{k},k}$, the image is contained in $\mathfrak{b}_{m\ell_{k},k}/\mathfrak{a}_{m\ell_{k},k}$. Moreover, one can check that $\ker(\phi_{k,m})\subset\mathfrak{a}_{m\ell_{k}\epsilon}(v)$. Hence
$\ell(\mathfrak{b}_{m\ell_{k},k}/\mathfrak{a}_{m\ell_{k},k})\geq\ell(R/\ker(\phi_{k,m})\geq\ell(R/\mathfrak{a}_{m\ell_{k}\epsilon}(v).$
Letting $m\to\infty$ we get $\mathrm{e}(\mathfrak{a}_{\bullet,k})-\mathrm{e}(\mathfrak{b}_{\bullet,k})\geq\epsilon^{n}\mathrm{vol}(v)>0$, which is a contradiction. ∎
Next we compare the two metrics $d_{1}$ and $d_{\infty}$.
Lemma 4.8.
Let $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}$ be filtrations satisfying $\mathfrak{m}^{C\bullet}\subset\mathfrak{a}_{\bullet}$ and $\mathfrak{m}^{C\bullet}\subset\mathfrak{b}_{\bullet}$ for some $C\in\mathbb{R}_{>1}$. Then there exists a constant $M:=M(C)>0$ such that
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\leq M\cdot d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}).$ (4.2)
As a consequence, the $d_{\infty}$-topology is finer than the $d_{1}$-topology.
Proof.
By Lemma 4.4 we may assume $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{h}$. Since $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=d_{1}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet})$ and $d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\geq\max\{d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet}),d_{\infty}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet})\}$, where $\mathfrak{c}_{\bullet}\coloneqq\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$, we may further assume that $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ (but replace $M$ by $2M$).
Denote $d:=d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})$. Then we know that
$0\leq\frac{\chi_{\mathfrak{b}}(f)}{\chi_{\mathfrak{a}}(f)}-1=\frac{\chi_{0}(f)}{\chi_{\mathfrak{a}}(f)}\cdot\frac{\chi_{\mathfrak{b}}(f)-\chi_{\mathfrak{a}}(f)}{\chi_{0}(f)}\leq Cd,$
which implies
$\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}\subset\mathfrak{a}_{\bullet/(1+Cd)}.$
Hence we get
$\mathrm{e}(\mathfrak{a}_{\bullet})\geq\mathrm{e}(\mathfrak{b}_{\bullet})\geq(1+Cd)^{-n}\mathrm{e}(\mathfrak{a}_{\bullet}),$
that is
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=$ $\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet})$
$\leq$ $\left(1-(1+Cd)^{-n}\right)\mathrm{e}(\mathfrak{a}_{\bullet})$
$\leq$ $nCd\cdot\mathrm{e}(\mathfrak{a}_{\bullet})$
$\leq$ $nC^{n+1}\mathrm{e}(\mathfrak{m})\cdot d,$
where in the last inequality we used the fact that $(1+x)^{-n}\geq 1-nx$ for $x\geq 0$. This proves (4.2) with $M\coloneqq 2nC^{n+1}\mathrm{e}(\mathfrak{m})$.
Given $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$, we may take $C\in\mathbb{Z}_{>1}$ such that $\mathfrak{m}^{C}\subset\mathfrak{a}_{1}$ which implies $\mathfrak{m}^{C\bullet}\subset\mathfrak{a}_{\bullet}$. Now for $0<\epsilon<1/2C$ and $\mathfrak{b}_{\bullet}\in\mathrm{Fil}$ with $d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})<\epsilon$, by definition we have $\mathfrak{m}^{2C\bullet}\subset\mathfrak{b}_{\bullet}$. Hence the second assertion follows from (4.2). ∎
To summarize, we have the following diagram for convergence of sequences
$d_{\infty}\text{-convergence}\left\{\begin{aligned} &\implies\text{weak convergence}\\ &\implies d_{1}\text{-convergence}\implies+\text{-convergence}\implies\text{coffecientwise}\end{aligned}\right.$
Here the upper line is obvious; the first two arrows of the lower line follow from Lemma 4.8 and Lemma 4.7 respectively; and the last arrow follows from the fact that $\mathrm{DivVal}_{X,x}\subset\mathrm{Val}^{+}_{X,x}$. It is not hard to see that the upper line is not an equivalence. We do not have any counterexamples for the other reverse arrows.
5. Proof of the main results
Proof of Theorem 1.1.
(1) By Proposition 3.2, $d_{1}$ is a pseudometric on $\mathrm{Fil}$.
By [BLQ24, Corollary 3.17], $\widetilde{\mathfrak{a}}_{\bullet}=\widetilde{\mathfrak{b}}_{\bullet}$ if and only if $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})=\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})$, which holds if and only if $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=0$ by definition. This proves $\mathrm{Fil}^{s}=\mathrm{Fil}/\sim_{1}$.
(2) Assume that $\mathfrak{a}_{\bullet,0}\neq\mathfrak{a}_{\bullet,1}$. We first assume that there exists $f\in R$ and $\lambda\in\mathbb{R}_{>0}$ such that $f\in\mathfrak{a}_{\lambda,0}\backslash\mathfrak{a}_{\lambda,1}$ and $f\notin\mathfrak{a}_{\lambda^{\prime},0}$ for $\lambda^{\prime}>\lambda$. Take $t\in(0,1)$ and $\mu,\nu$ with $(1-t)\mu+t\nu=\lambda$. We claim that $f\notin\mathfrak{a}_{\lambda,t}$. Indeed, if $\mu\leq\lambda$, then $\nu\geq\lambda$, so $f\notin\mathfrak{a}_{\nu,1}$ since $f\notin\mathfrak{a}_{\lambda,1}$. If $\mu>\lambda$, then by the choice of $f$ and $\lambda$ we have $f\notin\mathfrak{a}_{\mu,0}$. So in either case, $f\notin\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$. This shows that $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,1}$. If $\mathfrak{a}_{\bullet,1}\nsubseteq\mathfrak{a}_{\bullet,0}$, then we can repeat the above construction to show that $\mathfrak{a}_{\bullet,t}\neq\mathfrak{a}_{\bullet,0}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$.
Otherwise, $\mathfrak{a}_{\bullet,1}\subsetneq\mathfrak{a}_{\bullet,0}$. We claim that there exists $\lambda\in\mathbb{R}_{>0}$ and $\epsilon>0$ such that $\mathfrak{a}_{\lambda+\epsilon,0}\supsetneq\mathfrak{a}_{\lambda,1}$ and $\mathfrak{a}_{\lambda-\epsilon,1}\supsetneq\mathfrak{a}_{\lambda,1}$. Indeed if all the jumping numbers of $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ coincide then the claim is clear. Otherwise, choose a jumping number $a$ of $\mathfrak{a}_{\bullet,1}$ with $\mathfrak{a}_{a,1}\subsetneq\mathfrak{a}_{a,0}$ which is not a jumping number of $\mathfrak{a}_{\bullet,0}$. Then there exists $a^{\prime}>a$ such that $\mathfrak{a}_{a^{\prime},0}=\mathfrak{a}_{a,0}$. Let $\lambda:=(a+a^{\prime})/2$ and $\epsilon:=\lambda-a$, then $\mathfrak{a}_{\lambda-\epsilon,1}=\lambda_{a,1}\supsetneq\mathfrak{a}_{\lambda,1}$ since $a$ is a jumping number and $\mathfrak{a}_{\lambda+\epsilon,1}=\mathfrak{a}_{a^{\prime},1}=\mathfrak{a}_{a,1}\supsetneq\mathfrak{a}_{a,0}\supset\mathfrak{a}_{\lambda,0}$. Thus we showed $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,1}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$. Now it follows from Proposition 3.6 that $(\mathrm{Fil}^{s},d_{1})$ is convex.
When $R$ contains a field, by Lemma 3.10, the curve $\gamma:[0,1]\to\mathrm{Fil}^{s}$, $t\mapsto\widetilde{\mathfrak{a}}_{\bullet,t}$ is continuous. It computes the distance between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ by Proposition 3.6. Hence $\mathrm{Fil}^{s}$ is a geodesic metric space. ∎
Proof of Theorem 1.2.
The first assertion follows from Proposition 3.27.
Note that by 3.3, $(\mathrm{Fil}^{s}_{R,\mathfrak{m}},d_{1})$ is a rooftop metric space ([Xia23, Definition 5.2]. It is easy to verify that a decreasing (resp. increasing) Cauchy sequence $\mathfrak{a}_{\bullet,k}$ in $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ satisfies $\cap_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ (resp. $\cup_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$). Therefore by Proposition 3.15, Lemma 3.25 and [Xia23, Proposition 5.4], $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is complete. ∎
Proof of Theorem 1.3.
Assertion (1) follows from Lemma 4.2.
By Lemma 4.3, $d_{\infty}$ is non-degenerate when restricted to $\mathcal{N}^{h}$. Moreover, by Lemma 4.5, the metric space $(\mathcal{N}^{h},d_{\infty})$ is complete. Thus assertion (2) is proved. ∎
Proof of Theorem 1.5.
(1) By Definition-Lemma 2.2 and Definition-Lemma 2.3, the set of $\mathfrak{m}$-filtrations is a lattice. It is easy to see that the intersection and join of two linearly bounded $\mathfrak{m}$-filtrations are again linearly bounded, hence $(\mathrm{Fil},\subset,\cap,\vee)$ is a sublattice of all $\mathfrak{m}$-filtrations. Since $\mathfrak{a}\subset\mathfrak{m}$, we know that $\mathfrak{a}^{\lceil\lambda\rceil}\subset\mathfrak{m}^{\lfloor\lambda\rfloor}$ and the filtration $\mathfrak{a}^{\bullet}$ is linearly bounded. Now the map $\mathcal{I}_{\mathfrak{m}}\to\mathrm{Fil}$, $\mathfrak{a}\mapsto\mathfrak{a}^{\bullet}$ is clearly injective, and by Example 2.4, it is a join morphism.
(2) It follows from Definition-Lemma 2.2, Definition-Lemma 2.22 and Proposition 2.24 that $(\mathrm{Fil}^{s},\subset,\cap,\vee_{s})$ is a distributive lattice. By Proposition 2.23, the saturation from $\mathrm{Fil}$ to $\mathrm{Fil}^{s}$ is a joint morphism, which is clearly surjective. By Example 2.25, the map $\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}}\to\mathrm{Fil}^{s}$, $\mathfrak{a}\mapsto\widetilde{\mathfrak{a}^{\bullet}}$ is injective, and is a join morphism as the composition of two join morphisms. ∎
5.1. Continuity properties of log canonical thresholds
In this section, we assume $(R,\mathfrak{m})$ is a klt singularity over a field $\mathbbm{k}$ of characteristic 00.
Proposition 5.1.
Let $\mathfrak{b}_{\bullet,k}\in\mathrm{Fil}^{s}$ be a sequence of filtrations that converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$. Then
$\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\liminf_{k\to\infty}\mathrm{lct}(\mathfrak{b}_{\bullet,k}).$ (5.1)
Proof.
Consider the sequence $\mathfrak{a}_{\bullet,k}:=\cap_{m=k}^{\infty}\mathfrak{b}_{\bullet,m}\subset\mathfrak{b}_{\bullet,k}$ as in the proof of Proposition 3.18. Then the argument shows that $\mathfrak{a}_{\bullet,k}$ is an increasing sequence in $\mathrm{Fil}$ that converges to $\mathfrak{a}_{\bullet}$ weakly. Note that for any $k\in\mathbb{Z}_{>0}$, $\mathrm{lct}(\mathfrak{a}_{\bullet,k})\leq\mathrm{lct}(\mathfrak{b}_{\bullet,k})$, so $\lim_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k})\leq\liminf_{k\to\infty}\mathrm{lct}(\mathfrak{b}_{\bullet,k})$. Thus to prove (5.1), it suffices to show that
$\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\lim_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}).$ (5.2)
Assume to the contrary that $\mathrm{lct}(\mathfrak{a}_{\bullet})>\lim_{k}\mathrm{lct}(\mathfrak{a}_{\bullet,k})=:c$. Fix $\epsilon>0$ such that $\mathrm{lct}(\mathfrak{a}_{\bullet})>c+\epsilon$. Since $\mathrm{lct}(\mathfrak{a}_{\bullet})=\lim_{\lambda}\lambda\cdot\mathrm{lct}(\mathfrak{a}_{\lambda})$, there exists $\Lambda\in\mathbb{R}_{>0}$ such that $\lambda\cdot\mathrm{lct}(\mathfrak{a}_{\lambda})>c+\epsilon$ for any $\lambda\geq\Lambda$. As $\mathfrak{a}_{\bullet,k}$ converges to $\mathfrak{a}_{\bullet}$ weakly, that is, $\mathfrak{a}_{\Lambda}=\cup_{k}\mathfrak{a}_{\Lambda,k}$, there exists $k_{1}\in\mathbb{Z}_{>0}$ such that $\mathfrak{a}_{\Lambda}=\mathfrak{a}_{\Lambda,k_{1}}$.In particular,
$\Lambda\cdot\mathrm{lct}(\mathfrak{a}_{\Lambda,k_{1}})>c+\epsilon.$ (5.3)
Now take $k_{2}\in\mathbb{Z}_{>0}$ such that $\mathrm{lct}(\mathfrak{a}_{\bullet,k})=\sup_{\lambda}\lambda\cdot\mathrm{lct}(\mathfrak{a}_{\lambda,k})<c+\epsilon$ for any $k\geq k_{2}$. For $k\geq\max\{k_{1},k_{2}\}$, combining this with (5.3) we get
$c+\epsilon<\Lambda\cdot\mathrm{lct}(\mathfrak{a}_{\Lambda,k_{1}})=\Lambda\cdot\mathrm{lct}(\mathfrak{a}_{\Lambda,k})\leq\sup_{\lambda}\lambda\cdot\mathrm{lct}(\mathfrak{a}_{\lambda,k})<c+\epsilon,$
where the equality follows from $\mathfrak{a}_{\Lambda,k_{1}}=\mathfrak{a}_{\Lambda}=\mathfrak{a}_{\Lambda,k}$. This is a contradiction, so (5.2) holds and the proof is finished. ∎
Remark 5.2.
Note that in the above proof, one always has $\mathrm{lct}(\mathfrak{a}_{\bullet})\geq\lim_{k}\mathrm{lct}(\mathfrak{a}_{\bullet,k})$, hence equality holds indeed.
Proof of Theorem 1.4.
(1) follows immediately from Proposition 5.1.
To see (2), assume that $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the $+$-topology. By [Blu18, Theorem B.1], there exists $v\in\mathrm{Val}^{<+\infty}_{X,x}$ such that $\mathrm{lct}(\mathfrak{a}_{\bullet})=\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}$. By definition, we have $\lim_{k}v(\mathfrak{a}_{\bullet,k})=v(\mathfrak{a}_{\bullet})$, hence
$\limsup_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k})=\limsup_{k\to\infty}\inf_{w\in\mathrm{Val}^{<+\infty}_{X,x}}\frac{A_{X}(w)}{w(\mathfrak{a}_{\bullet,k})}\leq\lim_{k\to\infty}\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet,k})}=\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}=\mathrm{lct}(\mathfrak{a}_{\bullet}).$
(3) follows from Proposition 4.6. The proof is finished. ∎
6. Discussions
In this section, we discuss the relation to previous work, especially the work in the global case. We also propose some questions related to the main results and provide some toy examples.
6.1. Relation to global results
As was mentioned earlier, our definition for the metric $d_{1}$ is inspired by the Darvas metric $d_{1}$ on $\mathcal{H}(X,\omega)$, the space of smooth Kähler potentials of a compact Kähler manifold $(X,\omega)$, introduced in [Dar15], and is a local analogue of the metric $d_{1}$ on $R=R(X,-rK_{X})$, the section ring of a Fano manifold $X$, introduced in [BJ21]. We now explain an approach to equip the global object with a structure in our setting via the cone construction. We work over the complex numbers $\mathbb{C}$ for simplicity.
Let $(V,L)$ be a smooth polarized variety, that is, $V$ is a smooth projective variety and $L$ is an ample line bundle on $V$. Let $R(V,L)\coloneqq\oplus_{m=0}^{\infty}H^{0}(V,mL)=\oplus_{m=0}^{\infty}R_{m}$, then we know that $R_{0}=\mathbb{C}$. We may replace $L$ by $rL$ for some $r\in\mathbb{Z}_{>0}$ such that $R$ is generated by $R_{1}$ as an $R_{0}$-algebra, and in the sequel, we will always assume $L$ is positive enough such that the above condition is satisfied. In this case, the cone $C(V,L)\coloneqq\mathrm{Spec}R(V,L)$ has a normal, isolated singularity at the vertex 00 defined by $R_{+}(V,L)\coloneqq\oplus_{m>0}R_{m}$.
Denote the localization of $R(V,L)$ at $R_{+}(V,L)$ by $R$. Then we have a normal local domain $(R,\mathfrak{m})$ which is essentially of finite type over $\mathbb{C}$. Now the canonical saturated filtration $\widetilde{\mathfrak{m}^{\bullet}}=\mathfrak{a}_{\bullet}(\mathrm{ord}_{V})$, where $V\subset\mathrm{Bl}_{0}C(V,L)\to C(V,L)$ is identified as the exceptional divisor of the blow-up of the cone point. Any $v\in\mathrm{Val}_{V}$ defines a graded filtration $\mathcal{F}^{\bullet}_{v}$ on $R$ by
$\mathcal{F}^{\lambda}_{v}R_{m}\coloneqq\{s\in R_{m}\mid v(s)\geq\lambda\},$
which is linearly bounded when $v\in\mathrm{Val}^{<+\infty}_{V}$. This is not an $\mathfrak{m}$-filtration, but we could draw a ray from the canonical filtration $\widetilde{\mathfrak{m}^{\bullet}}$, by
$s=\sum s_{l}\in\mathfrak{a}_{\lambda,t}\Leftrightarrow\min\left\{\frac{\ell+tv(s_{\ell})}{1+t}\mid s_{\ell}\neq 0\right\}\geq\lambda,$
for $t\in[0,+\infty)$. Denote the corresponding norm by $\chi_{t}$, then $\chi_{0}=\mathrm{ord}_{V}$ by definition, and for any $s\in R$, we have $\lim_{t\to+\infty}\chi_{t}(s)=v(s)$. For any $t\in[0,+\infty)$, by definition we have $\mathfrak{m}^{\lceil(1+t)\lambda\rceil}\subset\mathfrak{a}_{\lambda,t}$. Moreover, since $\mathcal{F}_{v}$ is linearly bounded, there exists $C\in\mathbb{R}_{>0}$ such that for any $s\in R_{l}$, we have $v(s)\leq Cl$. So $s\in\mathfrak{a}_{\lambda,t}$ will force $\mathrm{ord}_{V}(s)\geq\frac{(1+t)\lambda}{1+C}$, that is, $\mathfrak{a}_{\lambda,t}\subset\mathfrak{m}^{\lfloor\lambda(1+t)/(1+C)\rfloor}$. This shows that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}$.  [^5] Indeed $\mathfrak{a}_{\bullet,t}=\mathfrak{a}_{\bullet}(v_{t})$ for some $v\in\mathrm{Val}^{<+\infty}_{X,x}$.
Thus we may view $v\in\mathrm{Val}^{<+\infty}_{V}$ as an element sitting on the boundary of $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$. We do not go further in this direction. See also [Fin23] and the references therein for some related results.
6.2. Examples
In this section we consider several examples. Note that we only work with linearly bounded filtrations, or equivalently, filtrations with positive multiplicity.
6.2.1. The metric space $(\mathrm{Fil}^{s},d_{1})$ in low dimensions
Example 6.1.
If $R$ is a DVR, then $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\cong\mathbb{R}_{>0}$. Let $\pi\in\mathfrak{m}$ a uniformizer. For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, let $c\coloneqq\mathrm{ord}_{\pi}(\mathfrak{a}_{\bullet})\in\mathbb{R}_{>0}$. Since $\mathrm{ord}_{\pi}$ is the only divisorial valuation up to scaling, for $\lambda\in\mathbb{R}_{>0}$ we have
$\mathfrak{a}_{\lambda}=\{f\in R\mid\mathrm{ord}_{\pi}(f)\geq\lambda\cdot c\}=(\pi^{\lceil c\lambda\rceil})=\mathfrak{m}^{\lceil c\lambda\rceil},$
that is, $\mathfrak{a}_{\bullet}=\mathfrak{m}^{c\bullet}$. Thus we have a bijection of sets $\phi:\mathrm{Fil}^{s}\to\mathbb{R}_{>0}$, $\mathfrak{m}^{c\bullet}\mapsto c$. It is easy to see that $\mathrm{e}(\mathfrak{m}^{c\bullet})=c\cdot\mathrm{e}(\mathfrak{m}^{\bullet})=c$, hence the bijection $\phi$ is an (order-reversing) isometry.
However, the metric space $(\mathrm{Fil}^{s}_{R,\mathfrak{m}},d_{1})$ becomes very complicated as the dimension goes up. When $R=\mathbb{C}[\![x,y]\!]$, by [FJ04], the valuation space $\mathrm{Val}_{R,\mathfrak{m}}$ is a cone over a tree, and $\mathrm{Val}^{+}_{R,\mathfrak{m}}$ is a subset of $\mathrm{Fil}^{s}$.
6.2.2. The metric on spaces of valuations.
By (2.9), Theorem 1.1 gives a metric on the valuation space $\mathrm{Val}^{+}_{X,x}$, defined by
$d_{1}(v_{0},v_{1}):=2\mathrm{e}(\mathfrak{a}_{\bullet}(v_{0})\cap\mathfrak{a}_{\bullet}(v_{1}))-\mathrm{vol}(v_{0})-\mathrm{vol}(v_{1}).$
Example 6.2.
Let $R=\mathbbm{k}[\![x,y]\!]$. Let $v_{0}$ be the weighted blowup along $x,y$ with weights $(2,1)$ and $v_{1}$ the weighted blowup with weights $(1,2)$. It is easy to see that $\mathrm{vol}(v_{i})=2$ for $i=0,1$. Computing the Newton-Okounkov body of the intersection, we get $\mathrm{e}(\mathfrak{a}_{\bullet}(v_{0})\cap\mathfrak{a}_{\bullet}(v_{1}))=8/3$. Hence $d_{1}(v_{0},v_{1})=4/3$. Similarly, let $v_{0}^{(l)}$ and $v_{1}^{(l)}$ be the weighted blowup with weights $(\ell,1)$ and $(1,\ell)$ respectively, then we can compute
$d_{1}(v_{0}^{(l)},v_{1}^{(l)})=\frac{4\ell^{2}}{\ell+1}-2\ell=\frac{2\ell^{2}-2\ell}{\ell+1}$
In particular, $\lim_{\ell\to\infty}d_{1}(v_{0}^{(l)},v_{1}^{(l)})=+\infty$, and hence the metric $d_{1}$ induces a topology on the dual complex of the simple normal crossing pair $(X=\mathrm{Spec}R,L_{0}+L_{1})$, where $L_{0}$ and $L_{1}$ are the lines $(x=0)$ and $(y=0)$ respectively, which is different from the Euclidean topology.
6.2.3. The metrics on spaces of ideals.
Similar to Theorem 1.1, one can also define a pseudo-metric $d_{1}$ on the set $\mathcal{I}_{\mathfrak{m}}$ by
$d_{1}(\mathfrak{a},\mathfrak{b})=2\mathrm{e}(\mathfrak{a}\cap\mathfrak{b})-\mathrm{e}(\mathfrak{a})-\mathrm{e}(\mathfrak{b}).$
By Rees’ Theorem, the Hausdorff quotient of $(\mathcal{I},d_{1})$ can be identified with $(\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}},d_{1})$. Note that with this distance, the inclusion $\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}}\hookrightarrow\mathrm{Fil}^{s}$ is not an embedding but just a contraction map, since in general we only have $\mathrm{e}(\mathfrak{a}\cap\mathfrak{b})\geq\mathrm{e}(\widetilde{\mathfrak{a}^{\bullet}}\cap\widetilde{\mathfrak{b}^{\bullet}})$.
Note that the spaces $\mathrm{Fil}^{s}$ and $\mathrm{Val}_{X,x}^{+}$ both allow an action by $\mathbb{R}_{\geq 0}$. If we consider the set of $\mathfrak{m}$-primary ideals “with coefficients”, that is, the set
$\mathcal{I}_{\mathfrak{m},A}:=\{\mathfrak{a}^{c}\mid\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}},\ c\in A\},$
where $A=\mathbb{Q}_{\geq 0}$ or $\mathbb{R}_{\geq 0}$. Clearly the following metric makes sense
$d_{1,A}(\mathfrak{a}^{c},\mathfrak{b}^{d}):=2\mathrm{e}(\mathfrak{a}^{c\bullet}\cap\mathfrak{b}^{d\bullet})-c^{n}\mathrm{e}(\mathfrak{a})-d^{n}\mathrm{e}(\mathfrak{b}).$
Warning. These are in general two different metrics on $\mathcal{I}_{\mathfrak{m}}$.
6.3. Questions
We propose a few questions related to the metric space $(\mathrm{Fil}^{s},d_{1})$.
6.3.1. Completeness under the Darvas metric
Recall that by [Xia23], in order to prove the completeness of the metric space $(\mathrm{Fil}^{s},d_{1})$, it is enough to check that a (bounded) monotonic sequence has a limit. By Proposition 3.15, this is true for increasing sequences. So it is natural to ask the following question:
Question 6.3.
Let $\{\mathfrak{a}_{\bullet,k}\}_{k\in\mathbb{Z}_{>0}}$ be a decreasing sequence in $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ such that $\mathfrak{a}_{\bullet}\coloneqq\cap\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, do we have
$\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet,k})=0?$
The answer to the question is likely no, but we do not have a counterexample.
6.3.2. More structural questions
We know in Example 6.1 that $\mathrm{Fil}^{s}$ is Euclidean when $R$ is a DVR. This is never the case in higher dimensions. Moreover, geodesics between two incomparable filtrations $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ are never unique, since the geodesic between $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ is distinct from the one obtained from combining the piecewise geodesics connecting $\mathfrak{a}_{\bullet}$, $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$ and $\mathfrak{b}_{\bullet}$. It might be interesting to ask the relationship between the topological or metric properties of $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ and the algebraic properties of the ring $R$. For example,
Question 6.4.
Is $\mathrm{Fil}^{s}$ always contractible? If not, is there a criterion in terms of the algebraic properties of $R$?

### v2

@@
 Abstract.
-We study the geometry of spaces of fitrations on a Noetherian local domain.
+We study the geometry of spaces of filtrations on a Noetherian local domain.
 We introduce a metric $d_{1}$ on the space of saturated filtrations, inspired by the Darvas metric in complex geometry, such that it is a geodesic metric space.
 In the toric case, using Newton-Okounkov bodies, we identify the space of saturated monomial filtrations with a subspace of $L^{1}_{\mathrm{loc}}$.
-We also consider several other topologies on such spaces and study the semi-continuity of the log canonical threshold function in the spirit of Kollár-Demailly.
+We also consider several other topologies on such spaces and study the semi-continuity of the log canonical threshold function in the spirit of Demailly-Kollár.
 Moreover, there is a natural lattice structure on the space of saturated filtrations, which is a generalization of the classical result that the ideals of a ring form a lattice.
@@
 Throughout this paper, we work with a Noetherian local domain $(R,\mathfrak{m},\kappa)$ which is analytically irreducible, that is, the $\mathfrak{m}$-adic completion $\hat{R}$ is a domain.
-As a generalization of ideals, graded sequences of ideals have been extensively studied in the past few years, and have played an important roles in many areas of algebraic geometry, see for example [ELS03, JM12, Cut13] and [BFJ08, BdFF12, BFJ14, Li18, Blu18, Liu18, Xu20, XZ21, LXZ22].
+As a generalization of ideals, graded sequences of ideals have been extensively studied in the past few years, and have played an important role in many areas of algebraic geometry; see for example [ELS03, JM12, Cut13] and [BFJ08, BdFF12, BFJ14, Li18, Blu18, Liu18, Xu20, XZ21, LXZ22].
 Filtrations on a local ring $(R,\mathfrak{m})$ are the continuously indexed version of graded sequences of ideals.
-To be more precise, an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ on $R$ is a collection $\{\mathfrak{a}_{\lambda}\}_{\lambda\in\mathbb{R}_{>0}}$ of $\mathfrak{m}$-primary ideals, which is decreasing, multiplicative and left continuous.
+To be more precise, an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ on $R$ is a collection $\{\mathfrak{a}_{\lambda}\}_{\lambda\in\mathbb{R}_{>0}}$ of $\mathfrak{m}$-primary ideals, which is decreasing, multiplicative, and left continuous.
 The three conditions mean $\mathfrak{a}_{\lambda}\subset\mathfrak{a}_{\mu}$ for $\lambda>\mu$, $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ and $\mathfrak{a}_{\lambda-\epsilon}=\mathfrak{a}_{\lambda}$ for $0<\epsilon\ll 1$, respectively.
 In this paper, we study the geometry of certain spaces of such filtrations and prove some structural results.
-The first main result is that there is a pseudometric $d_{1}$ on the set of $\mathfrak{m}$-filtrations with positive multiplicity, which is an analogue of the Darvas metric introduced in [Dar15] in complex geometry (see [BJ21] for the non-archimedean setting), such that a suitable retraction, the set of saturated filtrations, is a geodesic metric space.
+The first main result is that there is a pseudometric $d_{1}$ on the set of $\mathfrak{m}$-filtrations with positive multiplicity, which is an analogue of the Darvas metric introduced in [Dar15] in complex geometry (see [BJ25] for the non-archimedean setting), such that under a suitable retraction, the set of saturated filtrations is a geodesic metric space.
 In the toric setting, the subspace of saturated monomial filtrations can be identified with the space of cobounded convex sets of the dual cone with the symmetric difference metric, which is naturally a subspace of $L^{1}_{\mathrm{loc}}$.
@@
 In particular, the geodesic metric space $(\mathcal{H},d_{1})$ and its completion, $(\mathcal{E}^{1},d_{1})$, the space of Kähler potentials of finite energy, are closely related to Kähler-Einstein metrics.
-See for example [DR17, BBEGZ19, DL20].
+See, for example, [DR17, BBEGZ19, DL20].
 A comprehensive survey on this metric space and its applications, particularly in relation to Tian’s Properness Conjecture, can be found in [Dar19].
-More recently, Boucksom and Jonsson [BJ21] developed a similar construction in the non-archimedean setting , where they introduced a pseudo-metric $d_{1}$, also referred to as the Darvas metric, on the space $\mathcal{N}$ of norms on the section ring $R(X,-mrK_{X})$ of a projective Fano variety $X$.
+More recently, Boucksom and Jonsson [BJ25] developed a similar construction in the non-archimedean setting, where they introduced a pseudometric $d_{1}$, also referred to as the Darvas metric, on the space $\mathcal{N}$ of norms on the section ring $R(X,L)$ of a polarized variety $(X,L)$.
 Geodesics on this space have been independently studied by Reboulet [Reb22] and by Blum, Liu, Xu and Zhuang [BLXZ21], and these studies have found successful applications in the algebraic theory of $K$-stability and Kähler-Ricci soliton degenerations.
-For some related constructions, see also [Reb20, Wu22, Fin23].
+For some related constructions, see also [Reb20, Wu22, Fin25a].
 Our primary motivation is to introduce an analogue of the aforementioned constructions on the space of filtrations on a local ring $(R,\mathfrak{m})$.
 Given an $\mathfrak{m}$-filtration, following Ein, Lazarsfeld, and Smith [ELS03], the multiplicity of $\mathfrak{a}_{\bullet}$ is defined to be
-$\mathrm{vol}(\mathfrak{a}_{\bullet})\coloneqq\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\ell(R/\mathfrak{a}_{m})}{m^{n}/n!}=\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}},$
+$\mathrm{e}(\mathfrak{a}_{\bullet})\coloneqq\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\ell(R/\mathfrak{a}_{m})}{m^{n}/n!}=\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}},$
 where the existence of the above limit and the equality were proven in increasing generality by [ELS03, Mus02, LM09, Cut13, Cut14].
@@
 Denote the space of all $\mathfrak{m}$-filtrations on $R$ with positive multiplicity by $\mathrm{Fil}_{R,\mathfrak{m}}$, and define a function $d_{1}:\mathrm{Fil}_{R,\mathfrak{m}}\times\mathrm{Fil}_{R,\mathfrak{m}}\to\mathbb{R}_{\geq 0}$ by
-$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\coloneqq 2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}),$ (1.1)
-where $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$ is defined termwise. (1.1) is a similar expression to the Darvas metrics in the global case [Dar15, BJ21].
-Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}_{R,\mathfrak{m}}$, the geodesic $\mathfrak{a}_{\bullet,t}$ between them, introduced in [XZ21, BLQ24], is a segment of $\mathfrak{m}$-filtrations for $t\in[0,1]$ defined by
-$\mathfrak{a}_{\lambda,t}\coloneqq\sum_{\mu,\nu\geq 0,\ (1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}.$ (1.2)
-The definition is a local analogue of the geodesics studied in [BLXZ21].
-Our first main result is that the constructions above endow the Hausdorff quotient of $(\mathrm{Fil}_{R,\mathfrak{m}},d_{1})$ with the structure of a geodesic metric space.
+(1.1) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\coloneqq 2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}),$
+where $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$ is defined termwise. (1.1) is a similar expression to the Darvas metrics in the global case [Dar15, BJ25].
+An issue is that we can have $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})$ for $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$, which means that $d_{1}$ is not a metric on $\mathrm{Fil}_{R,\mathfrak{m}}$.
+An observation of [BLQ24] is that this issue can be resolved by considering saturated filtrations, that is, $\mathfrak{a}_{\bullet}=\widetilde{\mathfrak{a}_{\bullet}}$, where $\widetilde{\mathfrak{a}_{\bullet}}$ is the saturation of $\mathfrak{a}_{\bullet}$ defined by
+[…]
+for any $\lambda\in\mathbb{R}_{>0}$.
+Denote by $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$ the subset consisting of saturated filtrations.
+Note that there is a canonical element $\widetilde{\mathfrak{m}^{\bullet}}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, where $\mathfrak{m}^{\bullet}\coloneqq\{\mathfrak{m}^{\lfloor\lambda\rfloor}\}_{\lambda\in\mathbb{R}_{>0}}$ is the $\mathfrak{m}$-adic filtration.
+Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}_{R,\mathfrak{m}}$, the geodesic $\mathfrak{a}_{\bullet,t}$ between them, introduced in [XZ21, BLQ24], is a segment of $\mathfrak{m}$-filtrations for $t\in[0,1]$, defined by
+(1.2) $\mathfrak{a}_{\lambda,t}\coloneqq\sum_{\mu,\nu\geq 0,\ (1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$
+for any $\lambda>0$. The definition is a local analogue of the geodesics studied in [BLXZ21].
+Our first main result is that the Hausdorff quotient of $(\mathrm{Fil}_{R,\mathfrak{m}},d_{1})$ is naturally $(\mathrm{Fil}^{s}_{R,\mathfrak{m}})$, which is endowed with the structure of a geodesic metric space by the above constructions.
 Theorem 1.1.
-Let $(R,\mathfrak{m})$ be a Noetherian local domain that is analytically irreducible. Denote the set of all linearly bounded $\mathfrak{m}$-filtrations on $R$ by $\mathrm{Fil}_{R,\mathfrak{m}}$. Then
+Let $(R,\mathfrak{m})$ be a Noetherian local domain that is analytically irreducible.
+Denote the set of all linearly bounded $\mathfrak{m}$-filtrations on $R$ by $\mathrm{Fil}_{R,\mathfrak{m}}$ and let $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$ be the subset consisting of saturated ones.
+Then
   1. (1)
-    the function $d_{1}$ given by (1.1) is a pesudometric on $\mathrm{Fil}_{R,\mathfrak{m}}$,
+    the function $d_{1}$ given by (1.1) is a pseudometric on $\mathrm{Fil}_{R,\mathfrak{m}}$,
   2. (2)
@@
 Moreover, if $R$ contains a field, then it is a geodesic metric space, and a geodesic between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ can be given by $\mathfrak{a}_{\bullet,t}$ for $t\in[0,1]$, where $\mathfrak{a}_{\bullet,t}$ is defined by (1.2).
-Here $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is the subset of $\mathrm{Fil}_{R,\mathfrak{m}}$ consisting of saturated $\mathfrak{m}$-filtrations.
-For the definition of saturated filtrations, see section 2.2.
-For definitions related to metric spaces, see section 2.6.
+For more details about saturated filtrations, see section $2.$ For definitions related to metric spaces, see section 2.6.
 1.1.2. The toric case
 In the case where $(R,\mathfrak{m})$ is the local ring of a toric singularity, we can concretely describe the subspace $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ of saturated monomial filtrations.
-Let $\mathbbm{k}$ be a field, $N$ a free abelian group of rank $n\geq 1$ and $M=N^{*}$ its dual.
+Let $\mathbbm{k}$ be an algebraically closed field of characteristic 0.
+Let $N$ be a free abelian group of rank $n\geq 1$ and $M=N^{*}$ its dual.
 Let $\sigma\subset N_{\mathbb{R}}=N\otimes_{\mathbb{Z}}\mathbb{R}$ be a strongly convex rational polyhedral cone of maximal dimension.
-This gives us an affine normal toric variety $X_{\sigma}=\mathrm{Spec}R_{\sigma}=\mathrm{Spec}\mathbbm{k}[\sigma^{\vee}\cap M]$ with a unique torus invariant point $x$, where $\sigma^{\vee}\subset M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$.
+We have an affine normal toric variety $X_{\sigma}=\mathrm{Spec}R_{\sigma}=\mathrm{Spec}\mathbbm{k}[\sigma^{\vee}\cap M]$ with a unique torus invariant point $x$, where $\sigma^{\vee}\subset M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$.
 Let $R$ be the local ring of $X$ at $x$ and $\mathfrak{m}$ its maximal ideal.
-Let $\mathrm{Fil}^{s\mathrm{mon}}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ be the subset of all monomial filtrations (i.e.
-filtrations $\mathfrak{a}_{\bullet}$ such that every $\mathfrak{a}_{\lambda}$ is monomial).
-Via the Okounkov body construction, we identify the metric space $(\mathrm{Fil}^{S,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ with a subspace of the space of co-bounded subsets of $\sigma^{\vee}$ with the symmetric difference metric (sometimes also called the Fréchet-Nykodym-Aronszyan metric).
+Let $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ be the subset of all monomial filtrations (i.e.
+filtrations $\mathfrak{a}_{\bullet}$ such that every $\mathfrak{a}_{\lambda}$ is a monomial ideal).
+Via the Okounkov body construction, we identify the metric space $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ with a subspace of the space of cobounded subsets of $\sigma^{\vee}$ with the symmetric difference metric (sometimes also called the Fréchet-Nykodym-Aronszajn metric).
 Theorem 1.2.
@@
   2. (2)
-    the metric space $(\mathrm{Fil}^{s,\mathrm{mon}}_{X,x},d_{1})$ is complete.
+    the metric space $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ is complete.
 In the above theorem, $\mathcal{P}(\sigma^{\vee})$ denotes the set of all closed convex subsets of $\sigma^{\vee}$ with bounded complement, and the metric $d$ is defined to be the Euclidean volume of the symmetric difference of two subsets.
 For more details, we refer to Section 3.4.
-Note that $P\mapsto\chi_{P}$ gives an embedding $\mathcal{P}(\sigma^{\vee})\to L^{1}_{\mathrm{loc}}(\sigma^{\vee})$.
+Note that $P\mapsto\chi_{P}$ gives an embedding $\mathcal{P}(\sigma^{\vee})\to L^{1}_{\mathrm{loc}}(\sigma^{\vee})$, where $\chi_{P}$ is the characteristic function of $P\subset\sigma$.
 1.1.3. The supnorm metric and homogeneous norms
-As in the global case [BJ21], there is also a correspondence between filtrations and norms in the local setting, see Definition 2.7 for the definition of norms and Definition-Lemma 2.8, Lemma 2.12 and Lemma 2.29 for the dictionary.
+As in the global case [BJ25], there is also a correspondence between filtrations and norms in the local setting.
+Indeed, given $\mathfrak{a}\in\mathrm{Fil}_{R,\mathfrak{m}}$, we can define a function $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ on $R$ by $\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)\coloneqq\sup\{\lambda\in\mathbb{R}\mid f\in\mathfrak{a}_{\lambda}\}$.
+In particular, the canonical saturated filtration $\widetilde{\mathfrak{m}^{\bullet}}$ defines a canonical norm $\chi_{0}\coloneqq\mathrm{ord}_{\widetilde{\mathfrak{m}^{\bullet}}}$.
+See Definition 2.7 for the definition of norms and Definition-Lemma 2.8, Lemma 2.12 and Lemma 2.29 for the detailed dictionary.
 Despite the equivalence, it is sometimes more convenient to use one of the languages.
 As an example, we now consider another pseudometric $d_{\infty}$ on the space $\mathcal{N}_{R,\mathfrak{m}}$ of $\mathfrak{m}$-norms on $R$, which behaves like the uniform metric on the space of continuous functions on a compact space.
-This can also be viewed as the local analogue of the $d_{\infty}$ metric introduced in [BJ21].
+This can also be viewed as the local analogue of the $d_{\infty}$ metric introduced in [BJ25].
 Fix a homogeneous norm $\rho\in\mathcal{N}^{h}_{R,\mathfrak{m}}$, define
-$d_{\infty,\rho}(\chi,\chi^{\prime})\coloneqq\limsup_{\lambda\to\infty}\sup_{\rho(f)\geq\lambda}\frac{|\chi(f)-\chi^{\prime}(f)|}{\rho(f)}.$ (1.3)
+(1.3) $d_{\infty,\rho}(\chi,\chi^{\prime})\coloneqq\limsup_{\lambda\to\infty}\sup_{\rho(f)\geq\lambda}\frac{|\chi(f)-\chi^{\prime}(f)|}{\rho(f)}.$
 We will denote the function $d_{\infty,\chi_{0}}$ by $d_{\infty}$.
@@
 Here $\mathcal{N}^{h}_{R,\mathfrak{m}}$ is the subset of $\mathcal{N}_{R,\mathfrak{m}}$ of homogeneous norms. See Section 2.3 for the definition and more details.
-1.2. Weak topologies and semicontinuity of log canonical thresholds
+1.2. Weak topologies and semi-continuity of log canonical thresholds
 Besides the metrics introduced, we also consider the filtrations, or equivalently, norms, as functions on various spaces, and thus define some weak topologies on the spaces of filtrations.
@@
 Another point of view is to consider a filtration $\mathfrak{a}_{\bullet}$ as a function on certain valuation spaces by $v\mapsto v(\mathfrak{a}_{\bullet})$, where $v(\mathfrak{a}_{\bullet})\coloneqq\lim_{\lambda\to\infty}\frac{v(\mathfrak{a}_{\lambda})}{\lambda}$.
-We call the topology defined this way using the space $\mathrm{Val}^{+}_{R,\mathfrak{m}}$ of valuations with positive volume the $+$-topology.
-We prove that the log canonical threshold satisfies certain semicontinuity on the space of filtrations, and that it is locally Lipschitz continuous with respect to the $d_{\infty}$-topology, which is similar to [DK01, Theorem 3.3].
+We call the topology defined this way using the space $\mathrm{Val}^{+}_{R,\mathfrak{m}}$ of valuations with positive volume the ++-topology.
+We prove that the log canonical threshold satisfies certain semi-continuity on the space of filtrations, and that it is locally Lipschitz continuous with respect to the $d_{\infty}$-topology, which is similar to [DK01, Theorem 3.3].
 Theorem 1.4.
-Let $(X,x)=(\mathrm{Spec}R,\mathfrak{m})$ be a klt singularity over a field $\mathbbm{k}$ of characteristic 00. Then
+Let $(X,x)=(\mathrm{Spec}R,\mathfrak{m})$ be a klt singularity over a field $\mathbbm{k}$ of characteristic 0. Then
   1. (1)
@@
   2. (2)
-    if $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ converges to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the $+$-topology, then
+    if $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ converges to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the ++-topology, then
     $\mathrm{lct}(\mathfrak{a}_{\bullet})\geq\limsup_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}),$
@@
 The spaces of filtrations are equipped with a natural partial order by inclusion, that is, $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ if and only if $\mathfrak{a}_{\lambda}\subset\mathfrak{b}_{\lambda}$ for any $\lambda\in\mathbb{R}_{>0}$.
-One can similarly define the intersection of two filtration termwisely, and in Definition-Lemma 2.3 (Definition-Lemma 2.22, resp.), we define the join $\vee$ (saturated join $\vee_{s}$, resp.), which is the analogue of the sum of two ideals.
+One can similarly define the intersection of two filtrations termwise, and in Definition-Lemma 2.3 (Definition-Lemma 2.22, resp.), we define the join $\vee$ (saturated join $\vee_{s}$, resp.), which is the analogue of the sum of two ideals.
 The next theorem asserts that these operations define natural lattice structures on the spaces of $\mathfrak{m}$-filtrations.
@@
 Moreover, $\mathfrak{a}\mapsto\widetilde{\mathfrak{a}^{\bullet}}$ is an injective join morphism from $\mathcal{I}^{\mathrm{ic}}_{R,\mathfrak{m}}$, the set of integrally closed $\mathfrak{m}$-primary ideals, and the saturation $\mathrm{Fil}_{R,\mathfrak{m}}\to\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is a surjective join morphism.
-This theorem, together with Theorem 1.1, can be viewed as a higher-dimensional analogue of the results of [FJ04].
-We will frequently use the following characterization for saturated filtrations, which can be viewed as an analogue of the Berkovich Maximum Modulus Principle. See also Lemma 2.32.
+In the above theorem, $\widetilde{\mathfrak{a}^{\bullet}}$ denotes the saturation of the $\mathfrak{a}$-adic filtration $\mathfrak{a}^{\bullet}=\{\mathfrak{a}^{\lceil\lambda\rceil}\}$.
+The above theorem, together with Theorem 1.1, can be viewed as a higher-dimensional analogue of the results of [FJ04].
+We will frequently use the following characterization for saturated filtrations, which generalizes results of [Ree56a, Ree56b].
 Proposition 1.6 (= Proposition 2.20).
@@
 $\mathfrak{a}_{\bullet}=\cap_{v\in\Sigma}\mathfrak{a}_{\bullet}(v).$
-As an example of the applications, when $\mathfrak{a}$ is an $\mathfrak{m}$-primary ideal, we prove that the saturation of the $\mathfrak{a}$-adic filtration is determined by the set of Rees valuations of $\mathfrak{a}$, as expected.
+We remark that the above proposition can be viewed as an analogue of the Berkovich Maximum Modulus Principle. See also Lemma 2.32.
+As an application, when $\mathfrak{a}$ is an $\mathfrak{m}$-primary ideal, we prove that the saturation of the $\mathfrak{a}$-adic filtration is determined by the set of Rees valuations of $\mathfrak{a}$, as expected.
 Proposition 1.7 (= Proposition 2.27).
@@
 In Section 2, we recall some definitions and basic properties that will be needed.
-Along the way we also prove some basic facts about filtrations.
+Along the way, we also prove some basic facts about filtrations.
 In Section 3, we introduce the Darvas metric $d_{1}$ and consider its properties.
@@
 We write $c\coloneqq c(a,\alpha,\ldots)$ to mean that $c\in\mathbb{R}_{>0}$ is a constant depending only on $a,\alpha\ldots$.
-Acknowledgement.
+Acknowledgments.
 A large portion of this paper is derived from the author’s doctoral thesis.
@@
 Special thanks are also due to Rémi Reboulet and Mattias Jonsson for reviewing a preliminary version of this paper and offering valuable suggestions.
-Part of this work is done while the author visited the Northwestern University, the University of Utah, Fudan University and BICMR at Peking University.
-The author would like to thank their hospitality and the amazing environment they provided.
+Part of this work is done while the author visited Northwestern University, the University of Utah, Fudan University and BICMR at Peking University.
+The author would like to thank them for their hospitality and the amazing environment they provided.
+We are also grateful to the referees for the corrections and many valuable comments, which improved the quality of this paper.
 2. Preliminaries
 In this section, we recall definitions and basic properties that will be needed in the paper.
-Several results in Section 2.2, including a characterization of saturated filtrations Propositon 2.20, the lattice structure Definition-Lemma 2.22 and the distributivity Proposition 2.24, seem to be new.
+Several results in Section 2.2, including a characterization of saturated filtrations Proposition 2.20, the lattice structure Definition-Lemma 2.22 and the distributivity Proposition 2.24, seem to be new.
 2.1. Filtrations and norms
@@
   2. (2)
-    (left-continuouity) $\mathfrak{a}_{\lambda}=\mathfrak{a}_{\lambda-\epsilon}$ for any $\lambda\in\mathbb{R}_{>0}$ and $0<\epsilon\ll 1$, and
+    (left-continuous) $\mathfrak{a}_{\lambda}=\mathfrak{a}_{\lambda-\epsilon}$ for any $\lambda\in\mathbb{R}_{>0}$ and $0<\epsilon\ll 1$, and
   3. (3)
-    (multiplicativity) $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ for any $\lambda,\mu\in\mathbb{R}_{>0}$.
+    (multiplicative) $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ for any $\lambda,\mu\in\mathbb{R}_{>0}$.
 A filtration $\mathfrak{a}_{\bullet}$ is called an $\mathfrak{m}$-filtration if $\mathfrak{a}_{\lambda}$ is $\mathfrak{m}$-primary for any $\lambda\in\mathbb{R}_{>0}$.
@@
 Their intersection is $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}:=(\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})_{\lambda\in\mathbb{R}_{>0}}$.
-It is not hard to check that this is again an $\mathfrak{m}$-filtrations.
+It is not hard to check that this is again an $\mathfrak{m}$-filtration.
 Definition-Lemma 2.2.
@@
 $=$ $\cap_{l\in\mathbb{Z}_{>1/\lambda}}\cap_{i\in I}\mathfrak{a}_{\lambda-\frac{1}{l},i}=\cap_{i\in I}\mathfrak{a}_{\lambda-\frac{1}{l},i}$
-for some $l\in\mathbb{Z}_{>1/\lambda}$, where the first equality follows from the left-continuouity of $\mathfrak{a}_{\bullet,i}$, and the last equality uses the fact that $R/\mathfrak{b}_{\lambda}$ is Artinian.
+for some $l\in\mathbb{Z}_{>1/\lambda}$, where the first equality follows from the left-continuity of $\mathfrak{a}_{\bullet,i}$, and the last equality uses the fact that $R/\mathfrak{b}_{\lambda}$ is Artinian.
 This proves that condition (2) holds for $\{\cap_{i\in I}\mathfrak{a}_{\lambda,i}\}$, and thus it is an $\mathfrak{m}$-filtration.
@@
 Given an $\mathfrak{m}$-primary ideal $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$, we have an $\mathfrak{m}$-filtration $\mathfrak{a}^{\bullet}$, the $\mathfrak{a}$-adic filtration, defined by $\{\mathfrak{a}^{\lceil\lambda\rceil}\}_{\lambda\in\mathbb{R}_{>0}}$.
-Thus we get an injection from the set $\mathcal{I}_{\mathfrak{m}}$ to the set of all $\mathfrak{m}$-filtrations by $\mathfrak{a}\mapsto\mathfrak{a}^{\bullet}$ In particular, there is a canonical $\mathfrak{m}$-filtration $\mathfrak{m}^{\bullet}$.
+Thus we get an injection from the set $\mathcal{I}_{\mathfrak{m}}$ to the set of all $\mathfrak{m}$-filtrations defined by $\mathfrak{a}\mapsto\mathfrak{a}^{\bullet}$.
+In particular, there is a canonical $\mathfrak{m}$-filtration $\mathfrak{m}^{\bullet}$.
 For $\mathfrak{a},\mathfrak{b}\in\mathcal{I}_{\mathfrak{m}}$, we claim that $\mathfrak{a}^{\bullet}\vee\mathfrak{b}^{\bullet}=(\mathfrak{a}+\mathfrak{b})^{\bullet}$.
@@
 In Section 2.2, we will define a similar notion, the saturated join, and give a formula for it.
-The following proposition, which is a combination of Definition-Lemma 2.2 and Defnition-Lemma 2.3, justifies the term join.
+The following proposition, which is a combination of Definition-Lemma 2.2 and Definition-Lemma 2.3, justifies the term join.
 Proposition 2.5.
-The set of all $\mathfrak{m}$-filtrations with the partial order by inclusion form a lattice, where the meet of two elements is given by their intersection, and the join is defined as in Definition-Lemma 2.3.
+The set of all $\mathfrak{m}$-filtrations with the partial order by inclusion forms a lattice, where the meet of two elements is given by their intersection, and the join is defined as in Definition-Lemma 2.3.
 ∎
@@
 If $R$ is a $\mathbbm{k}$-algebra for a field $\mathbbm{k}$, then we will always assume that $\chi$ is a $\mathbbm{k}$-norm, that is, $\chi(af)=\chi(f)$ for any $a\in\mathbbm{k}^{\times}$ and $f\in R$.
-As in the global case, we adopt the additive terminology here. For a norm $\chi$ on $R$, $\|\cdot\|_{\chi}:=e^{-\chi(\cdot)}$ defines a non-Archimedean norm on $R$ in the sense of [BGR84].
+Here, as in the global case, we adopt the additive terminology here. For a norm $\chi$ on $R$, $\|\cdot\|_{\chi}:=e^{-\chi(\cdot)}$ defines a non-Archimedean norm on $R$ in the sense of [BGR84].
 There is a correspondence between seminorms and filtrations, given by the lemma below.
@@
 Given an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$, its order function $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is defined by
-$\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f):=\sup\{\lambda\in\mathbb{R}_{>0}\mid f\in\mathfrak{a}_{\lambda}\},$ (2.1)
+(2.1) $\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f):=\sup\{\lambda\in\mathbb{R}_{>0}\mid f\in\mathfrak{a}_{\lambda}\},$
 which is a seminorm on $R$. If $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda\ll 1$, then it is an $\mathfrak{m}$-seminorm.
 Conversely, given an $\mathfrak{m}$-seminorm $\chi$, define $\mathfrak{a}_{\bullet}(\chi)$ by
-$\mathfrak{a}_{\lambda}(\chi):=\{f\in R\mid\chi(f)\geq\lambda\}.$ (2.2)
+(2.2) $\mathfrak{a}_{\lambda}(\chi):=\{f\in R\mid\chi(f)\geq\lambda\}.$
 Then $\mathfrak{a}_{\bullet}(\chi)$ is an $\mathfrak{m}$-filtration, called the associated filtration of $\chi$.
@@
 Clearly the collection of $\mathfrak{m}$-primary ideals $\{\mathfrak{a}_{\lambda}(\chi)\}$ is decreasing, and it is multiplicative by condition (3) of Definition 2.7.
-To show left continuity, note that we can write $\mathfrak{a}_{\lambda}(\chi)$ as an decreasing intersection
+To show left continuity, note that we can write $\mathfrak{a}_{\lambda}(\chi)$ as a decreasing intersection
 $\mathfrak{a}_{\lambda}(\chi)=\cap_{\epsilon>0}\mathfrak{a}_{\lambda-\epsilon}(\chi),$
@@
 2.1.3. Divisorial valuations
+Recall that a valuation $v\in\mathrm{Val}_{R,\mathfrak{m}}$ defines a local ring $\mathcal{O}_{v}$, the valuation ring of $v$, by
+$\mathcal{O}_{v}\coloneqq\{f\in\mathrm{Frac}(R)\mid v(f)\geq 0\}\text{ with the maximal ideal }\mathfrak{m}_{v}=\{f\in\mathcal{O}_{v}\mid v(f)\geq 0\}.$
+The residue field of $v$ is $\kappa_{v}\coloneqq\mathcal{O}_{v}/\mathfrak{m}_{v}$.
 A valuation $v\in\mathrm{Val}_{R,\mathfrak{m}}$ is divisorial if
@@
 By [ELS03, Mus02, LM09, Cut13, Cut14] in increasing generality, the above limit exists and
-$\mathrm{e}(\mathfrak{a}_{\bullet})=\lim_{m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}}=\inf_{m}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}};$ (2.3)
+(2.3) $\mathrm{e}(\mathfrak{a}_{\bullet})=\lim_{m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}}=\inf_{m}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}};$
 see in particular [Cut14, Theorem 6.5].
@@
 By Krull intersection theorem, a linearly bounded $\mathfrak{m}$-filtration satisfies $\cap_{\lambda\in\mathbb{R}_{\geq 0}}\mathfrak{a}_{\lambda}=\{0\}$.
-In particular, by Definitin-Lemma 2.8, we have the following
+In particular, by Definition-Lemma 2.8, we have the following
 Lemma 2.12.
-An $\mathfrak{m}$-filtration is linearly bounded if and only if $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is a linearly bounded $\mathfrak{m}$-norm.
-We will consider filtrations with positive multiplicity, which are exactly those which are linearly bounded by the following lemma.
+An $\mathfrak{m}$-filtration is linearly bounded if and only if $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is a linearly bounded $\mathfrak{m}$-norm.∎
+We will consider filtrations with positive multiplicity, which are exactly the linearly bounded ones by the following lemma.
 Lemma 2.13.
@@
 By Definition-Lemma 2.8 and Lemma 2.12, $\mathcal{N}$ is a subset of $\mathrm{Fil}$.
-There is a partial order $\leq$ on $\mathcal{N}$ defined pointwisely on $R$, that is,
-$\chi\leq\chi^{\prime}\text{ if and only if }\chi(f)\leq\chi^{\prime}(f)\text{ for any }f\in R,$
+There is a natural pointwise partial order $\leq$ on $\mathcal{N}$ on $R$, defined by
+[…]
 which corresponds to the partial order defined by inclusion on filtrations.
@@
 Conversely, assume that $\mathfrak{a}_{\bullet}\nsubseteq\mathfrak{b}_{\bullet}$.
-Then there exists $\lambda\in\mathbb{R}_{>0}$ and $f\in\mathfrak{a}_{\lambda}$ such that $f\notin\mathfrak{b}_{\lambda}$.
-This implies that there exists $\epsilon>0$ and $v\in\mathrm{DivVal}_{X,x}$ such that $v(f)<\lambda\cdot v(\mathfrak{b}_{\bullet})-\epsilon$.
+Then there exist $\lambda\in\mathbb{R}_{>0}$ and $f\in\mathfrak{a}_{\lambda}$ such that $f\notin\mathfrak{b}_{\lambda}$.
+This implies that there exist $\epsilon>0$ and $v\in\mathrm{DivVal}_{X,x}$ such that $v(f)<\lambda\cdot v(\mathfrak{b}_{\bullet})-\epsilon$.
 Hence for any $m\in\mathbb{Z}_{>0}$, we have
@@
 In other words, $\widetilde{\mathfrak{a}}_{\bullet}$ is the unique maximal $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}\supset\mathfrak{a}_{\bullet}$ such that $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})$.
-We now prove a formula for saturations.
+We now prove a formula for saturations, which can be viewed as an analogue of the Berkovich maximal modulus principle, where one considers the subset of divisorial valuations of the Berkovich spectrum.
 Lemma 2.19.
@@
 Since $\mathrm{DivVal}_{R,\mathfrak{m}}\subset\mathrm{Val}^{+}_{R,\mathfrak{m}}$, the proof is finished. ∎
+In particular, $\mathfrak{a}_{\bullet}(v)$ is saturated for any $v\in\mathrm{Val}^{+}_{R,\mathfrak{m}}$.
+By the above proposition, saturated filtrations (or norms) can be considered as the local analogue of maximal norms in the sense of [BJ25, Definition 6.16].
 Corollary 2.21.
@@
 If $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}$, then $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}=\widetilde{\mathfrak{a}_{\bullet}\vee\mathfrak{b}_{\bullet}}=\widetilde{\mathfrak{a}}_{\bullet}\vee_{s}\widetilde{\mathfrak{b}}_{\bullet}$, and
-$\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}}).$ (2.4)
+(2.4) $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}}).$
 Proof.
@@
 To prove the formula (2.4), we first rewrite Lemma 2.19 as
-$\mathfrak{a}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a}_{\bullet})})$ (2.5)
+(2.5) $\mathfrak{a}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a}_{\bullet})})$
 for $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$. Let $\mathfrak{c}_{\bullet}$ be the $\mathfrak{m}$-filtration defined by the right-handed side of (2.4). Since
@@
 Hence $v(\mathfrak{c}_{\bullet})\leq\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$ for any $v\in\mathrm{Val}_{R,\mathfrak{m}}$.
-By (2.4), for any $v\in\mathrm{DivVal}_{X,x}$ we have $v(\mathfrak{c}_{\bullet})\geq v(\mathfrak{a}_{\bullet}(v/\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}))=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$.
+By (2.4), for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$ we have $v(\mathfrak{c}_{\bullet})\geq v(\mathfrak{a}_{\bullet}(v/\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}))=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$.
 Thus
-$v(\mathfrak{c}_{\bullet})=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}.$ (2.6)
-We now verify that $\mathfrak{c}_{\bullet}\in\mathrm{Fil}$.
-Take $C\in\mathbb{R}_{>0}$ such that $\mathfrak{a}_{C\bullet}\subset\mathfrak{m}^{\bullet}$ and $\mathfrak{b}_{C\bullet}\subset\mathfrak{m}^{\bullet}$.
-Since $\mathfrak{m}^{\bullet}\subset\widetilde{\mathfrak{m}}^{\bullet}$, for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$, we have $Cv(\mathfrak{a}_{\bullet})\geq v(\mathfrak{m}^{\bullet})=v(\widetilde{\mathfrak{m}}^{\bullet})$ and similarly $Cv(\mathfrak{b}_{\bullet})\geq v(\widetilde{\mathfrak{m}}^{\bullet})$.
-By (2.6), $Cv(\mathfrak{c}_{\bullet})\geq v(\widetilde{\mathfrak{m}}^{\bullet})$.
-So by Lemma 2.16 $\mathfrak{c}_{\bullet}\subset\widetilde{\mathfrak{m}}^{\bullet}$ is linearly bounded.
+(2.6) $v(\mathfrak{c}_{\bullet})=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}.$
+[…]
 By Corollary 2.21, $\mathfrak{c}_{\bullet}\in\mathrm{Fil}^{s}$, which implies $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}\subset\mathfrak{c}_{\bullet}$.
@@
 $v(f)/\lambda\geq$ $\max\{v(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}),v(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})\}$
-$\geq$ $\max\{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\},\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{c}_{\bullet})\}\}$
+$=$ $\max\{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\},\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{c}_{\bullet})\}\}$
 $\geq$ $\min\{v(\mathfrak{a}_{\bullet}),\max\{v(\mathfrak{b}_{\bullet}),v(\mathfrak{c}_{\bullet})\}\},$
@@
 By Proposition 2.23 again, we get $f\in(\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet}))_{\lambda}$.
-This shows that $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{c}_{\bullet}\cap\mathfrak{b}_{\bullet})\supset(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})\cap\mathfrak{b}_{\bullet}$ and the equality must hold.
+This shows that $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})\supset(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet})\cap(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})$ and the equality must hold.
 ∎
@@
 We continue to consider the adic filtration $\mathfrak{a}^{\bullet}$ of $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$. By definition,
-$(\widetilde{\mathfrak{a}^{\bullet}})_{\lambda}=\{f\in\mathfrak{m}\mid v(f)\geq\lambda\cdot v(\mathfrak{a}^{\bullet})\text{ for any }v\in\mathrm{DivVal}_{R,\mathfrak{m}}\}.$
+[…]
 Since $v(\mathfrak{a}^{\bullet})=\lim_{\lambda\to\infty}v(\mathfrak{a}^{\lceil\lambda\rceil})/\lambda=\lim_{\lambda\to\infty}\lceil\lambda\rceil v(\mathfrak{a})/\lambda=v(\mathfrak{a})$, for $m\in\mathbb{Z}_{>0}$ we have
-$(\widetilde{\mathfrak{a}^{\bullet}})_{m}=\{f\in\mathfrak{m}\mid v(f)\geq mv(\mathfrak{a})=v(\mathfrak{a}^{m})\text{ for any }v\in\mathrm{DivVal}_{R,\mathfrak{m}}\}.$
+[…]
 Hence $(\widetilde{\mathfrak{a}^{\bullet})}_{m}=\overline{\mathfrak{a}^{m}}$ is the integral closure of $\mathfrak{a}^{m}$.
@@
 Let $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$ be an $\mathfrak{m}$-primary ideal. Then
-$\widetilde{\mathfrak{a}^{\bullet}}=\cap_{v\in RV(\mathfrak{a})}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a})}),$ (2.7)
+(2.7) $\widetilde{\mathfrak{a}^{\bullet}}=\cap_{v\in RV(\mathfrak{a})}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a})}),$
 where $RV(\mathfrak{a})$ is the set of Rees valuations of $\mathfrak{a}$.
@@
 (2) Choose a set $\{f_{1},\ldots,f_{r}\}$ of generators of $\mathfrak{m}$ and $D\in\mathbb{Z}_{>0}$ such that $\mathfrak{m}^{D}\subset\mathfrak{a}_{1}$.
-Then $f_{i}^{D}\in\mathfrak{a}_{1}$ and hence $f_{i}\in\mathfrak{a}_{1/D}$ by assumption.
+Then $f_{i}^{D}\in\mathfrak{a}_{1}$, and hence $f_{i}\in\mathfrak{a}_{1/D}$ by assumption.
 So $\mathfrak{m}\subset\mathfrak{a}_{1/D}$ and hence $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda<1/D$.
@@
 Definition 2.30.
-[CP24, Definition 3.5](c.f.
-[BFJ14, Section 4]) Given a filtration $\mathfrak{a}_{\bullet}$ on $R$, its asymptotic Samuel function $\widehat{\mathrm{ord}}_{\mathfrak{a}_{\bullet}}:R\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is defined by $\widehat{\mathrm{ord}}_{\mathfrak{a}_{\bullet}}(f):=\lim_{k\to\infty}\frac{\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f^{k})}{k}$ for any $f\in R$.
-Denote the corresponding filtration by $\widehat{\mathfrak{a}}_{\bullet}$, called the homogenization of $\mathfrak{a}_{\bullet}$.
-The limit exists by Feteke’s lemma, see [CP24, Theorem 3.4]. Similar to Theorem 2.18 for the saturation, the homogenization has the following property.
+(c.f.
+[HS06, Definition 6.9.3] and [BFJ14, Section 4]) Given a norm $\chi$ on $R$, its homogenization $\widehat{\chi}:R\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is defined by $\widehat{\chi}(f):=\lim_{k\to\infty}\frac{\chi(f^{k})}{k}$ for any $f\in R$.
+[…]
+The limit exists by Fekete’s lemma, see, for example, [CP24, Theorem 3.4].
+Hence it is not hard to see that $\widehat{\mathfrak{a}}_{\bullet}$ is the smallest homogeneous filtration that contains $\mathfrak{a}_{\bullet}$.
+Moreover, the homogenization has the following property.
 Theorem 2.31.
-[CP24, Theorem 5.5] Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration.
-Then $\widehat{\mathfrak{a}}_{\bullet}$ is the unique largest $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}$ such that $\widehat{\mathrm{ord}_{\mathfrak{a}_{\bullet}}}=\widehat{\mathrm{ord}_{\mathfrak{b}_{\bullet}}}$.
-[…]
+[CP24] Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration. Then
+  1. (1)
+    $\widehat{\mathfrak{a}}_{\bullet}$ is the unique largest $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}$ such that $\widehat{\mathrm{ord}_{\mathfrak{a}_{\bullet}}}=\widehat{\mathrm{ord}_{\mathfrak{b}_{\bullet}}}$, and
+  2. (2)
+    $\widehat{\mathfrak{a}}_{\bullet}\subset\widetilde{\mathfrak{a}}_{\bullet}$.
+Before proceeding, we briefly recall the definition of the Berkovich analytification.
+Fix $\chi\in\mathcal{N}$.
+By linear boundedness, we know that the $\mathfrak{m}$-adic completion $\hat{R}$ is complete with respect to $\chi$.
+Let $\mathcal{M}=\mathcal{M}(R,\chi)$ be the Berkovich spectrum of the Banach ring $(\hat{R},\chi)$, that is, the set of multiplicative $\mathfrak{m}$-seminorms $w$ on $\hat{R}$ such that $w\geq\chi$.
+Note that the homogenization corresponds to the spectral radius construction of [Ber90], since we adopted the multiplicative notation.
+We thank the referee for pointing out a mistake in the statement of the previous version.
+Next, we give a characterization for homogeneous norms similar to Proposition 2.20, which is an easier local analogue of [Fin25b, Theorem 3.17] and [BJ22, Theorem 2.16].
 Lemma 2.32.
 A norm $\chi\in\mathcal{N}$ is homogeneous if and only if it is of the form
-$\chi=\inf_{v\in I}v,$
-where $I$ is a non-empty subset of $X^{\mathrm{an}}_{x}$.
+$\chi=\inf_{w\in\Sigma}w,$
+where $\Sigma\subset\mathcal{M}$ is a non-empty subset.
 Proof.
-If $\chi(f)=\inf_{v\in I}v(f)$, then $\chi$ is homogeneous, since the same is true for each $v$.
-The reverse implication is essentially a restatement of the Berkovich maximum modulus principle, applied to the seminormed ring $(R,\chi)$. See, for example, [Jon20, Theorem 2.2.1]
+If $\chi(f)=\inf_{w\in I}w(f)$, then $\chi$ is homogeneous, since the same is true for each $w$.
+The reverse implication is essentially a restatement of the Berkovich Maximum Modulus principle, see [Ber90, Theorem 1.3.1]. ∎
 Denote the subspace of $\mathrm{Fil}$ consisting of homogeneous filtrations by $\mathrm{Fil}^{h}$.
+For adic filtrations, it is not hard to see that homogenization and saturation coincide.
+Lemma 2.33.
+Let $\mathfrak{a}\in\mathrm{Fil}_{R,\mathfrak{m}}$. Then $\widetilde{\mathfrak{a}^{\bullet}}=\widehat{\mathfrak{a}^{\bullet}}$.
+Proof.
+By Example 2.25, $(\widetilde{\mathfrak{a}^{\bullet}})_{m}=\overline{\mathfrak{a}^{m}}$ is the integral closure of $\mathfrak{a}^{m}$ for any $m\in\mathbb{Z}_{>0}$.
+The same calculation implies $(\widehat{\mathfrak{a}^{\bullet}})_{m}=\overline{\mathfrak{a}^{m}}$ by [HS06, Proposition 6.10].
+So the lemma follows from a simple calculation since both filtrations are homogeneous by Lemma 2.29.
+∎
 2.4. Log canonical thresholds and normalized volumes
@@
 We will write $\mathrm{lct}(\mathfrak{a}_{\bullet})$ when there is no ambiguity.
-For more properties of lct, we refer to [JM12].
-We refer to [LLX20, Zhu23] for a comprehensive survey on normalized volumes.
+For more properties of lct, we refer to [JM12]. We refer to [LLX20, Zhu25] for a comprehensive survey on normalized volumes.
 Lemma 2.35.
-(c.f. [Zhu24, Lemma 3.4]) There exists $c_{0}=c_{0}(n)>0$ depending only one $n$ such that for any $n$-dimensional klt singularity $x\in X$ and $v\in\mathrm{Val}_{X,x}$,
-$\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}\leq c_{0}\cdot\frac{\widehat{\mathrm{vol}}_{x,X}(v)}{\widehat{\mathrm{vol}}(x,X)}\cdot\mathrm{lct}(X;\mathfrak{a}_{\bullet})$ (2.8)
+(c.f. [Zhu24, Lemma 3.4]) There exists $c_{0}=c_{0}(n)>0$ depending only on $n$ such that for any $n$-dimensional klt singularity $x\in X$ and $v\in\mathrm{Val}_{X,x}$,
+(2.8) $\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}\leq c_{0}\cdot\frac{\widehat{\mathrm{vol}}_{x,X}(v)}{\widehat{\mathrm{vol}}(x,X)}\cdot\mathrm{lct}(X;\mathfrak{a}_{\bullet})$
 for any $\mathfrak{a}_{\bullet}\in\mathrm{Fil}_{x}(X)$.
@@
 Definition 2.37.
-A point $z$ in a metric space $(S,d)$ is said to be between two points $x$ and $y$, if all three points are distinct, and $d(x,y)=d(x,z)+d(z,x)$.
+A point $z$ in a metric space $(S,d)$ is said to be between two points $x$ and $y$, if all three points are distinct, and $d(x,y)=d(x,z)+d(z,y)$.
 A metric space $(S,d)$ is convex if for any two distinct points $x,y\in S$, there exists a point $z$ between $x$ and $y$.
@@
   • the space of quasi-monomial valuations when $R$ contains a field, denoted by $\mathrm{QMVal}_{R,\mathfrak{m}}$,
-  • the space of valuations with finite log discrepancy, when $x\in X=\mathrm{Spec}R$ is a klt singularity over a field of characteristic 00, denoted by $\mathrm{Val}^{<+\infty}_{X,x}$, and
+  • the space of valuations with finite log discrepancy, when $x\in X=\mathrm{Spec}R$ is a klt singularity over a field of characteristic 0, denoted by $\mathrm{Val}^{<+\infty}_{X,x}$, and
   • the space of valuations with positive volume, denoted by $\mathrm{Val}^{+}_{X,x}$, which is equal to the space of linearly bounded valuations by Lemma 2.13.
@@
 We have the following inclusions
-$\mathrm{DivVal}_{X,x}\subset\mathrm{QMVal}_{X,x}\subset\mathrm{Val}^{<+\infty}_{X,x}\subset\mathrm{Val}^{+}_{X,x}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{h}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$ (2.9)
+(2.9) $\mathrm{DivVal}_{X,x}\subset\mathrm{QMVal}_{X,x}\subset\mathrm{Val}^{<+\infty}_{X,x}\subset\mathrm{Val}^{+}_{X,x}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{h}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$
 where the third inclusion follows from Li’s properness estimate [Li18, Theorem 1.4]. By Lemma 2.17, to compute the saturation of a filtration, one can use any of the above spaces of valuations.
 2.7.1. Weak topologies on $\mathrm{Fil}$
-There are more ways to define a topology on the space $\mathrm{Fil}$ (and $\mathrm{Fil}^{s}$ and $\mathrm{Fil}^{s}$, respectively), depending on the point of view.
+There are more ways to define a topology on the space $\mathrm{Fil}$ (as well as on $\mathrm{Fil}^{s}$ and $\mathrm{Fil}^{h}$), depending on the point of view.
 For example, a filtration $\mathfrak{a}_{\bullet}$ acts naturally on the ring $R$ as a function $\chi_{\mathfrak{a}_{\bullet}}$ (or dually).
-The product topology, or the “weak-* topology” induced by this action will be called the weak topology, which is the coarest topology such that for any $f\in R\backslash\{0\}$, the function $\phi_{f}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)$ is continuous.
-In other words, a subbase of the weak topology is given by $\{\phi_{f}^{-1}(U)\mid U\subset\mathbb{R}_{\geq 0}\text{ is open, }f\in R\backslash\{0\}\}$.
+The product topology, or the “weak-* topology” induced by this action will be called the weak topology, which is the coarsest topology such that for any $f\in R\backslash\{0\}$, the function $\phi_{f}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)$ is continuous.
+In other words, a subbase of the weak topology is given by $\phi_{f}^{-1}(U)\mid U\subset\mathbb{R}_{\geq 0}\text{ is open, }f\in R\backslash\{0\}\}$.
 Under this topology, all valuation spaces in (2.9) equipped with the weak topology, defined in [JM12], are subspaces of $\mathrm{Fil}$.
-It is not hard to check that a sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}$ converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$ if and only of $\chi_{k}(f)\to\chi(f)$ for any $f\in R\backslash\{0\}$.
+It is not hard to check that a sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}$ converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$ if and only if $\chi_{k}(f)\to\chi(f)$ for any $f\in R\backslash\{0\}$.
 Similarly, a valuation has a natural action on the spaces of filtrations (or dually), that is, any $v\in\mathrm{Val}_{X,x}$ gives a function $\mathrm{ev}_{v}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto v(\mathfrak{a}_{\bullet})$.
-Hence one can consider the “weak topology” induced by this action, that is, the coarest topology on $\mathrm{Fil}^{s}$ such that $\mathrm{ev}_{v}$ is continuous for any $v\in X^{{}^{\mathrm{an}}}$, $\mathrm{Val}_{X,x}$, $\mathrm{Val}^{+}_{X,x}$ or $\mathrm{DivVal}_{X,x}$, respectively.
-Similarly one can check that a sequence $\{\mathfrak{a}_{\bullet,k}\}$ converges to $\mathfrak{a}_{\bullet}$ under these topologies if any only if $v(\mathfrak{a}_{\bullet,k})\to v(\mathfrak{a}_{\bullet})$ for any $v$ in the corresponding valuation space.
-In this paper, we will mainly consider the topology defined using $\mathrm{Val}^{+}_{X,x}$, and call it the $+$-topology.
+Hence one can consider the “weak topology” induced by this action, that is, the coarsest topology on $\mathrm{Fil}^{s}$ such that $\mathrm{ev}_{v}$ is continuous for any $v\in X^{{}^{\mathrm{an}}}$, $\mathrm{Val}_{X,x}$, $\mathrm{Val}^{+}_{X,x}$ or $\mathrm{DivVal}_{X,x}$, respectively.
+Similarly, one can check that a sequence $\{\mathfrak{a}_{\bullet,k}\}$ converges to $\mathfrak{a}_{\bullet}$ under these topologies if and only if $v(\mathfrak{a}_{\bullet,k})\to v(\mathfrak{a}_{\bullet})$ for any $v$ in the corresponding valuation space.
+In this paper, we will mainly consider the topology defined using $\mathrm{Val}^{+}_{X,x}$, and call it the ++-topology.
 The product topology defined using $\mathrm{DivVal}_{X,x}$ as above was considered in [BdFFU15] for $b$-divisors.
@@
 Remark 2.39.
-Since $v(\mathfrak{a}_{\bullet})=v(\widetilde{\mathfrak{a}}_{\bullet})$ by [BLQ24, Proposition 3.9] for $v\in\mathrm{DivVal}_{X,x}$, the functions $\mathrm{ev}_{v}$ do not distinguish $\mathfrak{a}_{\bullet}$ and its saturation.
-Hence the $+$-topology (and the coefficientwise topology) only makes sense on $\mathrm{Fil}^{s}$.
-Similar issues appear if we define the topology on $\mathrm{Fil}$ using other valuation spaces.
-But all the topologies are well-defined on $\mathrm{Fil}^{s}$.
+[…]
 Later we will define metrics on the space $\mathrm{Fil}^{s}$. See Section 4.3 for a comparison of these topologies.
@@
 Given $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}$. If $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$, then let
-$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}):=\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}).$ (3.1)
+(3.1) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}):=\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}).$
 In general, define
-$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}):=d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{b}_{\bullet}).$ (3.2)
+(3.2) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}):=d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{b}_{\bullet}).$
 Proposition 3.2.
 $d_{1}$ satisfies the triangle inequality, that is, for $\mathfrak{a}_{\bullet}$, $\mathfrak{b}_{\bullet}$, $\mathfrak{c}_{\bullet}\in\mathrm{Fil}$, we have
-$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$ (3.3)
+(3.3) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$
 Hence $d_{1}$ is a pseudometric on $\mathrm{Fil}$.
@@
 which will follow immediately from the following comparison on colengths
-$\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{c}_{\lambda})\leq\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})+\ell(R/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})-\ell(R/\mathfrak{b}_{\lambda})$ (3.4)
+(3.4) $\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{c}_{\lambda})\leq\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})+\ell(R/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})-\ell(R/\mathfrak{b}_{\lambda})$
 for any $\lambda\in\mathbb{R}_{>0}$. But we compute
@@
 Using the identification between $\mathrm{Fil}$ and $\mathcal{N}$ Lemma 2.8, we get a pseudometric $d_{1}$ on $\mathcal{N}$.
-The following lemma asserts that $(\mathrm{Fil},d_{1})$ is a rooftop metric space in the sense of [Xia23], which is an analogue of [Dar19, Proposition 3.35] and [BJ21, Lemma 3.2].
+The following lemma asserts that $(\mathrm{Fil},d_{1})$ is a rooftop metric space in the sense of [Xia23], which is an analogue of [Dar19, Proposition 3.35] and [BJ25, Lemma 3.2].
 Lemma 3.3.
 For $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}\in\mathrm{Fil}$, we have
-$d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$ (3.5)
+(3.5) $d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$
 Proof.
@@
 $\mathfrak{a}_{\lambda,t}=\sum_{(1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}.$
-The goal of this section is to justify the term geodesic, in the sense that it computes the distance between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$.
-In other words, it is a distance-minimizing continuous curve with respect to the metric $d_{1}$.
-We first give a formula for the geodesic between two saturated filtrations, and use it to show that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}^{s}$ when $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$..
+The goal of this subsection is to justify the term geodesic, in the sense that it computes the distance between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$.
+In other words, it is a distance-minimizing, continuous curve with respect to the metric $d_{1}$.
+We first give a formula for the geodesic between two saturated filtrations, and use it to show that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}^{s}$ when $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$.
 Proposition 3.4.
 Let $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$. Then for any $t\in(0,1)$, we have
-$\mathfrak{a}_{\bullet,t}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{c_{t,v}}),$ (3.6)
+(3.6) $\mathfrak{a}_{\bullet,t}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{c_{t,v}}),$
 where
@@
 Since $\mathfrak{a}_{\lambda,0}\cap\mathfrak{a}_{\lambda,1}\subset\mathfrak{a}_{\lambda,s}$ for any $s\in[0,1]$, it suffices to show that for any $(\mu,\nu)\in\mathbb{R}_{\geq 0}^{2}$ satisfying $(1-t^{\prime})\mu+t^{\prime}\nu=\lambda$ and $\mu>\lambda$, we have
-$\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\lambda,t}.$ (3.7)
+(3.7) $\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\lambda,t}.$
 To see this, let $\mu^{\prime}:=\frac{(1-t^{\prime})\mu+(t^{\prime}-t)\nu}{1-t}$, then $(1-t)\mu+t\nu=(1-t^{\prime})\mu+t^{\prime}\nu=\lambda$, and we have
@@
 Let $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$ and let $\{\mathfrak{a}_{\bullet,t}\}_{t\in[0,1]}$ be the geodesic between them. Then for any $0\leq t<t^{\prime}\leq 1$, we have
-$d_{1}(a_{\bullet,0},\mathfrak{a}_{\bullet,t^{\prime}})=d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{a}_{\bullet,t^{\prime}}).$ (3.8)
+(3.8) $d_{1}(a_{\bullet,0},\mathfrak{a}_{\bullet,t^{\prime}})=d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{a}_{\bullet,t^{\prime}}).$
 Proof.
@@
 We claim that for any $\lambda\in\mathbb{R}_{>0}$,
-$\mathfrak{a}_{\lambda,t}/\mathfrak{b}_{\lambda}\cong\mathfrak{c}_{\lambda}/\mathfrak{b}^{\prime}_{\lambda}.$ (3.9)
+(3.9) $\mathfrak{a}_{\lambda,t}/\mathfrak{b}_{\lambda}\cong\mathfrak{c}_{\lambda}/\mathfrak{b}^{\prime}_{\lambda}.$
 Granting the claim for now, we know that $d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet,t})=d_{1}(\mathfrak{b}^{\prime}_{\bullet},\mathfrak{c}_{\bullet})$. So we get
@@
 3.2.2. Duistermaat-Heckman measures and continuity
-In this section $R$ contains a field $\mathbbm{k}$. We briefly recall the measure $\mu$ introduced in [BLQ24, Section 4] encoding multiplicities and its properties.
+In this subsection, we assume that $R$ contains a field $\mathbbm{k}$. We first briefly recall the measure $\mu$ introduced in [BLQ24, Section 4] encoding multiplicities and its properties.
 Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$. Define the function $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ by
@@
 Proof.
-The first two statements are [BLQ24, Proposition 4.7] and the last two are [BLQ24, Proposition 4.9]
+The first two statements are [BLQ24, Proposition 4.7] and the last two are [BLQ24, Proposition 4.9]. ∎
+For a rectangle, the measure $\mu$ is controlled by the Lebesgue measure.
 Lemma 3.9.
 Fix $(a,b)\in\mathbb{R}^{2}$.
-Let $A:=[a+\alpha]\times[b+\beta]\subset\mathbb{R}^{2}$ be a rectangle, where $\alpha,\beta\in\mathbb{R}_{>0}$.
+Let $A:=[a,a+\alpha]\times[b,b+\beta]\subset\mathbb{R}^{2}$ be a rectangle, where $\alpha,\beta\in\mathbb{R}_{>0}$.
 Then there exists $M:=M(a,b)\in\mathbb{R}_{>0}$ such that $\mu(A)\leq M\alpha\beta$ when $\alpha,\beta\leq 1$.
@@
 where the second and the third equality follow from Lemma 3.8(3), and the first inequality follows from Lemma 3.8(2).
-Hence proof is finished with $M:=n^{2}(a+1)^{n-1}(b+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})$.
+Hence the proof is finished with $M:=n^{2}(a+1)^{n-1}(b+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})$.
 ∎
@@
 Suppose $R$ contains a field.
-Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$, the map $[0,1]\to\mathrm{Fil}$, $t\mapsto\mathfrak{a}_{\bullet,t}$ is Lipschitz continuous, where $[0,1]$ is equipped with the Euclidean topology and $\mathrm{Fil}$ is equipped with the topology induced by the pseudometric $d_{1}$.
+Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$, the map $0,1]\to\mathrm{Fil}$, $t\mapsto\mathfrak{a}_{\bullet,t}$ is Lipschitz continuous, where $[0,1]$ is equipped with the Euclidean topology and $\mathrm{Fil}$ is equipped with the topology induced by the pseudometric $d_{1}$.
 Proof.
 Given $0\leq t_{1}<t_{2}\leq 1$, denote $\Delta_{j}:=\{(x,y)\in\mathbb{R}^{2}_{>0}\mid(1-t_{j})x+t_{j}y\leq 1\}$ for $j=1,2$. An easy calculation using Lemma 3.8(1) gives
-$d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})=\mu(\Delta_{1}\backslash\Delta_{2})+\mu(\Delta_{2}\backslash\Delta_{1}).$ (3.10)
-First assume that $10<t_{1}<t_{2}<$ 1. By Lemma 3.8(4) we have
+(3.10) $d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})=\mu(\Delta_{1}\backslash\Delta_{2})+\mu(\Delta_{2}\backslash\Delta_{1}).$
+First assume that $0<t_{1}<t_{2}<1$. By Lemma 3.8(4) we have
 $\mu(\Delta_{1}\backslash\Delta_{2})=\mu((\Delta_{1}\backslash\Delta_{2})\cap\{y\leq Dx\})=\mu(\Delta)$
@@
 $\mathrm{Area}(\Delta)\leq S:=\frac{D(D-1)(t_{2}-t_{1})}{(1-t_{1}+t_{1}D)(1-t_{2}+t_{2}D)}\leq D(D-1)(t_{2}-t_{1}).$
-Note that $(x,y)\in\Delta$ satisfies $x\leq 1$ and $y\leq D$.
-So $\Delta$ can be covered by finitely many rectangles $R_{j}$ of the form $[a_{j},a_{j}^{\prime}]\times[b_{j},b_{j}^{\prime}]$, satisfying
+Note that $(x,y)\in\Delta$ satisfies $x\leq 1$ and $y\leq D$. So $\Delta$ can be covered by finitely many rectangles $R_{j}$ of the form $a_{j},a_{j}^{\prime}]\times[b_{j},b_{j}^{\prime}]$, satisfying
 $a_{j}\leq 1,\ b_{j}\leq D,\text{ and }\sum_{j}\mathrm{Area}(R_{j})\leq 2S.$
@@
 Similarly $\mu(\Delta_{1}\backslash\Delta_{1})$ satisfies the same estimate. Hence by (3.10), we have
-$d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n+1}(D+1)^{n-1}D(D-1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})(t_{2}-t_{1}).$ (3.11)
+(3.11) $d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n+1}(D+1)^{n-1}D(D-1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})(t_{2}-t_{1}).$
 Now assume that $t_{1}=0$. We may assume $t_{2}<\frac{1}{2}$. We have
@@
 where the second inequality follows from Lemma 3.9. Hence
-$d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n}(2^{n-1}(D-1)^{2}+1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})t_{2}.$ (3.12)
+(3.12) $d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n}(2^{n-1}(D-1)^{2}+1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})t_{2}.$
 The case when $t_{2}=1$ can be treated similarly. Combining the cases together, we get
@@
 3.3. Continuity along monotonic sequences
-In this section we prove some convergence result along monotonic results.
+In this subsection we prove some convergence results along monotonic sequences.
+The results are partially inspired by the study of the completion of certain metric spaces in the global setting, for example, [Dar17, Xia23, BJ25].
 3.3.1. Okounkov bodies and continuity along increasing sequences
@@
 $\lim_{m\to\infty}\frac{\#\Gamma_{m}}{m^{n}}=\mathrm{vol}(\Delta(\Gamma)).$
+We need the following formula, which relates the co-length of a primary ideal to the points of its semigroups.
 Proposition 3.13.
@@
 Then there exists $t=t(R)\in\mathbb{Z}_{>0}$ depending only on $R$, and $2t$ semigroups $\bar{\Gamma}^{(j)},\Gamma^{(j)}\subset\mathbb{Z}_{\geq 0}^{n+1}$, $j=1,\ldots,t$ satisfying the conditions of Lemma 3.12, such that
-$\ell(R/\mathfrak{a}_{m})=\sum_{j=1}^{t}(\#\bar{\Gamma}^{(j)}_{m}-\#\Gamma^{(j)}_{m}).$ (3.13)
+(3.13) $\ell(R/\mathfrak{a}_{m})=\sum_{j=1}^{t}(\#\bar{\Gamma}^{(j)}_{m}-\#\Gamma^{(j)}_{m})$.
 Proof.
@@
 Denote $\kappa^{\prime}:=R^{\prime}/\mathfrak{n}$, where $\mathfrak{n}$ is the maximal ideal of $R^{\prime}$, then we know that $t:=[\kappa^{\prime}:\kappa]\in\mathbb{Z}_{>0}$.
-We proceed to define a valuation $v$ on the quotient field $K:=Q(R)=Q(R^{\prime})$.
-Fix a regular system of parameters $(x_{1},\ldots,x_{n})$ of $R^{\prime}$ and $\alpha_{1},\ldots,\alpha_{n}\in\mathbb{R}_{>1}$, which are linearly independent over $\mathbb{Q}$.
-By Cohen’s structure theorem, the $\mathfrak{n}$-completion $(\widehat{R^{\prime}},\hat{\mathfrak{n}})$ of $R^{\prime}$ has a coefficient ring $C$ as a regular local ring.
-Hence for any $f\in R^{\prime}$ and $d\in\mathbb{Z}_{>0}$, we can write
+[…]
 $f=\sum_{i_{1}+\ldots i_{n}<d}c_{i_{1},\ldots,i_{n}}x_{1}^{i_{1}}\cdots x_{n}^{i_{n}}+g_{d}\in\widehat{R^{\prime}},$
@@
 $v(f):=\min\{\sum_{j=1}^{n}\alpha_{j}i_{j}\mid c_{i_{1},\ldots,i_{n}}\neq 0\}$
-is independent on the choice of the units and $d$, hence this defines a function $v:K^{\times}\to\mathbb{R}$ by $v(f/g)=v(f)-v(g)$ for any $f,g\in R^{\prime}$.
+is independent of the choice of the units and $d$, hence this defines a function $v:K^{\times}\to\mathbb{R}$ by v(f/g)=v(f)v(g)v(f/g)=v(f)-v(g) for any $f,g\in R^{\prime}$.
 One can check that $v$ is a real valuation.
 Denote its value group by $\Gamma_{v}:=v(K^{\times})$ and its valuation ring by $(\mathcal{O}_{v},\mathfrak{m}_{v})$ and for any $\lambda\in\mathbb{R}_{\geq 0}$, denote
-$\mathfrak{b}_{\lambda}:=\{f\in\mathcal{O}_{v}\mid v(f)\geq\lambda\}\text{ and }\mathfrak{b}_{>\lambda}:=\{f\in\mathcal{O}_{v}\mid v(f)>\lambda\}.$
+[…]
 Since $\alpha_{j}$ are linearly independent over $\mathbb{Q}$, we know that
@@
 Then for any $m\in\mathbb{Z}_{>0}$ we have
-$\mathfrak{b}_{cm}\cap R\subset\mathfrak{m}^{c_{1}m}\subset\mathfrak{a}_{1}^{m}\subset\mathfrak{a}_{m}.$ (3.14)
+(3.14) $\mathfrak{b}_{cm}\cap R\subset\mathfrak{m}^{c_{1}m}\subset\mathfrak{a}_{1}^{m}\subset\mathfrak{a}_{m}.$
 To simplify the notation, for $\bm{\beta}:=(\beta_{1},\ldots,\beta_{n})\in\mathbb{R}_{\geq 0}^{n}$, we denote $\xi(\bm{\beta}):=\sum_{i=1}^{n}\alpha_{i}\beta_{i}$ and $|\bm{\beta}|:=\sum_{i=1}^{n}\beta_{i}$.
@@
 This proves (3.13) and finishes the proof. ∎
+It is not hard to see that taking the closed convex cone commutes with taking the union.
 Lemma 3.14.
@@
 The last equality follows immediately. ∎
+Now we are ready to show that the multiplicity is continuous along increasing sequences.
+The reader may want to compare the following result with similar statements in the global setting, for example, [Dar17, Proposition 6.11] and [BJ25, Lemma 4.7].
 Proposition 3.15.
@@
 Let $\Gamma^{(j)}$ and $\bar{\Gamma}^{(j)}$ be the semigroups of $\mathfrak{a}_{\bullet}$.
-By the proof of Proposition 3.13 it is easy to see that $\bar{\Gamma}_{,k}^{(j)}$ is indeed independent on $k$, that is, $\bar{\Gamma}_{,k}^{(j)}=\bar{\Gamma}^{(j)}$ for any $1\leq j\leq t$ and $k\in\mathbb{Z}_{>0}$.
+By the proof of Proposition 3.13 it is easy to see that $\bar{\Gamma}_{,k}^{(j)}$ is indeed independent of $k$, that is, $\bar{\Gamma}_{,k}^{(j)}=\bar{\Gamma}^{(j)}$ for any $1\leq j\leq t$ and $k\in\mathbb{Z}_{>0}$.
 We claim that for any $1\leq j\leq t$,
-$\Gamma^{(j)}=\cup_{k}\Gamma_{,k}^{(j)}.$ (3.15)
+(3.15) $\Gamma^{(j)}=\cup_{k}\Gamma_{,k}^{(j)}.$
 Indeed, by definition $\Gamma_{,k}^{(j)}\subset\Gamma^{(j)}$ for any $k$ and $j$, so $\cup_{k}\Gamma_{,k}^{(j)}\subset\Gamma^{(j)}$.
@@
 $\dim_{\kappa}\mathfrak{a}_{m,K}\cap\mathfrak{b}_{\xi(\bm{\beta})}/\mathfrak{a}_{m,K}\cap\mathfrak{b}_{>\xi(\bm{\beta})}\geq j.$
-This show that $(\bm{\beta},m)\in\Gamma_{,K}^{(j)}$. Thus we get $\Gamma^{(j)}\subset\cup_{k}\Gamma_{,k}^{(j)}$ and finishes the proof of the claim.
+This shows that $(\bm{\beta},m)\in\Gamma_{,K}^{(j)}$. Thus we get $\Gamma^{(j)}\subset\cup_{k}\Gamma_{,k}^{(j)}$ and the proof of the claim is finished.
 Now we can apply Lemma 3.12, Proposition 3.13 and Lemma 3.14 to get
@@
 $=$ $n!\sum_{j=1}^{t}\lim_{k\to\infty}(\mathrm{vol}(\Delta(\Gamma^{(j)}))-\mathrm{vol}(\Delta(\Gamma_{,k}^{(j)})$
-$=$ $0$
+$=0.$
 This finishes the proof. ∎
@@
 Example 3.17.
-Let $R=k[\![x]\!]$ with $\mathfrak{m}=(x)$. For $k\in\mathbb{Z}_{>0}$, define $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}$ by
+Let R=k[[x]]R=k[\![x]\!] with $\mathfrak{m}=(x)$. For $k\in\mathbb{Z}_{>0}$, define $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}$ by
 $\mathfrak{a}_{\lambda,k}:=\left\{\begin{aligned} &\mathfrak{m}^{\lceil 2\lambda\rceil},&\lambda\leq k,\\ &\mathfrak{m}^{2k},&k<\lambda\leq 2k,\\ &\mathfrak{m}^{\lceil\lambda\rceil},&\lambda>2k.\end{aligned}\right$.
@@
 Proposition 3.18.
-Let $x\in X$ be a klt singularity over a field of characteristic 00.
+Let $x\in X$ be a klt singularity over a field of characteristic 0.
 Let $v_{k}\in\mathrm{Val}_{X,x}$, $k\in\mathbb{Z}_{>0}$ be a sequence of valuations such that $A_{X}(v_{k})=1$ and $\widehat{\mathrm{vol}}_{x,X}(v_{k})<V$ for some $V>0$.
@@
 is sequentially compact under the weak topology. Hence there is a subsequence of $\{v_{k}\}$ which converges to some $v\in\mathrm{Val}_{X,x}$. From now on, we assume that $v_{k}\to v$ weakly.
-For $k,l\in\mathbb{Z}_{>0}$, define
-$\mathfrak{a}_{\bullet,k}^{(l)}:=\bigcap_{j=k}^{k+l}\mathfrak{a}_{\bullet}(v_{j}).$
-Then for any $k\in\mathbb{Z}_{>0}$, $\{\mathfrak{a}_{\bullet,k}^{(l)}\}_{l\in\mathbb{Z}_{>0}}$ is a decreasing sequence of filtrations. Let
-$\mathfrak{a}_{\bullet,k}:=\cap_{l}\mathfrak{a}_{\bullet,k}^{(l)}\subset\mathfrak{a}_{\bullet}(v_{k})$
-Clearly the sequence of filtrations $\{\mathfrak{a}_{\bullet,k}\}_{k\in\mathbb{Z}_{>0}}$ is increasing.
-Let $\chi_{k}=\mathrm{ord}_{\mathfrak{a}_{\bullet,k}}$ be the corresponding norm.
-By definition, $\chi_{k}=\inf_{j\geq k}v_{j}\leq v$.
+Let $\chi_{k}\coloneqq\inf_{j\geq k}v_{j}\leq v$.
 Since $v_{j}\to v$ weakly, given any $f\in\mathfrak{m}$ and any $\epsilon>0$, there exists $J\in\mathbb{Z}_{>0}$ such that $v_{j}(f)>v(f)-\epsilon$ for any $j\geq J$.
@@
 $v(f)\geq\chi_{k}(f)=\inf_{j\geq k}v_{j}(f)\geq v(f)-\epsilon$
-as long as $k\geq J$. This proves $\chi_{k}\to v$ weakly as an increasing sequence. So we may apply Proposition 3.15 to get $\lim_{k\to\infty}d_{1}(\chi_{k},v)=0$, and the proof is finished. ∎
+as long as $k\geq J$.
+This proves $\chi_{k}\to v$ weakly as an increasing sequence.
+So we may apply Proposition 3.15 to get $\lim_{k\to\infty}d_{1}(\chi_{k},v)=0$.
+By Proposition 2.20, $\mathfrak{a}_{\bullet,k}\coloneqq\mathfrak{a}(\chi_{k})=\cap_{j\geq k}\mathfrak{a}(v_{j})$ is saturated, and the proof is finished.
+∎
 3.3.2. Normalized volumes and continuity along decreasing sequences
@@
 Lemma 3.21.
-(c.f. [Blu18, Proposition 3.7]) Let $\epsilon,V>0$ and $C\in\mathbb{Z}_{>1}$ be constants. Then for there exists $\Lambda=\Lambda(C,V,\epsilon)\in\mathbb{R}_{>0}$ such that the following holds.
+(c.f. [Blu18, Proposition 3.7]) Let $\epsilon,V>0$ and $C\in\mathbb{Z}_{>1}$ be constants. Then there exists $\Lambda=\Lambda(C,V,\epsilon)\in\mathbb{R}_{>0}$ such that the following holds.
 Let $\mathfrak{a}_{\bullet}=\cap_{i\in I}\mathfrak{a}_{\bullet}(v_{i})$, where $\{v_{i}\}_{i\in I}\subset\mathrm{Val}^{+}_{X,x}$ is a set of valuations such that
@@
 By the valuative characterization of multiplier ideals [BdFFU15, Theorem 1.2], we have
-$\mathcal{J}(X;\lambda\cdot\mathfrak{a}_{\bullet})\subset\mathfrak{a}_{\lambda-E},$ (3.16)
+(3.16) $\mathcal{J}(X;\lambda\cdot\mathfrak{a}_{\bullet})\subset\mathfrak{a}_{\lambda-E},$
 where $E:=\frac{c_{0}CV\mathrm{lct}(X;\mathfrak{m})}{\widehat{\mathrm{vol}}(x,X)}>0$ depends only on $C$ and $V$. Replacing the inclusion
@@
 3.4. The toric case and Newton-Okounkov bodies
-In this section, we work over a field $\mathbbm{k}$ and follow the notation in [Ful93] for toric varieties.
-Our goal is to identify the subspace $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ of saturated monomial filtrations on a toric singularity with a subspace of the Fréchet-Nikodym-Aronszyan metric on cobounded sets of the dual cone.
-See [KK14, Section 6] and [Blu18, Section 8] for some relevant constructions.
+We deal with the toric case in this subsection.
+Throughout, we work over an algebraically closed field $\mathbbm{k}$ of characteristic 0 and follow the notation in [Ful93] for toric varieties.
+Our goal is to identify the subspace $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ of saturated monomial filtrations on a toric singularity with a subspace of the Fréchet-Nykodym-Aronszajn metric on cobounded sets of the dual cone.
+See [JM12, Section 8], [KK14, Section 6] and [Blu18, Section 8] for some relevant constructions.
 Let $N$ be a free abelian group of rank $n\geq 1$ and $M=N^{*}$ its dual.
@@
 Let $R$ be the local ring of $X_{\sigma}$ at $x$ and $\mathfrak{m}$ its maximal ideal.
-Recall that there is a $1$-to-$1$ correspondence between toric valuations $v_{u}\in\mathrm{Val}^{\mathrm{toric}}_{R}$ and $u\in\sigma$, given by
+Recall that there is a 11-to-11 correspondence between toric valuations $v_{u}\in\mathrm{Val}^{\mathrm{toric}}_{R}$ and $u\in\sigma$, given by
 $v_{u}(\sum_{m\in\sigma^{\vee}\cap M}c_{m}\chi^{m})=\min\{\langle u,m\rangle\mid c_{m}\neq 0\},$
@@
 $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$).
-Given an $\mathfrak{m}$-primary monomial ideal $\mathfrak{a}$, the Newton polyhedron $P(\mathfrak{a})$ of $\mathfrak{a}$ is the convex hull of $\{m\in\sigma^{\vee}\cap M\mid\chi^{m}\in\mathfrak{a}\}$.
-Moreover, we have $\mathrm{e}(\mathfrak{a})=\mathrm{vol}(\sigma^{\vee}\backslash P(\mathfrak{a}))$.
-Recall that the Newton-Okounkov body $P(\mathfrak{a}_{\bullet})$ of a monomial $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$ is defined to be
+[…]
 $P(\mathfrak{a}_{\bullet})\coloneqq\overline{\cup_{\lambda\in\mathbb{R}_{>0}}\{m/\lambda\mid m\in P(\mathfrak{a}_{\lambda})\}}\subset\sigma^{\vee},$
-which is a convex set such that $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})$ is bounded. It is known that the covolume of the Newton-Okounkov body computes the multiplicity of the filtration.
+[…]
 Lemma 3.22.
 [KK14, Theorem 6.5] Let $\mathfrak{a}_{\bullet}$ be a monomial $\mathfrak{m}$-filtration on $R$. Then we have
-$\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{vol}(\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})).$
+[…]
 The next lemma gives an alternative characterization for $P(\mathfrak{a}_{\bullet})$.
 Lemma 3.23.
-For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$, $P(\mathfrak{a}_{\bullet})=\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
+For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$, $P(\mathfrak{a}_{\bullet})=\cap_{u\in\mathrm{Int}(\sigma)\cap N}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
 Proof.
-For $m\in\sigma^{\vee}\cap M$ with $\chi^{m}\in\mathfrak{a}_{\lambda}$, we have $\langle u,m\rangle=v_{u}(m)\geq v_{u}(\mathfrak{a}_{\lambda})\geq\lambda v(\mathfrak{a}_{\bullet})$ by definition, that is, $m/\lambda\in H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
-Hence $P(\mathfrak{a}_{\bullet})\subset\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
-Conversely, assume $p\in\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
-Assume $p\notin P(\mathfrak{a}_{\bullet})$.
-Since $P(\mathfrak{a}_{\bullet})$ is convex, there exists $u\in N_{\mathbb{R}}$ and $b\in\mathbb{R}_{>0}$ such that $\langle u,p\rangle<b$ but $P(\mathfrak{a}_{\bullet})\subset H_{u}(b)$.
-Since $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})$ is bounded, we know that $u\in\mathrm{Int}(\sigma)$.
-Now $b>\langle u,p\rangle\geq v_{u}(\mathfrak{a}_{\bullet})$ by the choice of $p$.
-Hence there exists $\lambda\in\mathbb{R}_{>0}$ and $\chi^{m}\in\mathfrak{a}_{\lambda}$ such that $v_{u}(\chi^{m})=v_{u}(\mathfrak{a}_{\lambda})<\lambda\cdot b$.
-In particular, $m/\lambda\in P(\mathfrak{a}_{\bullet})\backslash H_{u}(b)$, a contradiction.
-So $\cap_{u\in\mathrm{Int}(\sigma)}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))\subset P(\mathfrak{a}_{\bullet})$ and the proof is finished.
-∎
-Next we show that the saturation of a monomial filtration can be computed using only toric valuations.
+For $m\in\sigma^{\vee}\cap M$ with $\chi^{m}\in\mathfrak{a}_{\lambda}$, we have $\langle u,m\rangle=v_{u}(\chi^{m})\geq v_{u}(\mathfrak{a}_{\lambda})\geq\lambda v(\mathfrak{a}_{\bullet})$ by definition, that is, $m/\lambda\in H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
+Hence $P(\mathfrak{a}_{\bullet})\subset\cap_{u\in\mathrm{Int}(\sigma)\cap N}H_{u}(v_{u}(\mathfrak{a}_{\bullet}))$.
+[…]
+Next we show that the saturation of a monomial filtration can be computed using only toric divisorial valuations.
 Lemma 3.24.
 A monomial filtration $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{\mathrm{mon}}_{R,\mathfrak{m}}$ is saturated if and only if it is of the form
-$\mathfrak{a}_{\bullet}=\cap_{u\in\mathrm{Int}(\sigma)}\mathfrak{a}_{\bullet}(v_{u}/a_{u})$ (3.17)
-where $a_{u}:\mathrm{Int}(\sigma)\to\mathbb{R}_{>0}$ is a homogeneous function.
+(3.17) $\mathfrak{a}_{\bullet}=\cap_{u\in\mathrm{Int}(\sigma)\cap N}\mathfrak{a}_{\bullet}(v_{u}/a_{u})$
+where $a_{u}:\mathrm{Int}(\sigma)\to\mathbb{R}_{>0}$ is a homogeneous function for any $u\in\mathrm{Int}(\sigma)\cap N$.
 Proof.
@@
 Conversely, assume that $\mathfrak{a}_{\bullet}$ is saturated. Then we know that
-$\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\coloneqq\cap_{u\in\mathrm{Int}(\sigma)}\mathfrak{a}_{\bullet}(\frac{v_{u}}{v_{u}(\mathfrak{a}_{\bullet})})$
+$\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\coloneqq\cap_{u\in\mathrm{Int}(\sigma)\cap N}\mathfrak{a}_{\bullet}(\frac{v_{u}}{v_{u}(\mathfrak{a}_{\bullet})})$
 by Lemma 2.19.
-Moreover, $v_{u}(\mathfrak{a}_{\bullet}^{\prime})=v_{u}(\mathfrak{a}_{\bullet})$ since $\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\subset\mathfrak{a}_{\bullet}(v_{u}/v_{u}(\mathfrak{a}_{\bullet}))$.
+Moreover, $v_{u}(\mathfrak{a}_{\bullet}^{\prime})=v_{u}(\mathfrak{a}_{\bullet})$ since $\mathfrak{a}_{\bullet}\subset\mathfrak{a}^{\prime}_{\bullet}\subset\mathfrak{a}_{\bullet}(v_{u}/v_{u}(\mathfrak{a}_{\bullet}))$ for any $u\in\mathrm{Int}(\sigma)\cap N$.
 So $P(\mathfrak{a}_{\bullet})=P(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.23, and hence $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{a}^{\prime}_{\bullet})$ by Lemma 3.22.
-Now $\mathfrak{a}^{\prime}_{\bullet}$ is saturated by Proposition 2.20 again, so by Theorem 2.18, we get $\mathfrak{a}_{\bullet}=\mathfrak{a}^{\prime}_{\bullet}$.
+Now $\mathfrak{a}^{\prime}_{\bullet}$ is saturated by Proposition 2.20 again, and so by Theorem 2.18, we get $\mathfrak{a}_{\bullet}=\mathfrak{a}^{\prime}_{\bullet}$.
 ∎
@@
 Then
-$v_{u}(\cap_{i\in I}\mathfrak{a}_{\bullet,i})=\sup_{i\in I}v_{u}(\mathfrak{a}_{\bullet,i})$ (3.18)
+(3.18) $v_{u}(\cap_{i\in I}\mathfrak{a}_{\bullet,i})=\sup_{i\in I}v_{u}(\mathfrak{a}_{\bullet,i})$
 for any $u\in\mathrm{Int}(\sigma)$. As a result, we have $P(\cap_{i\in I}\mathfrak{a}_{\bullet,i})=\cap_{i\in I}P(\mathfrak{a}_{\bullet,i})$.
@@
 $C\coloneqq\{p\}+(\sigma^{\vee}\backslash H_{w}(\delta c^{\prime}_{w}))\subset\sigma^{\vee},$
-where $+$ denotes the Minkowski sum.
+where ++ denotes the Minkowski sum.
 Take $p_{i}\in C\cap M_{\mathbb{Q}}$ such that $p_{i}\to p$ and $k_{i}\in\mathbb{Z}_{>0}$ such that $k_{i}p_{i}\in M$.
@@
 $C\coloneqq\{p\}+(\sigma^{\vee}\backslash H_{w}(\delta a_{w}))\subset\sigma^{\vee}$
-where $+$ denotes the Minkowski sum.
+where ++ denotes the Minkowski sum.
 Then there exists $p_{j}\in C\cap M_{\mathbb{Q}}$ such that $p_{j}\to p$.
@@
 Then for any $u\in\mathrm{Int}(\sigma)$, we have $\langle u,p\rangle\leq\langle u,p_{j}\rangle$ for any $u\in\mathrm{Int}(\sigma)$ and
-$\langle w,p_{j}\rangle<a_{w}(1-2\delta)+\delta a_{w}=a_{w}(1-\delta).$ (3.19)
+(3.19) $\langle w,p_{j}\rangle<a_{w}(1-2\delta)+\delta a_{w}=a_{w}(1-\delta).$
 By the first inequality, for any $u\in\mathrm{Int}(\sigma)$ we have
@@
 Let $\mathcal{P}(\sigma^{\vee})\coloneqq\{P\subsetneq\sigma^{\vee}\mid P\text{ is closed and convex, }\sigma^{\vee}\backslash P\text{ is bounded}\}$.
-We prove that taking Newton-Okounkov body gives a bijection between $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ and $\mathcal{P}(\sigma^{\vee})$.
+We prove that taking the Newton-Okounkov body gives a bijection between $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ and $\mathcal{P}(\sigma^{\vee})$.
+Note that the following lemma also justifies the statement that “if $P(\mathfrak{a}_{\bullet})=P(\mathfrak{b}_{\bullet})$, then $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ should be regarded as equisingular” in [JM12, Section 8].
 Lemma 3.26.
 $P:\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}\to\mathcal{P}(\sigma^{\vee})$ is a bijection, whose inverse is given by $P\mapsto\mathfrak{a}_{\bullet}(P)$, where
-$\mathfrak{a}_{\lambda}(P)\coloneqq\mathrm{span}\{\chi^{m},m\in\sigma^{\vee}\cap M\mid\langle m,u\rangle\geq\lambda\cdot\inf_{p\in P}\langle p,u\rangle\text{ for any }u\in\mathrm{Int}(\sigma)\},$ (3.20)
+(3.20) $\mathfrak{a}_{\lambda}(P)\coloneqq\mathrm{span}\{\chi^{m},m\in\sigma^{\vee}\cap M\mid\langle m,u\rangle\geq\lambda\cdot\inf_{p\in P}\langle p,u\rangle\text{ for any }u\in\mathrm{Int}(\sigma)\},$
 Proof.
@@
 $P(\mathfrak{a}_{\bullet})\supset P(\mathfrak{m}^{d\bullet)}=P(\widetilde{\mathfrak{m}^{d\bullet}})=\cap_{i=1}^{k}P(\mathfrak{a}_{\bullet}(v_{u_{i}})=\cap_{i=1}^{k}H_{u_{i}}(1).$
-Thus $\sigma^{\vee}\backslash P(\mathfrak{a}_{\bullet})\subset\cup_{i=1}^{k}\sigma^{\vee}\backslash H_{u_{i}}(1)$, which is bounded.
-This proves $P:\mathrm{Fil}^{s,\mathrm{mon}}_{X,x}\to\mathcal{P}(\sigma^{\vee})$.
+[…]
 Now let $P\in\mathcal{P}(\sigma^{\vee})$.
@@
 By Lemma 3.26, $P$ is a bijection.
-For simplicity, denote $\mathrm{covol}(P)\coloneqq\mathrm{vol}(\sigma^{\vee}\backslash P)$ below. For $\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, we have
+[…]
 $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=$ $2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet})$
@@
 $=$ $2\mathrm{covol}(P(\mathfrak{a}_{\bullet})\cap P(\mathfrak{b}_{\bullet}))-\mathrm{covol}(P(\mathfrak{a}_{\bullet}))-\mathrm{covol}(P(\mathfrak{b}_{\bullet}))$
-$=$ $\mathrm{vol}(P(\mathfrak{a}_{\bullet})\backslash P(\mathfrak{b}_{\bullet}))+\mathrm{vol}(P(\mathfrak{b}_{\bullet})\backslash P(\mathfrak{a}_{\bullet}))$
+[…]
 $=$ $\mathrm{vol}(P(\mathfrak{a}_{\bullet})\Delta P(\mathfrak{b}_{\bullet}))=d(P(\mathfrak{a}_{\bullet}),P(\mathfrak{b}_{\bullet})),$
@@
 4. Supnorm metrics on the space of filtrations
+In this section, we study the basic properties of $d_{\infty}$, prove the local Lipschitz continuity of log canonical thresholds, and compare several different topologies on the filtration spaces.
 4.1. Supnorm metrics and homogenization
-Throughout this section we will use the language of norms.
+Throughout this subsection, we will use the language of norms. Recall that $\mathcal{N}$ is the set of linearly bounded $\mathfrak{m}$-norms on $R$.
 Definition 4.1.
-Fix a norm $\rho\in\mathcal{N}^{h}$. We define a function $d_{\infty,\rho}:\mathcal{N}\times\mathcal{N}\to\mathbb{R}_{\geq 0}$ by
+Fix a norm $\rho\in\mathcal{N}$. We define a function $d_{\infty,\rho}:\mathcal{N}\times\mathcal{N}\to\mathbb{R}_{\geq 0}$ by
 $d_{\infty,\rho}(\chi,\chi^{\prime}):=\limsup_{\lambda\to\infty}\sup_{\rho(f)\geq\lambda}\left\{\frac{|\chi(f)-\chi^{\prime}(f)|}{\rho(f)}\right\}.$
 We will denote $d_{\infty,\chi_{0}}=:d_{\infty}$.
+It is not hard to see that the $d_{\infty}$-topology is independent of the choice of the reference norm.
 Lemma 4.2.
@@
 By definition, $d_{\infty,\rho}$ is symmetric and non-negative for any $\rho$.
-Given $\chi,\chi^{\prime},\chi^{\prime\prime}\in\mathcal{N}$, since $|\chi(f)-\chi^{\prime\prime}(f)|\leq|\chi(f)-\chi^{\prime}(f)|+|\chi^{\prime}(f)+\chi^{\prime\prime}(f)|$, the triangle inequality holds.
+Given $\chi,\chi^{\prime},\chi^{\prime\prime}\in\mathcal{N}$, since $|\chi(f)-\chi^{\prime\prime}(f)|\leq|\chi(f)-\chi^{\prime}(f)|+|\chi^{\prime}(f)-\chi^{\prime\prime}(f)|$, the triangle inequality holds.
 This proves that $d_{\infty,\rho}$ is a pseudometric.
-The second claim follows from the fact that given $\rho,\rho^{\prime}\in\mathcal{N}$, there exists $M\in\mathbb{R}_{>0}$ such that $M^{-1}\rho\leq\rho^{\prime}\leq M\rho$. ∎
+The second claim follows from the fact that given $\rho,\rho^{\prime}\in\mathcal{N}$, there exists $M\gg 0$ such that $M^{-1}\rho\leq\rho^{\prime}\leq M\rho$. ∎
+On $\mathcal{N}^{h}$, we do not need to take $\limsup$ in the definition of $d_{\infty}$.
 Lemma 4.3.
@@
 By Lemma 4.3, we may assume that $d=\sup\frac{\widehat{\chi}(f)-\widehat{\chi}^{\prime}(f)}{\chi_{0}(f)}$.
-For any $\epsilon>0$, there exits $f\in\mathfrak{m}$ such that
+For any $\epsilon>0$, there exists $f\in\mathfrak{m}$ such that
 $(d-\epsilon)\chi_{0}(f)<\widehat{\chi}(f)-\widehat{\chi}^{\prime}(f)=\lim_{k\to\infty}\frac{\chi(f^{k})}{k}-\lim_{k\to\infty}\frac{\chi^{\prime}(f^{k})}{k}.$
@@
 since $\chi_{0}$ is homogeneous. Similar to the proof of Lemma 4.3 we get $d_{\infty}(\chi,\chi^{\prime})\geq d$. ∎
+The following lemma shows that $(\mathcal{N}^{h},d_{\infty})$ is complete.
 Lemma 4.5.
@@
 ∎
-4.2. Locally Lipschitz continuity of log canonical thresholds
-In this section we prove that the log canonical thresholds is locally Lipschitz with respect to the $d_{\infty}$-topology.
+4.2. Local Lipschitz continuity of log canonical thresholds
+In this subsection we prove that the log canonical threshold is locally Lipschitz continuous with respect to the $d_{\infty}$-topology.
 Proposition 4.6.
@@
 Proof.
-Since $\mathrm{lct}(\chi)=\mathrm{lct}(\widehat{\chi})$ for any $\chi\in\mathcal{N}$, by Lemma 4.4, we may assume that $\chi,\chi^{\prime}\in\mathcal{N}^{h}$.
-Fix $c=c(\chi):=\inf_{f\in\mathfrak{m}}\frac{\chi(f)}{\chi_{0}(f)}>0$ and denote $a:=\mathrm{lct}(\chi)>0$.
+[…]
 Assume that $d_{\infty}(\chi,\chi^{\prime})<\delta$ for some $\delta\in(0,c)$. By Lemma 4.3, we know that for any $f\in R$,
@@
 Now we compute
-$\frac{v(f)}{\chi^{\prime}(f)}=$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\chi(f)-\chi^{\prime}(f)}{\chi^{\prime}(f)}\right)$ (4.1)
+(4.1) $\frac{v(f)}{\chi^{\prime}(f)}=$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\chi(f)-\chi^{\prime}(f)}{\chi^{\prime}(f)}\right)$
 $=$ $\frac{v(f)}{\chi(f)}\left(1+\frac{\chi_{0}(f)}{\chi^{\prime}(f)}\frac{\chi(f)-\chi^{\prime}(f)}{\chi_{0}(f)}\right)$
@@
 [^3] Since the product topology is in general not first countable and hence not obviously sequential, we avoid using the term finer for them, but only compare convergence for sequences.
-First we show that the a $d_{1}$-converging seqeunce converges in the $+$-topology.
+First we show that a $d_{1}$-converging sequence converges in the ++-topology.
 Lemma 4.7.
-Let $\mathfrak{a}_{\bullet}$,$\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ for $k\in\mathbb{Z}_{>0}$.
-If $\lim_{k}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{a}_{\bullet})=0$, then $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}$ in the $+$-topology.
+Let $\mathfrak{a}_{\bullet}$$\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ for $k\in\mathbb{Z}_{>0}$.
+If $\lim_{k}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{a}_{\bullet})=0$, then $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}$ in the ++-topology.
 Proof.
-We claim that if $\mathfrak{a}_{\bullet,k}\subset\mathfrak{b}_{\bullet,k}$ satisfy $\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{b}_{\bullet,k})=0$, then for any $v\in\mathrm{Val}^{+}_{X,x}$, we have $\lim_{k\to\infty}(v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{b}_{\bullet,k}))=0$.
+We claim that if $\mathfrak{a}_{\bullet,k}\subset\mathfrak{b}_{\bullet,k}$ satisfies $\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet,k},\mathfrak{b}_{\bullet,k})=0$, then for any $v\in\mathrm{Val}^{+}_{X,x}$, we have $\lim_{k\to\infty}(v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{b}_{\bullet,k}))=0$.
 For any $v\in\mathrm{Val}^{+}_{X,x}$, applying the claim to $\mathfrak{a}_{\bullet,k},\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet}$ and $\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet}$ separately yields
 $\lim_{k\to\infty}|v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{a}_{\bullet})|\leq\lim_{k\to\infty}|v(\mathfrak{a}_{\bullet,k})-v(\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet})|+\lim_{k\to\infty}|v(\mathfrak{a}_{\bullet,k}\cap\mathfrak{a}_{\bullet})-v(\mathfrak{a}_{\bullet})|=0.$
-Now it remains to prove the claim, which follows essentially from the same argument as [BLQ24, Proosition 3.12].
+Now it remains to prove the claim, which follows essentially from the same argument as [BLQ24, Proposition 3.12].
 We include a sketch for the reader’s convenience.
-Assume to the contrary that there exists $\epsilon>0$ and $v\in\mathrm{Val}^{+}_{X,x}$ such that $v(\mathfrak{a}_{\bullet,k})>v(\mathfrak{b}_{\bullet,k})+\epsilon$ for any $k$.
+Assume to the contrary that there exist $\epsilon>0$ and $v\in\mathrm{Val}^{+}_{X,x}$ such that $v(\mathfrak{a}_{\bullet,k})>v(\mathfrak{b}_{\bullet,k})+\epsilon$ for any $k$.
 Then for $k\in\mathbb{Z}_{>0}$, we may choose $\ell_{k}\in\mathbb{Z}_{>0}$ and $f_{k}\in\mathfrak{b}_{\ell_{k},k}$ such that
@@
 Then there exists a constant $M:=M(C)>0$ such that
-$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\leq M\cdot d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}).$ (4.2)
+(4.2) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\leq M\cdot d_{\infty}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}).$
 As a consequence, the $d_{\infty}$-topology is finer than the $d_{1}$-topology.
@@
 ∎
-To summarize, we have the following diagram for convergence of sequences
-$d_{\infty}\text{-convergence}\left\{\begin{aligned} &\implies\text{weak convergence}\\ &\implies d_{1}\text{-convergence}\implies+\text{-convergence}\implies\text{coffecientwise}\end{aligned}\right.$
+To summarize, we have the following diagram for the convergence of sequences
+$d_{\infty}\text{-convergence}\left\{\begin{aligned} &\implies\text{weak convergence}\\ &\implies d_{1}\text{-convergence}\implies+\text{-convergence}\implies\text{coefficientwise convergence}\end{aligned}\right$.
 Here the upper line is obvious; the first two arrows of the lower line follow from Lemma 4.8 and Lemma 4.7 respectively; and the last arrow follows from the fact that $\mathrm{DivVal}_{X,x}\subset\mathrm{Val}^{+}_{X,x}$.
@@
 This proves $\mathrm{Fil}^{s}=\mathrm{Fil}/\sim_{1}$.
-(2) Assume that $\mathfrak{a}_{\bullet,0}\neq\mathfrak{a}_{\bullet,1}$.
-We first assume that there exists $f\in R$ and $\lambda\in\mathbb{R}_{>0}$ such that $f\in\mathfrak{a}_{\lambda,0}\backslash\mathfrak{a}_{\lambda,1}$ and $f\notin\mathfrak{a}_{\lambda^{\prime},0}$ for $\lambda^{\prime}>\lambda$.
+(2) Since $\mathfrak{a}_{\bullet,0}\neq\mathfrak{a}_{\bullet,1}$, we may assume that $\mathfrak{a}_{\bullet,0}\not\subseteq\mathfrak{a}_{\bullet,1}$.
+Then there exist $f\in R$ and $\lambda\in\mathbb{R}_{>0}$ such that $f\in\mathfrak{a}_{\lambda,0}\backslash\mathfrak{a}_{\lambda,1}$, and $f\notin\mathfrak{a}_{\lambda^{\prime},0}$ for $\lambda^{\prime}>\lambda$.
 Take $t\in(0,1)$ and $\mu,\nu$ with $(1-t)\mu+t\nu=\lambda$.
@@
 If $\mu>\lambda$, then by the choice of $f$ and $\lambda$ we have $f\notin\mathfrak{a}_{\mu,0}$.
-So in either case, $f\notin\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$.
-This shows that $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,1}$.
-If $\mathfrak{a}_{\bullet,1}\nsubseteq\mathfrak{a}_{\bullet,0}$, then we can repeat the above construction to show that $\mathfrak{a}_{\bullet,t}\neq\mathfrak{a}_{\bullet,0}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$.
+In either case, $f\notin\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$.
+Thus $f\notin\mathfrak{a}_{\lambda,t}$ and hence $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,0}$.
+If $\mathfrak{a}_{\bullet,1}\nsubseteq\mathfrak{a}_{\bullet,0}$, then we can repeat the above construction to show that $\mathfrak{a}_{\bullet,t}\neq\mathfrak{a}_{\bullet,1}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ for any $t\in(0,1)$.
 Otherwise, $\mathfrak{a}_{\bullet,1}\subsetneq\mathfrak{a}_{\bullet,0}$.
-We claim that there exists $\lambda\in\mathbb{R}_{>0}$ and $\epsilon>0$ such that $\mathfrak{a}_{\lambda+\epsilon,0}\supsetneq\mathfrak{a}_{\lambda,1}$ and $\mathfrak{a}_{\lambda-\epsilon,1}\supsetneq\mathfrak{a}_{\lambda,1}$.
-Indeed if all the jumping numbers of $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ coincide then the claim is clear.
+We claim that there exist $\lambda\in\mathbb{R}_{>0}$ and $\epsilon>0$ such that $\mathfrak{a}_{\lambda+\epsilon,0}\supsetneq\mathfrak{a}_{\lambda,1}$ and $\mathfrak{a}_{\lambda-\epsilon,1}\supsetneq\mathfrak{a}_{\lambda,1}$.
+Indeed, if all the jumping numbers of $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$ coincide then the claim is clear.
 Otherwise, choose a jumping number $a$ of $\mathfrak{a}_{\bullet,1}$ with $\mathfrak{a}_{a,1}\subsetneq\mathfrak{a}_{a,0}$ which is not a jumping number of $\mathfrak{a}_{\bullet,0}$.
 Then there exists $a^{\prime}>a$ such that $\mathfrak{a}_{a^{\prime},0}=\mathfrak{a}_{a,0}$.
-Let $\lambda:=(a+a^{\prime})/2$ and $\epsilon:=\lambda-a$, then $\mathfrak{a}_{\lambda-\epsilon,1}=\lambda_{a,1}\supsetneq\mathfrak{a}_{\lambda,1}$ since $a$ is a jumping number and $\mathfrak{a}_{\lambda+\epsilon,1}=\mathfrak{a}_{a^{\prime},1}=\mathfrak{a}_{a,1}\supsetneq\mathfrak{a}_{a,0}\supset\mathfrak{a}_{\lambda,0}$.
-Thus we showed $\mathfrak{a}_{\lambda,t}\neq\mathfrak{a}_{\lambda,1}$ and hence $\mathfrak{a}_{\bullet,t}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$.
-Now it follows from Proposition 3.6 that $(\mathrm{Fil}^{s},d_{1})$ is convex.
+Let $\lambda:=(a+a^{\prime})/2$ and $\epsilon:=\lambda-a$, then
+$\mathfrak{a}_{\lambda-\epsilon,1}=\mathfrak{a}_{a,1}\supsetneq\mathfrak{a}_{\lambda,1}$
+since $a$ is a jumping number of $\mathfrak{a}_{\bullet,1}$, and
+$\mathfrak{a}_{\lambda+\epsilon,0}=\mathfrak{a}_{a^{\prime},0}=\mathfrak{a}_{a,0}\supsetneq\mathfrak{a}_{a,1}\supset\mathfrak{a}_{\lambda,1}$
+by construction. So we get
+$\mathfrak{a}_{\lambda,1}\subsetneq\mathfrak{a}_{\lambda+\epsilon,0}\cap\mathfrak{a}_{\lambda-\epsilon,1}\subset\sum_{\mu+\nu=2\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}=\mathfrak{a}_{\lambda,1/2},$
+and hence $\mathfrak{a}_{\bullet,1/2}$ is between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$. Now it follows from Proposition 3.6 that $(\mathrm{Fil}^{s},d_{1})$ is convex.
+Otherwise, $\mathfrak{a}_{\lambda,0}\subsetneq\mathfrak{a}_{\lambda,1}$ for any $\lambda\in\mathbb{R}_{>0}$.
 When $R$ contains a field, by Lemma 3.10, the curve $\gamma:[0,1]\to\mathrm{Fil}^{s}$, $t\mapsto\widetilde{\mathfrak{a}}_{\bullet,t}$ is continuous.
@@
 The first assertion follows from Proposition 3.27.
-Note that by 3.3, $(\mathrm{Fil}^{s}_{R,\mathfrak{m}},d_{1})$ is a rooftop metric space ([Xia23, Definition 5.2].
+Note that by Lemma 3.3, $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ is a rooftop metric space ([Xia23, Definition 5.2].
 It is easy to verify that a decreasing (resp.
-increasing) Cauchy sequence $\mathfrak{a}_{\bullet,k}$ in $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ satisfies $\cap_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ (resp.
-$\cup_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$).
-Therefore by Proposition 3.15, Lemma 3.25 and [Xia23, Proposition 5.4], $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is complete.
+increasing) Cauchy sequence $\mathfrak{a}_{\bullet,k}$ in $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ satisfies $\cap_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ (resp.
+$\cup_{k}\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$).
+Therefore by Proposition 3.15, Lemma 3.25 and [Xia23, Proposition 5.4], $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}$ is complete.
 ∎
@@
 (2) It follows from Definition-Lemma 2.2, Definition-Lemma 2.22 and Proposition 2.24 that $(\mathrm{Fil}^{s},\subset,\cap,\vee_{s})$ is a distributive lattice.
-By Proposition 2.23, the saturation from $\mathrm{Fil}$ to $\mathrm{Fil}^{s}$ is a joint morphism, which is clearly surjective.
+By Proposition 2.23, the saturation from $\mathrm{Fil}$ to $\mathrm{Fil}^{s}$ is a join morphism, which is clearly surjective.
 By Example 2.25, the map $\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}}\to\mathrm{Fil}^{s}$, $\mathfrak{a}\mapsto\widetilde{\mathfrak{a}^{\bullet}}$ is injective, and is a join morphism as the composition of two join morphisms.
@@
 5.1. Continuity properties of log canonical thresholds
-In this section, we assume $(R,\mathfrak{m})$ is a klt singularity over a field $\mathbbm{k}$ of characteristic 00.
+In this section, we assume $(R,\mathfrak{m})$ is a klt singularity over a field $\mathbbm{k}$ of characteristic 0.
 Proposition 5.1.
 Let $\mathfrak{b}_{\bullet,k}\in\mathrm{Fil}^{s}$ be a sequence of filtrations that converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$. Then
-$\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\liminf_{k\to\infty}\mathrm{lct}(\mathfrak{b}_{\bullet,k}).$ (5.1)
+(5.1) $\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\liminf_{k\to\infty}\mathrm{lct}(\mathfrak{b}_{\bullet,k}).$
 Proof.
@@
 Thus to prove (5.1), it suffices to show that
-$\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\lim_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}).$ (5.2)
+(5.2) $\mathrm{lct}(\mathfrak{a}_{\bullet})\leq\lim_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k}).$
 Assume to the contrary that $\mathrm{lct}(\mathfrak{a}_{\bullet})>\lim_{k}\mathrm{lct}(\mathfrak{a}_{\bullet,k})=:c$.
@@
 As $\mathfrak{a}_{\bullet,k}$ converges to $\mathfrak{a}_{\bullet}$ weakly, that is, $\mathfrak{a}_{\Lambda}=\cup_{k}\mathfrak{a}_{\Lambda,k}$, there exists $k_{1}\in\mathbb{Z}_{>0}$ such that $\mathfrak{a}_{\Lambda}=\mathfrak{a}_{\Lambda,k_{1}}$.In particular,
-$\Lambda\cdot\mathrm{lct}(\mathfrak{a}_{\Lambda,k_{1}})>c+\epsilon.$ (5.3)
+(5.3) $\Lambda\cdot\mathrm{lct}(\mathfrak{a}_{\Lambda,k_{1}})>c+\epsilon.$
 Now take $k_{2}\in\mathbb{Z}_{>0}$ such that $\mathrm{lct}(\mathfrak{a}_{\bullet,k})=\sup_{\lambda}\lambda\cdot\mathrm{lct}(\mathfrak{a}_{\lambda,k})<c+\epsilon$ for any $k\geq k_{2}$.
@@
 Proof of Theorem 1.4.
-(1) follows immediately from Proposition 5.1.
-To see (2), assume that $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the $+$-topology.
+Assertion (1) follows immediately from Proposition 5.1.
+To see assertion (2), assume that $\mathfrak{a}_{\bullet,k}\to\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the ++-topology.
 By [Blu18, Theorem B.1], there exists $v\in\mathrm{Val}^{<+\infty}_{X,x}$ such that $\mathrm{lct}(\mathfrak{a}_{\bullet})=\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}$.
@@
 $\limsup_{k\to\infty}\mathrm{lct}(\mathfrak{a}_{\bullet,k})=\limsup_{k\to\infty}\inf_{w\in\mathrm{Val}^{<+\infty}_{X,x}}\frac{A_{X}(w)}{w(\mathfrak{a}_{\bullet,k})}\leq\lim_{k\to\infty}\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet,k})}=\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}=\mathrm{lct}(\mathfrak{a}_{\bullet}).$
-(3) follows from Proposition 4.6. The proof is finished. ∎
+Assertion (3) follows from Proposition 4.6. The proof is finished. ∎
 6. Discussions
@@
 6.1. Relation to global results
-As was mentioned earlier, our definition for the metric $d_{1}$ is inspired by the Darvas metric $d_{1}$ on $\mathcal{H}(X,\omega)$, the space of smooth Kähler potentials of a compact Kähler manifold $(X,\omega)$, introduced in [Dar15], and is a local analogue of the metric $d_{1}$ on $R=R(X,-rK_{X})$, the section ring of a Fano manifold $X$, introduced in [BJ21].
+As was mentioned earlier, our definition for the metric $d_{1}$ is inspired by the Darvas metric $d_{1}$ on $\mathcal{H}(X,\omega)$, the space of smooth Kähler potentials of a compact Kähler manifold $(X,\omega)$, introduced in [Dar15], and is a local analogue of the metric $d_{1}$ on $R=R(X,-rK_{X})$, the section ring of a Fano manifold $X$, introduced in [BJ25].
 We now explain an approach to equip the global object with a structure in our setting via the cone construction.
@@
 We may replace $L$ by $rL$ for some $r\in\mathbb{Z}_{>0}$ such that $R$ is generated by $R_{1}$ as an $R_{0}$-algebra, and in the sequel, we will always assume $L$ is positive enough such that the above condition is satisfied.
-In this case, the cone $C(V,L)\coloneqq\mathrm{Spec}R(V,L)$ has a normal, isolated singularity at the vertex 00 defined by $R_{+}(V,L)\coloneqq\oplus_{m>0}R_{m}$.
+In this case, the cone $C(V,L)\coloneqq\mathrm{Spec}R(V,L)$ has a normal, isolated singularity at the vertex 0 defined by $R_{+}(V,L)\coloneqq\oplus_{m>0}R_{m}$.
 Denote the localization of $R(V,L)$ at $R_{+}(V,L)$ by $R$.
@@
 Denote the corresponding norm by $\chi_{t}$, then $\chi_{0}=\mathrm{ord}_{V}$ by definition, and for any $s\in R$, we have $\lim_{t\to+\infty}\chi_{t}(s)=v(s)$.
-For any $t\in[0,+\infty)$, by definition we have $\mathfrak{m}^{\lceil(1+t)\lambda\rceil}\subset\mathfrak{a}_{\lambda,t}$.
+For any $t\in[0,+\infty)$, by definition, we have $\mathfrak{m}^{\lceil(1+t)\lambda\rceil}\subset\mathfrak{a}_{\lambda,t}$.
 Moreover, since $\mathcal{F}_{v}$ is linearly bounded, there exists $C\in\mathbb{R}_{>0}$ such that for any $s\in R_{l}$, we have $v(s)\leq Cl$.
@@
 We do not go further in this direction.
-See also [Fin23] and the references therein for some related results.
+See also [Fin25a] and the references therein for some related results.
 6.2. Examples
-In this section we consider several examples. Note that we only work with linearly bounded filtrations, or equivalently, filtrations with positive multiplicity.
+In this section, we consider several examples. Note that we only work with linearly bounded filtrations, or equivalently, filtrations with positive multiplicity.
 6.2.1. The metric space $(\mathrm{Fil}^{s},d_{1})$ in low dimensions
@@
 If $R$ is a DVR, then $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\cong\mathbb{R}_{>0}$.
-Let $\pi\in\mathfrak{m}$ a uniformizer.
+Let $\pi\in\mathfrak{m}$ be a uniformizer.
 For $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, let $c\coloneqq\mathrm{ord}_{\pi}(\mathfrak{a}_{\bullet})\in\mathbb{R}_{>0}$.
@@
 6.2.3. The metrics on spaces of ideals.
-Similar to Theorem 1.1, one can also define a pseudo-metric $d_{1}$ on the set $\mathcal{I}_{\mathfrak{m}}$ by
+Similar to Theorem 1.1, one can also define a pseudometric $d_{1}$ on the set $\mathcal{I}_{\mathfrak{m}}$ by
 $d_{1}(\mathfrak{a},\mathfrak{b})=2\mathrm{e}(\mathfrak{a}\cap\mathfrak{b})-\mathrm{e}(\mathfrak{a})-\mathrm{e}(\mathfrak{b}).$
-By Rees’ Theorem, the Hausdorff quotient of $(\mathcal{I},d_{1})$ can be identified with $(\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}},d_{1})$.
-Note that with this distance, the inclusion $\mathcal{I}^{\mathrm{ic}}_{\mathfrak{m}}\hookrightarrow\mathrm{Fil}^{s}$ is not an embedding but just a contraction map, since in general we only have $\mathrm{e}(\mathfrak{a}\cap\mathfrak{b})\geq\mathrm{e}(\widetilde{\mathfrak{a}^{\bullet}}\cap\widetilde{\mathfrak{b}^{\bullet}})$.
+[…]
 Note that the spaces $\mathrm{Fil}^{s}$ and $\mathrm{Val}_{X,x}^{+}$ both allow an action by $\mathbb{R}_{\geq 0}$.
@@
 $\mathcal{I}_{\mathfrak{m},A}:=\{\mathfrak{a}^{c}\mid\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}},\ c\in A\},$
-where $A=\mathbb{Q}_{\geq 0}$ or $\mathbb{R}_{\geq 0}$. Clearly the following metric makes sense
+where $A=\mathbb{Q}_{\geq 0}$ or $\mathbb{R}_{\geq 0}$. Clearly, the following metric makes sense.
 $d_{1,A}(\mathfrak{a}^{c},\mathfrak{b}^{d}):=2\mathrm{e}(\mathfrak{a}^{c\bullet}\cap\mathfrak{b}^{d\bullet})-c^{n}\mathrm{e}(\mathfrak{a})-d^{n}\mathrm{e}(\mathfrak{b}).$
 Warning. These are in general two different metrics on $\mathcal{I}_{\mathfrak{m}}$.
+6.3. The structure of finitely generated filtrations
+Recall that a filtration $\mathfrak{a}_{\bullet}$ is finitely generated, or Noetherian, if the associated graded ring $\mathrm{gr}_{\mathfrak{a}_{\bullet}}R\coloneqq\oplus_{\lambda\in\mathbb{R}_{\geq 0}}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ is a finitely generated $\kappa$-algebra.
+The finite generation of valuation ideals has been extensively studied since [BCHM10]. See [LXZ22, XZ25] for recent progress based on the theory of special complements.
+As another application of Proposition 3.15, we show that finitely generated filtrations are dense in $\mathrm{Fil}_{R,\mathfrak{m}}$, which is the local analogue of [BJ25, Corollary 3.19].
+Proposition 6.3.
+The subset of finitely generated filtrations is dense in $\mathrm{Fil}_{R,\mathfrak{m}}$ with respect to $d_{1}$.
+Proof.
+The following construction is a local analogue of [Xu25, Definition-Lemma 3.56]. For $\mathfrak{a}\in\mathrm{Fil}_{R,\mathfrak{m}}$ and $k\in\mathbb{Z}_{>0}$, let
+$\mathfrak{a}_{\lambda,k}\coloneqq\left\{\begin{aligned} &\mathfrak{a}_{\lambda},&\text{ if }\lambda\leq k,\\ &\sum_{\vec{\mu}}\mathfrak{a}_{\mu_{1}}\cdot\ldots\cdot\mathfrak{a}_{\mu_{s}},&\text{ otherwise,}\end{aligned}\right$.
+where the sum is taken over all vectors $\vec{\mu}=(\mu_{1},\ldots,\mu_{s})\in\mathbb{R}_{>0}^{s}$ satisfying $\sum_{i}\mu_{i}\geq\lambda$ and $\mu_{i}\leq k$ for any $1,\ldots,s$.
+Note that the sum is finite since the set of jumping numbers of $\mathfrak{a}_{\bullet}$ contained in $(0,k]$ is finite.
+It is easy to check that $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}_{R,\mathfrak{m}}$.
+Moreover, $\mathrm{gr}_{\mathfrak{a}_{\bullet,k}}R$ is generated by elements of degree at most $k$, and hence is finitely generated again by the finiteness of jumping numbers.
+Clearly $\{\mathfrak{a}_{\bullet,k}\}$ is an increasing sequence and $\cup_{k}\mathfrak{a}_{\bullet,k}=\mathfrak{a}_{\bullet}$.
+So by Proposition 3.15, we know that $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet,k})\to 0$, which finishes the proof.
+∎
+Remark 6.4.
+In general, the saturation of a finitely generated filtration is not finitely generated.
+Hence, the above proposition says nothing about the structure or distribution of finitely generated saturated filtrations, which remains rather mysterious.
+Indeed, in [LX24], Y.
+Liu and Xu proposed several conjectures regarding the structure of Kollár valuations, which are special classes of finitely generated filtrations.
+See [ABB+23, LXZ22, Pen25] for some special cases.
+Moreover, in view of [LXZ22, Section 4] and [XZ25, Section 4.3], it seems that the specialty of quasi-monomial valuations inside a single dual complex is related to the geodesics.
 6.4. Questions
@@
 So it is natural to ask the following question:
-Question 6.3.
+Question 6.5.
 Let $\{\mathfrak{a}_{\bullet,k}\}_{k\in\mathbb{Z}_{>0}}$ be a decreasing sequence in $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ such that $\mathfrak{a}_{\bullet}\coloneqq\cap\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, do we have
 $\lim_{k\to\infty}d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet,k})=0?$
-The answer to the question is likely no, but we do not have a counterexample.
+We remark that convergence along decreasing sequences or nets has been studied in the complex analytic setting [Dar17, Proposition 6.11] and the global non-archimedean setting [BJ22, Theorem 12.5].
+The answer to the above question is likely no, but we do not have a counterexample.
 6.4.2. More structural questions
@@
 For example,
-Question 6.4.
-Is $\mathrm{Fil}^{s}$ always contractible? If not, is there a criterion in terms of the algebraic properties of $R$?
+Question 6.6.
+Is $\mathrm{Fil}^{s}$ always contractible? If not, is the contractibility related to the algebraic properties of $R$?
+Statement for conflict of interest. The author states that there is no conflict of interest.